A Bibliography of Publications in International Journal of Computational Geometry and Applications

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 August 2025 Version 1.53

Title word cross-reference

(1|1) [YLL18]. 1 [DE12, LPC00, Ngu12]. 1.5 [GLS10]. 14 [Res14]. 2 [ADM11, BS16, DEG^{+}03, Dey97, EMM98, ELPZ07, For95, kLkHsL^{+}23, ÓWW00, RW11, RR00]. 3 [AAH^{+}15, AK99, BHP01, BBCS99, BS16, BS17, CM11, CSY97, CK97b, FM99, GGBK21, HSS05, JJ06, LWŻ12, RW11, TW06, Zhu04a]. 30 [O’R97a]. 31 [O’R97b]. 32 [O’R97c]. 33 [O’R98]. 34 [AO98]. 35 [O’R99b]. 36 [O’R99a]. 37 [DO00]. 38 [O’R00a]. 39 [O’R00b]. 4 [BS17]. 40 [O’R00c]. 41 [O’R01]. 42 [MO01]. 43 [O’R02]. 44 [O’R03]. A [BXHN03]. c [BK17]. C^{1} [HREK07]. χ [BDH^{+}12]. d [AB09, AK99, BK02, Gav09b]. δ [BDH^{+}12]. E [BDH^{+}12]. ϵ [DGRS08]. \frac{2}{3} [WTX02]. K [BKN^{+}11, AKKS14, AGM^{+}12, CHU14, DHT15, DKMM23, ESS11, FN05, FS08, KK10, MNP^{+}00, MRM15, Pap99, Wan15, WZ16, WZ19]. L [BRD09]. L_{1} [WZ19, Wan15]. L_{2} [Rab05]. L_{∞} [PX15, PL01]. R^{d} [MRM15]. R^{N} [FKMW22]. n [AH19]. O [BS00]. O(n log^{∗} n) [Dev92]. O(n log n) [ADS00]. ω [BDH^{+}12]. Ω(n) [Dev92]. π/2 [BDD^{+}12]. r [LWŻ12]. θ [BvR19]. V [San09].

-Angle [BDD^{+}12]. -Approximation [LWŻ12]. -Block [San09]. -Center [WZ16, BKN^{+}11]. -Centerpoints [MRM15]. -Centroid [YLL18]. -Clustering [KK10]. -Colorability [AAH^{+}15]. -Colored [BS16, BS17]. -Complexes [ÓWW00]. -Connected [CK97b]. -Continuous [HREK07]. -Convex [BS00]. -D [CM11]. -Dimensional [AB09, AK99, BK02, Gav09b, JJ06].

-Enclosing [MNP^{+}00]. -Extensions [Ngu12]. -Flats [CHU14]. -Graphs [BvR19]. -Level [AGM^{+}12]. -Manifolds [Dey97]. -Maps [BS17]. -Means [FS08, WZ16]. -Median [WZ16]. -Modem [DHT15]. -Omino [AH19]. -Packed [BK17]. -Pairs [Pap99]. -patches [BXHN03]. -Piercing [AK99]. -Plane [DE12]. -Ranges [FN05]. -Sampling [DGRS08]. -Searcher [LPC00]. -Sets [ESS11]. -Space [CSY97]. -splines [BXHN03]. -Star [LWŻ12]. -Visibility [BRD09].

/Max [EHP18].

2-Approximate [GSZ11]. 2-Central [HZX24]. 2-Centres [DK08]. 2-Dot [JPV21]. 2-Manifolds [DMMH11]. 2-Pseudomanifolds [DMMH11]. 2-Trees [BFL21].

3-Coloured [BHLL10].

Abstract [BK14, BKL17, MMR01]. Acyclic [DGL^{+}00, Fra08]. Adaptive [BD05, CL17, EW00]. Advancing [HS02]. Advantages [AAH^{+}11]. Aggregate [Wan15, WZ19]. Aggregate-MAX [Wan15]. Aggregated [GJS09]. Algebraic [CCD06, MS07a, SV01]. Algorithm [AL11, AKM^{+}17, ADS00, ACDL02, AFN11, ACM01, BGK^{+}09, BL03, BM02, BCHS07, Che10, CER97, DN97, EFKP13, HZX24, HH12, KYZ14, LSS02, LWŻ12, LHX24, MMNM07, MS07a, NY98, OGB11, Sha01, SI94, TV01, THI99, TO21, TMPD97, TW06, WZ20, WTX02, WDBB09, dF18, CL93, TMPD95]. Algorithms [Als97, AR19, AS01, ACKT01, BD05, BG05, BBL08, CY17, CD03, CHL^{+}04, CSX05, CFM^{+}01, DDCN13, Dey97, EFS09, ECHS11, FG04, For95, Gui22, IST20, JW24, JH04a, LSS98, Maf14, MS10, MTT99, MPW05, MS14, RW11, SV15, STÜ07, WCMS04, Wu09, ZP01, Dev92].

Aligned [BKN^{+}11, iN23]. Almost [AACT17, DR02, KK10, WLW01]. Alternating [KKY00]. Amidst [BL03, CCK^{+}06]. Among [CW12b, LYW97]. Analyses [STÜ07]. Analysis [BDIZ03, CWW08, Cho99, FOG00, Gui22, SOR06]. Anchored [DBGV06, FSS^{+}97]. Angle [BDD^{+}12, DE12, KLV21, KV23, Mit97]. Angles [CDRR05, FMHT14, WG21]. Angularity [DMOW98]. Anisotropic [SYI00]. Anisotropy [ACFV10]. Annulus [Cha02, DBHM^{+}03]. Any [CM10, VO98]. Application [CEK^{+}07, DG99, Epp97, Gui22, MHW00, NS09, TW06, KNA94]. Applications [Ata99, BS12, BCHS07, CHW02, CLX03, CHW^{+}08, Cho99, DBGV06, DK06, FIS08, IM12, KTT02, NN09, PL01, SPPK08, WCMS04, Wu09]. Approach [BMT00, CMO03, CKMK03, KT03, MC91, MS06, MH00, PL04, Pet98, SM06]. Approaches [CHL^{+}06]. Approximate [AMV13, Ber05, BDH^{+}04, CJVW12, CSY97, DR20, GSZ11, KS11, MS07a, MS10, MST13]. Approximating [AS18, BYM^{+}18, Cha02, CD03, NN09, VO98, Zhu97, Zhu04a]. Approximation [AFN11, ACM01, BXHN03, BGK^{+}09, BG05, BCHS07, BK17, DDCN13, DK08, EFS09, GRS08, HH08, LWŻ12, LR00, MNP^{+}00, MHS07, WTX02, WCMS04, ZP01]. Approximations [dFdSdF17]. Arabesque [Nab21]. Arbitrarily [MR03]. Arbitrary [AM07, HQYD22]. Arc [WJA20]. Arcs [GBRT13]. Area [BDJ10, BHLM03, BHLL10, CDG^{+}09, Fra08, GR03a, GSa20, HL98, HSKK98, KPS13, MGR09, TWC06]. Area-Efficient [GR03a]. Areas [AACKM11, KSN99]. Arithmetic [Gav09b, JS09, dF18]. Arithmetic-Degree [dF18]. Arm [Kan97b]. Arrangement [BEW03, MS07a]. Arrangements [ACGK17, GHH^{+}98, GM98, HL04, KYZ14, LHHHP03, SS11, dBHOvK97]. Art [BI21, CJK^{+}06, KM11, WK07]. Assembly [GM99, GHH^{+}98, JMM98]. Assessment [San09]. Assignment [Mit00]. Asteroidality [CWW02]. Asteroidality/Tubularity [CWW02]. Asymptotically [RS11]. Attractors [MF06]. Attributes [BDIZ03]. Augmenting [WZ20]. Author [Ano97, Ano98, Ano99, Ano00, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Ano07, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13a, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, Ano23]. Automatic [BBCS99, KT03]. Aware [EFKM08]. Axes [WG21]. Axis [CDKW05, EMM98, GRS08, MGD15, Seg99, SFM07, WIEH05, Zhu97]. Axis-Parallel [CDKW05, MGD15, Seg99, Zhu97].

Back [kLkHsL^{+}23]. Balanced [AGLN03, KK05, KU10]. Ball [CLRW10, FG04]. Ball-Map [CLRW10]. Balls [BG11a, FG04, Gui22, NN09]. Bands [HH08]. Barrier [JW24]. Based [ADM11, AL01, ACKT01, BBR09, Ber00, CSX05, CW12a, CGJS11, DGRS08, EFKP13, GLL^{+}99, GGBK21, HH08, HH12, KS05, MF06, MH00, Sch00, Tou05]. be [CCMS19]. Beltrami [Xu06]. Bends [ECHS11, EC15]. Benefit [FOG00]. Bernstein [Pro22]. Best [BDE02]. Between [AS08b, BHP01, Ber05, Bes02, CLR07, CLRW10, GMMW19, Gui22, Tan02, Wan09, CT97]. Beyond [AMV13]. Bézier [Rab05, ZWG06]. Biarc [HH08]. Bichromatic [CGG^{+}12, PS19]. Bilateral [MG98]. Binary [DK12]. Bipartite [JMMR23]. Bipartitions [DK99]. Bisecting [BKL17]. Bisection [KLV21]. Bisector [EHP18]. Bisectors [FR98]. Bites [DG98]. Bitmap [KC97]. Black [BD05]. Black-Box [BD05]. Block [CHW^{+}08, San09]. Blue [AC01, BK18, HSS05]. Boat [NS09].

Boat-Sail [NS09]. Bodies [Sit06]. BOOLE [KMG^{+}01]. Boolean [KMG^{+}01]. Bottleneck [CARB15, DKMM23]. Bound [Ata99, BS05, BHLL10, DHT15, KS02, KPS13, Ror19]. Boundaries [DMMH11]. Boundary [AAH^{+}11, DG99, KU10, KMG^{+}01, NZ06, STYK01]. Boundary-Optimal [NZ06]. Bounded [AR19, BL03, BSX09, CL13, DK08, FOX08, GOG11, LW04, RSS^{+}05]. Bounded-Velocity [DK08]. Boundedness [BM12]. Bounds [Afs13, AHM^{+}06, BK17, CMO03, CER97, DG16, LOS01]. Box [BD05, FM99, ZE02]. Boxes [AK99, SU13, Zhu97]. Braid [Sto21]. Branching [HSKK98]. Bregman [AMV13]. Brep [Van91, MG98]. Brep-index [Van91]. Bridge [BG05, Tan02]. BSP [SPP08]. BSPs [DMS10]. Buffer [DG01]. Buildings [EHP18]. Butterfly [KS99].

CAD [BBCS99]. Calculations [BBR09]. Cameras [KM11]. Cartesian [LSB04, SOR06]. Cartograms [DMS10]. Cascading [BFS01]. Case [DKS05, TV01]. Catalog [ADM11]. Catalog-Based [ADM11]. Cell [ACGK17, HREK07]. Cells [GHH^{+}98]. Cellular [LSB04]. Center [BHLM03, BKN^{+}11, GKS99, WZ16, WZ18]. Centerpoints [MRM15]. Centers [AKKS14]. Central [ADS00, HZX24]. Centre [DK06]. Centres [DK08]. Centroid [ESS11, YLL18]. Chains [BBB^{+}10, DLMS13]. Characteristics [GW04]. Chessboard [SPPK08]. Chief [Lee03]. Chimneys [CDD^{+}12]. Choices [PW01]. Circle [BFMFP^{+}14, BE00, Epp97, KKS05, WTX02]. Circles [AS01, BCD^{+}00, HL04, KKS05, SW01]. Circular [AAH^{+}11, DH13, WJA20]. City [BKC09, GSW08]. Class [RS11]. Classes [BV05]. Classification [AGM^{+}12]. Close [SYI00]. Closed [BKL17, GSa20, HREK07, SVY16]. Closest [Bes03]. Cloud [MNG04]. Clouds [ULVH10]. Clustering [BVL11, BBG^{+}11, CSX05, KK10, MMNM07, WCMS04]. Clusters [Guh05]. Collections [Sit06]. Collision [GR03b, KSS02]. Color [DGN09]. Color-Spanning [DGN09]. Colorability [AAH^{+}15]. Colored [BS16, BS17, DP02]. Coloring [FK18, dBLM^{+}19]. Colorings [AS08a]. Coloured [BHLL10]. Column [AO98, DO00, MO01, O’R97a, O’R97b, O’R97c, O’R98, O’R99b, O’R99a, O’R00a, O’R00b, O’R00c, O’R01, O’R02, O’R03, O’R04a, O’R04b, O’R06, O’R07]. Combinations [KMG^{+}01]. Combinatorial [AHO^{+}14, CR01, CER97, DFL^{+}18, FG04, MS06, Nab21, SZP10]. Common [Rab05, SU13, Wan09, YCCV17]. Commuting [BBG^{+}11]. Compact [BBCK05, Kan97a]. Compass [KL10a, VR04]. Compatible [CLR07, CLRW10]. Competitive [BDDT17, GR10]. Complete [BMKS00, BG14, Emi98, OGB11]. Completion [ZG06]. Complex [DGRS08, ELPZ07, GRS08]. Complexes [ALS12, CC06, EW00, GK20, Maf14, ÓWW00]. Complexity [BBR09, GR10, GM99, GMV99]. Complicated [RS07]. Component [CWW08]. Compressive [GIPR12]. Computable [CCK^{+}06]. Computation [BFS01, EMM98, FR98, GC97, Gui22, Hiy08, LS08, Löf11, vKLSW18]. Computational [AO98, AAH^{+}11, DO00, JS09, MO01, O’R97a, O’R97b, O’R97c, O’R98, O’R99b, O’R99a, O’R00a, O’R00b, O’R00c, O’R01, O’R02, O’R03, O’R04a, O’R04b, O’R06, O’R07, Pet98]. Computing [AKS^{+}12, AKKS14, AH11, ABD^{+}11, AS08b, AL01, AEK05, BSC99, BSC00, DG13, Bes03, BMT99, BCD^{+}00, BL03, BMSS11, BHLM03, CK97a, DMOW98, DR02, Emi98, FSS^{+}97, Gav09b, GKK^{+}10, GKS99, KG14, Kir07, KS99, KYZ14, Kra20, MB02, MR03, TV01, WLW01, WZ18, WNGK^{+}12, dF18]. Concepts [PW01]. Conceptual [SOR06]. Condition [KU10]. Conditional [BK17]. Configuration [HLM99]. Configurations [BK07]. Conflict [AS08a, FK18, dBLM^{+}19]. Conflict-Free [AS08a, FK18, dBLM^{+}19]. Conflicting [SZP10]. Conforming [MMG01]. Congruence [BK02]. Conic [GW04]. Conjectures [MRM15]. Conjugacy [Sto21]. Connected [AACT17, CK97b]. Connecting [AC01, BG05]. Connection [ACGK17]. Conquer [PL04]. Consistency [SOR06]. Constrained [DDL^{+}10, GOG11, GBRT13, KS99, RSS^{+}05, TW06, WZ16, ZG06, DEG^{+}03]. Constraint [GBRT13, JW24, JTNM06, SM06, SZP10, TW06, ZG06]. Constraints [AAMT15, BvR19, CARB15, CWW02, MS06, VB05, Yan06, YLL18, DEG^{+}03]. Constructing [BDGT13, CDWK01, DN97, GSW08, GOG11, THI99]. Construction [BKC09, BET99, GSZ11, HDY07, LW04, LHHHP03, MSB19, Wen02]. Constructive [Goo98]. Contact [FPNZ98, LM97]. Contain [BSX09]. Containers [AS18]. Containing [EEM11, KS13]. Containment [BHP01]. Continuous [BDBF^{+}14, EFS09, HREK07, WIEH05]. Contours [DG03, HSKK98]. Contraction [Goo98]. Contractions [AGL09]. Controlled [HL04]. Convex [AH11, AS18, AFN11, BRD09, BHLO11, BBC^{+}02, BDM^{+}20, BHLM03, BS00, Cha12, CWKC98, CDWK01, CL17, CT97, Cho99, CK97b, DKS05, DKMM23, Emi98, GSa20, GHH^{+}98, HS02, HDY07, KS02, KPS13, LR00, MS99, MGR09, MHW00, NY98, PS19, RR00, Sha01, TWC06, TO21, Tat23, VO98, Żak10, Zhu97, KNA94]. Convexity [Ror19]. Convolution [MS07b]. Coordinate [Yan06]. Coresets [FS08, dFdSdF17]. Corners [DW02]. Corrigendum [THI99].

Cost [FOG00, LWZ17, WKG10]. Cost/Benefit [FOG00]. Counting [AB09]. Countries [SV10]. Counts [BHLO11]. Cover [AACKM11, BS05, BS00, CHW02, DFLON12, EC15, GSa20, KPS13]. Coverage [AMP10, ACS18, LWZ17, iN23]. Covered [GHH^{+}98]. Covering [ACFV10, Col04, JW24, Jia15, Kei97, KBA11, LWŻ12, Mit97, Por09]. Creation [ESG98]. Criteria [AAK^{+}06]. Critical [DGRS08]. Cross [EW00]. Cross-Sections [EW00]. Crossing [CARB15, DE12, KSN99, Pap99, TSN97]. Crossings [KKY00]. CSG [MG98]. Cubes [ABD^{+}18]. Culling [DP03]. Cumulative [Gui22]. Curvature [Aj24, BL03, CGJS11, GBRT13, Maf14]. Curvature-Based [CGJS11]. Curvature-Constrained [GBRT13]. Curve [BD05, CGJS11, FR98, GGBK21, HL97, HV91, Sch16]. Curve/Curve [FR98]. Curved [AS08b, Sha01]. Curves [BKL17, BK17, CD03, DW02, GSa20, GMMW19, HREK07, HH08, MS07a, SVY16, SV01]. Curvilinear [APS00]. Cut [TO21]. Cut-Locus [TO21]. Cutting [DL06, DH13]. Cycles [AFK^{+}10, Dey97, DL07, KKY00, WNGK^{+}12]. Cyclic [ACW22]. Cylinder [Cha02, FSS^{+}97]. Cylindrical [Ber04, Zhu04a]. Cylindricity [DP03].

D [BBCS99, DEG^{+}03, ADM11, BS16, CM11, EMM98, ELPZ07, FM99, For95, GLS10, GGBK21, HSS05, kLkHsL^{+}23, RW11, RR00, TW06, Zhu04a]. D-Range [kLkHsL^{+}23]. Dams [SV10]. Data [ACC^{+}12, AKKS14, ALS12, BYM^{+}18, CSX05, CW12a, EGS08, FIS08, GJS09, JS09, MTT99, MNG04, Tou05, WCMS04]. Database [Bar98, JMM98]. Dataflow [SPP08]. DBSCAN [dBGR19]. DCEL [Bar98]. Deceiving [San09]. Decision [AMM^{+}98]. Decompose [TW06]. Decomposition [CR01, FM01, GGBK21, HL98, JTNM06, KS02, SM06, WK07, WDBB09, ZG06, FM97]. Decompositions [Sha97a, Sha97b]. Deficiency [Sha01]. Definition [vKLSW18]. Deformation [CC06]. Deforming [Ber04]. Degeneracy [MS22]. Degenerate [MSB19]. Degree [AHO^{+}14, AR19, BSX09, HLW13, LW04, Rab05, dF18]. Delaunay [ABG^{+}09, ACH^{+}12, BDG13, BDG14, BSX09, BDDT17, Dev02, DEG^{+}03, DN18, For95, GJ21, GOG11, LS08, MPW05, Müc98, MMG01, RW11, STÜ07]. Deletion [AFK^{+}10, Dev02]. Density [CSX05]. Density-Based [CSX05]. Departure [San09]. Dependent [GJ21, MJ12]. Depth [DFL^{+}18, KMW00]. Depth-First [KMW00]. Derived [GJS03]. Descending [AL11]. Design [AAMT15, PW01, SOR06]. Destroying [SV10]. Detect [Dey97]. Detecting [BBG^{+}11]. Detection [CWW08, GR03b, KSS02, MSB19, MS22, Wu09]. Determinant [CKMK03]. Determination [LM97]. Determine [VB05]. Determined [BK07]. Determining [Che98, Gav09a]. Detour [WNGK^{+}12]. Developments [SU13]. Diagram [BKC09, BS12, BBB^{+}10, DG98, DBGV06, ETT08, Gav09b, GJ21, GSW08, HDY07, KS05, KKS05, NS09, PL01, PL04, PD13, PX15, SPPK08, SV16]. Diagrams [AAC^{+}99, AGMR98, BC06, BK14, BKL17, EH19, GJS03, MMR01, Sug92, SI94, VO98]. Diameter [Cha02, MB02, Poo09, Jan93]. Diameter-4 [Poo09]. Diameters [Als97]. Diamond [BSX09]. Diamonds [BDE02]. Differential [CP05]. difficult [Dev92]. Digital [BBCS99]. Dihedral [KV23]. Dilation [AFK^{+}10, CL13, DG16, EBGK^{+}07, GKK^{+}10]. Dilation-Bounded [CL13]. Dilation-Optimal [AFK^{+}10]. Dimension [CWW08, CVY11, VO98]. Dimensional [AB09, AS18, AK99, BSC00, BK02, CD03, Emi98, Gav09b, HQYD22, JJ06, KS05, Kir07, Müc98, dBGR19].

Dimensions [AM07, ALS12, BBCK05, DB92, EEM11, HDY07, IMTI02]. Directed [DGL^{+}00, Fra08]. Direction [JJ10, Ngu12]. Direction-Length [JJ10, Ngu12]. Directional [CvO01, FOX08]. Directions [BNS10, VR04]. Disc [CCK^{+}06]. Disconnected [BK14]. Discrete [AKS^{+}12, BDIZ03, BBB^{+}10, DFLON12, DDCN13, EFS09, WKG10, WZ20, WCLS07, Wu09, Xu06]. Discs [AS08a, CWKC98]. Disjoint [KBA11]. Disk [BDJ10, DG13, BHLL10, DFLON12, KS13, dFdSdF17]. Disk-Shaped [DG13]. Disks [AFN11, BDP08, BVL11, CDG^{+}09, CDJ^{+}15, DDCN13, Gui22, iN23]. Dispersion [BGK^{+}09]. Dissections [Żak10]. Distance [AKS^{+}12, AS08b, BHP01, BBB^{+}10, BKST00, BK17, DGRS08, GMMW19, Gui22, KS11, Kra20, Maf14, MJ12, Yan06, YLL18]. Distances [BK07, Cha01, KN20]. Distant [AEK05]. Distributed [Gui22, LSS98, LDHX20]. Distribution [BK07, Gui22]. Distributions [MTT99]. Divide [PL04]. Division [HL98]. Dog [DG98]. Domain [GGBK21, MS99]. Dominance [GJSD97]. Dominating [CDJ^{+}15]. Doo [WQS05]. Door [KZ10, LPC00]. Dot [JPV21]. Double [BFMFP^{+}14, GKS99]. Double-Ray [GKS99]. Doubly [DMMH11]. Drawing [BMT00, BGT99, DE12, DGL^{+}00]. Drawings [CK97b, Fra08, GR03a, HLW13, MHN06, NPR17, Sud04]. Duality [ABR14]. Dynamic [BI21, BG14, Cha12, DBGV06, EGS08, FIS08, IST20, kLkHsL^{+}23, LM97, dBLM^{+}19]. Dynamically [GM98]. Dynamization [CT92].

Easy [DR02]. Eccentricity [DK06]. Edge [AFK^{+}10, AGL09, BHLO11, BFL21, CARB15, Che98, GHN^{+}03, HS02, KLV21, kLkHsL^{+}23, SM00, Tan99]. Edge-Crossing [CARB15]. Edge-Length [BFL21]. Edges [AT18, GMMW19]. Editor [CL09, DBKU14, Aga99, Asa09, Bar05, Bar13, Efr08, Fle06, For97, Her01, Hon18, Kim09, KS07, Lee03, Mit04, Rok09, Sug03, Tam03, Ten00, Tok02, Tok19, Zha07, Zhu04b, dBS02]. Editors’ [CÜ05, AV14, AF98, AC08, AMS97, ANO13b, CHL13, CO12, GM06, Gav05, kHsLtT23, HN11, HV12, KS16, LM98, MR05, SK08, dBDE17]. Efficiency [FOG00]. Efficient [ACKT01, AM07, ALS12, CD03, Dey97, GR03a, GJS09, HZX24, KNA94, KC97, LW04, LM97, LHX24, LR00, MS22, Tat23, VB05, WCMS04, Wu09, WDBB09, ZP01]. Element [MHW00]. Elements [DNW^{+}09]. Eliminating [HV91]. Ellipses [ETT08]. Ellipsoids [SYI00]. Embeddability [BV13, DDL^{+}10]. Embedded [ADF13, CP05]. Embedding [ADF13, BFMFP^{+}14, DL07, EBGK^{+}07]. Embeddings [KK05]. Empty [DBHM^{+}03, FSS^{+}97, KS13, MR03]. Enclosed [MGD15]. Enclosing [BMSS11, Cha02, FG04, MNP^{+}00, NN09]. Enclosure [BMSN19, GJSD97]. Energy [EFKM08]. Energy-Aware [EFKM08]. Engineering [FPNZ98, TV01]. Entities [vKLSW18]. Enumerating [Cha01, CR01, IMTI02]. Enumeration [Tat23, KNA94]. Envelopes [CNTV10]. Environment [ABC^{+}15, Bar98, CL93]. Environments [DEH^{+}05, LM97]. Equilateral [ADD^{+}13]. Equitable [BK18]. Equivalence [APS00]. Equivalent [ÓWW00]. Errata [EC15, Sha97a]. Error [BYM^{+}18, CMO03, KL10a]. Error-Prone [KL10a]. Estimate [KLV21]. Estimating [CFL15, MNG04, RW11]. Estimation [MNP^{+}00]. Euclidean [AR19, BC06, CSY97, DN97, DK08, EFS09, ETT08, Gav09b, Gui22, KKS05, dBGR19]. Euler [GK20]. Evaluation [FPNZ98, KMG^{+}01, MS22, WQS05]. Evaluations [DP03]. Evasion [ABC^{+}15, GLL^{+}99]. Even [BDH^{+}04]. Every [DE12]. Exact [AL11, AS01, BG05, BFS01, DD00, ETT08, Gui22, RR00, Tat23]. Exchangeability [Sto21]. Existence [Löf11]. Expansive [HLM99]. Expected [ELPZ07]. Expensive [KN20]. Experimental [DGL^{+}00, LHHHP03]. Explicit [Gav09b]. Extended [TO21]. Extending [DMMH11]. Extensions [Ngu12]. Exterior [BRD09]. External [CFM^{+}01, Nek13]. External-Memory [CFM^{+}01]. Externally [BMT99]. Extra [BM02]. Extract [GW04]. Extracting [DG03]. Extraction [HREK07]. Extraneous [HV91]. Extreme [Guh05]. Eye [NŻ20].

Face [AHO^{+}14, BHLO11, DMMH11]. Faces [Res14]. Facets [CR01]. Facility [BMKS00, BKST00, DK06]. Factor [WTX02, WNGK^{+}12]. Factor- [WTX02]. Families [Fra08]. Far [AAMT15]. Far-Field [AAMT15]. Faraway [LS08]. Farthest [BD05, PD13, WZ19]. Fast [DN97, DW02, FS08, GMMW19, HH12, MMNM07, MSB19, Nek13, TW06, ZE02]. Faster [Epp97, GSW08, dBGR19]. Fault [BI22]. Fault-Tolerant [BI22]. Feature [CCD06, JH04b, RW11]. Features [GIPR12, JMM98]. Fidelity [Mit00]. Field [AAMT15]. Filling [GGBK21]. Finding [ADS00, AM07, BD05, BDGW10, BG05, CWKC98, CM10, DKMM23, EEM11, FKNN17, FMR05, KZ10, KS13, LYW97, LLCC11, Mit97, Tan02, YCCV17]. Finite [CFL15, MHW00]. Finite-Element [MHW00]. First [KMW00]. Fitting [AAK^{+}06, CW12a, Da 11, ULVH10]. Fixed [BBL08, CVY11]. Flashlight [LSS02]. Flats [CHU14, Da 11]. Flexible [Sch16]. Flipping [GHN^{+}03]. Flips [AHO^{+}14]. Floating [Gav09b, JS09]. Floating-Point [Gav09b, JS09]. Flooding [NZ06, SV10]. Floodlight [BGL^{+}97]. Floodlights [AECSU98, BDBF^{+}14]. Flow [DGRS08, GJS03, GRS08, MH00]. Flow-Complex-Based [DGRS08]. Folding [ADD^{+}13, ABD^{+}18, BDGT13, FOX08]. Footprints [EHP18]. Forest [kLkHsL^{+}23]. Forests [KK05]. Foreword [Aga99, AV14, AF98, AC08, AMS97, Asa09, ANO13b, Bar05, Bar13, CHL13, CL09, CO12, CÜ05, DBKU14, Efr08, Fle06, For97, GM06, Gav05, Her01, kHsLtT23, HN11, Hon18, HV12, Kim09, KS16, KS07, LM98, MR05, Mit04, Rok09, Sug03, SK08, Tam03, Ten00, Tok02, Tok10, Tok19, Zha07, Zhu04b, dBS02, dBDE17]. Form [APS00, CM11, HREK07, MG98]. Formed [Sha99, Sit06]. Four [AHO^{+}14]. FPT [ECHS11, EC15]. FPT-Algorithms [ECHS11]. FPTAS [Kir07]. Frames [MS03]. Frameworks [JJ10, Ngu12, OP10]. Fréchet [AKS^{+}12, BBB^{+}10, BK17, GMMW19, Sch16, SVY16]. Free [ACCS04, AS08a, CM11, FK18, HREK07, MG98, MSB19, dBLM^{+}19]. Free-Form [CM11, HREK07, MG98]. Friend [BDE02]. Fully [kLkHsL^{+}23, dBLM^{+}19]. Fully-Dynamic [dBLM^{+}19]. Function [CW12a, Gui22, JJ06]. Functions [Aj24, BKST00, Kra20]. Furthest [MMR01].

Gabriel [KG14]. Galleries

[BI21, CJK^{+}06, KM11]. Gallery [WK07]. General [BCHS07, Emi98, FKNN17, IMTI02]. Generalization [Zer12]. Generalized [CHL^{+}06, HH12, LOS01, WKG10, Wen02]. Generalizing [BV05]. Generation [MTT99, Sch00, TW00]. Generic [JJ10, Ngu12]. Geodesic [AHK^{+}14, BVL11, FKMW22]. Geodesic-Preserving [AHK^{+}14]. Geodesics [Tat23]. Geoexploration [PW01]. Geometric [APS00, AMM^{+}98, AHM^{+}06, AGR16, BGK^{+}09, BI22, BFS01, CLLP09, CS06, CDK01, CHL^{+}04, CSX05, CHL^{+}06, Che10, CFM^{+}01, EHP19, FOX08, GKK^{+}10, GJ21, GW04, Guh05, GJS09, GIPR12, JMMR23, JTNM06, KL10b, KTT02, LSS98, MST13, MTT99, MJ12, Pet98, SOR06, Sha97a, Sha97b, SZP10, TV01, Tou05, TW06, ULVH10, VB05, XYZK10, ZG06, dFdSdF17]. Geometrical [SM06]. Geometry [AO98, CP05, DO00, ESG98, Goo98, JS09, MO01, O’R97a, O’R97b, O’R97c, O’R98, O’R99b, O’R99a, O’R00a, O’R00b, O’R00c, O’R01, O’R02, O’R03, O’R04a, O’R04b, O’R06, O’R07, WCLS07, Wu09]. Geosheet [LSS98]. Ghost [CDD^{+}12]. Global [JJ10, Maf14, Ngu12, Yan06]. Good [DB92, VR04]. GPDOF [TW06]. Graph [ACC^{+}12, ABG^{+}09, BMT00, BGT99, DE12, NPR17, ÓWW00, Roy16]. Graphics [HHMK14]. Graphs [ADD^{+}13, ABG^{+}09, ACW22, ADF13, ABR14, BDJ10, BV13, BEW03, BvR19, BS00, CK97b, DGL^{+}00, DL07, EBGK^{+}07, EHP18, FM99, FK18, Fra08, DDL^{+}10, GKK^{+}10, HH12, JMMR23, JPV21, KL10b, KG14, MHN06, SM00, Tou05, dF18, dFdSdF17, BDD^{+}12, BDH^{+}12]. Greedily [NPR17]. Greedy [GSZ11]. Grid [BFMFP^{+}14, CY17, CK97b, DIL10, EvKSS15, KNA94]. Grid-Unfolding [CY17]. Grids [EW00]. Group [SM06]. Groups [vKLSW18]. Growing [CM10]. guarantee [FMR05]. Guaranteed [CMO03]. Guard [BRD09, FLMS18, THL98]. Guarding [BI21, BNS10, CJK^{+}06, DKK09, KM11]. Guards [AMP10, PLC02, Tan99]. Guest [Aga99, AV14, AF98, AC08, AMS97, Asa09, ANO13b, Bar05, CHL13, CO12, CÜ05, Efr08, Fle06, GM06, Gav05, Her01, kHsLtT23, HN11, Hon18, HV12, Kim09, KS16, KS07, LM98, MR05, Rok09, Sug03, SK08, Tam03, Ten00, Tok02, Tok19, Zha07, Zhu04b, dBS02, dBDE17]. Guided [DNW^{+}09].

Half [Vig12]. Half-Planes [Vig12].

Hamilton [KKY00]. Hamiltonian [Nar99]. Hard [BHP01, BG11a, BZ14, BDH^{+}04, GKK^{+}10, Roy16]. Hardness [KG14, MHS07]. Harm [BMKS00]. Hausdorff [AS08b, BHP01, KS11, Kra20, PL04, PX15]. Hazard [Gui22]. HDBSCAN [dBGR19]. Heavy [AHP08]. Helly [CPRS18]. Helly-Type [CPRS18]. Hexahedra [KV23]. Hexahedral [Sch00]. Hidden [GMV99]. Hidden-Surface [GMV99]. Hierarchical [AM07]. Hierarchy [Ber04]. High [ALS12, HQYD22, HLW13, Mit00, MH00]. High-Degree [HLW13]. Higher [ABG^{+}09]. Hinged [CVG^{+}07]. Histogram [FM97]. Holes [SM00]. Homeomorphic [ACDL02]. Homeomorphism [CLRW10, ÓWW00]. Homologous [Dey97]. Homology [ACW22, CFL15]. Homothetic [AK99]. Homotopic [CJVW12]. Homotopy [SFM07]. Homotopy-Preserving [SFM07]. Horizons [AEK05]. Hull [ACCS04, CWKC98, KPS13, NY98]. Hulls [Cha12, Emi98, Pet98, RR00]. Hybrid [CKMK03]. Hypercube [Ata99]. Hypersphere [BM12].

I/O [Afs13]. Identification [CCD06].

Identifying [BBR09]. if [DR02]. II [JH04b]. Ill [BBR09]. Ill-Posed [BBR09]. Illumination [AECSU98, DHT15, EFKM08]. Image [ACKT01, CWW02, WCLS07]. Immobilizing [CVG^{+}07, CSU99]. Implementation [AM07, Emi98, FS08, MMNM07, Müc98]. Implicit [ESG98, HYSC18]. Impossibilities [BHMW11]. Imprecise [AKS^{+}12, GLS10, Löf11]. Improved [Afs13, Ata99, BM02, BK17, CY17, EvKSS15, GC97, Goo98, KLV21]. Improvement [FOG00]. Improving [ACH^{+}12, Tou05]. Inaccurate [CK97a].

Incidence [MS06]. Incidences [Sit06]. Incongruent [SU13]. Incremental [SI94, THI99]. Independent [BDJ10, GJ21]. Index [Ano97, Ano98, Ano99, Ano00, Ano01, Ano02, Ano03, Ano04, Ano05, Ano06, Ano07, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13a, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, Ano23, Van91]. Indicators [Ber00]. Induced [EHP18]. Inducing [SS11]. Inequality [AMV13]. Infimaximal [MS03]. Inflating [BG11a]. Information [CK97a, DDE^{+}07, Wis00]. Inner [MHN06]. Input [AKS^{+}12, EHP19]. Inscribed [Gav09a]. Instance [Tou05]. Instance-Based [Tou05]. Instances [BBR09]. Integer [DD00, KNA94]. Intended [VB05]. Intensity [CHW^{+}08, WDBB09]. Intensity-Modulated [CHW^{+}08, WDBB09]. Interconnection [LYW97]. Interiors [Bin02]. Interpolation [Hiy08, Pro22]. Intersecting [HS02, MGR09]. Intersection [BDJ10, BDGT13, CFM^{+}01, DG99, FK18, GJS09, HYSC18, JMMR23, MC91, dFdSdF17]. Intersections [CGG^{+}12, ZE02]. Interval [LWZ17, MF06, Mit00]. Interval-Based [MF06]. Intervals [dBLM^{+}19]. Intervisibility [MVV07]. Invariant [KS11, LD15]. Invariants [GW04]. Inverse [ZWG06]. Involving [BM12]. Irregularities [Guh05]. ISODATA [MMNM07]. Isomorphic [KU99]. Isomorphism [ÓWW00]. Isosceles [BMSS11]. Isothetic [MGR09]. Iteration [BM12].

Joint [Guh05].

Keep [NŻ20]. Kinematics [ZWG06]. Kinetic [KSS02, dBLM^{+}19]. Kinodynamic [RS11]. Kleitman [Zer12]. Krieger [Zer12].

Label [Gav09a, KT03, ZP01]. Labeling [BG14, CLL05, DMM02, KSY^{+}01, KNN^{+}02, SW01, WTX02, ZP01]. Labels [KSY^{+}01]. Laguerre [HYSC18]. Laguerre-Intersection [HYSC18]. Laplace [Xu06]. Large [DKMM23]. Largest [BCD^{+}00, DBHM^{+}03, FSS^{+}97, KS13, MR03, YCCV17]. Lattice [Lab08, Poo09]. Layer [LYW97]. Layered [FM01, Sud04, WCLS07]. Layout [EvKSS15]. Leaf [CHL^{+}04, CHL^{+}06]. Leapfrog [ABC^{+}15]. Learning [NN09, Tou05]. Least [GSa20]. Leaving [KL10a]. Lebesgue [BS05]. Lemmas [AGR16]. Length [BFL21, JJ10, Ngu12]. Level [AGM^{+}12, BYM^{+}18, FN05]. Library [Gui22]. Like [BS12, MS03]. Likely [CCMS19, SV16]. Line [ACGK17, BS12, BBB^{+}19, BMT99, CLL05, CW12b, CFM^{+}01, DK99, GR10, GR03a, JW24, KMW00, LHHHP03, MGR09, PD13, SS11, WLW01, WZ16, Wis00, CL93]. Line-Constrained [WZ16]. Line-of-Sight [BBB^{+}19]. Line-Segment [PD13]. Linear [AGM^{+}12, LWŻ12, WZ20, dFdSdF17]. Linear-Time [LWŻ12, WZ20, dFdSdF17]. Lines [CDKW05, DL06, LHHHP03, MS03]. Link [ADS00, CT97]. Linked [DMMH11]. List [DMMH11]. Lists [DG99]. Lithographic [SPPK08]. Local [BDM^{+}20, GIPR12, RW11]. Locally [KG14]. Locating [AMP10, AFN11, CW12b]. Location [BMKS00, BKST00, CL17, DG98, DK06, IM12, CT92]. Locations [KZ10]. Locus [TO21]. Logarithmic [KS99]. Long [GMMW19]. Longest [KLV21]. Longest-Edge [KLV21]. Look [MS03]. Low [LW04, dF18, dBGR19]. Low-Dimensional [dBGR19]. Lower [Afs13, AHM^{+}06, BS05, BHLL10, BK17, DG16, KPS13, LOS01]. Luggage [AHP08].

Machine [Afs13, LDHX20, NN09]. Machining [WIEH05]. Maintaining [DDE^{+}07, Jan93]. Maintenance [Sha97a, Sha97b]. Make [KZ10]. Making [MS03]. Manhattan [GSZ11]. Manifolds [CLR07, Dey97, DMMH11]. Manipulation [MST13]. Manufacturing [FM01]. Many [CM10]. Map [CLR07, CLRW10, EvKSS15, KSY^{+}01]. Mapping [RS07]. Mappings [CMO03]. Maps [BCHS07, BS16, BS17, SV15]. Marginal [DLOP06]. Maskless [SPPK08]. Matability [BS08]. Matching [AAR97, BK17, CARB15, CHW^{+}08, EvKSS15, JH04b, Sch16]. Matchings [BHP16]. Matrices [CKMK03]. Matrix [WDBB09]. Matroid [JJ06]. MAX [Wan15]. Max-Volume [EHP18]. Maximal [AFN11]. Maximally [GHH^{+}98]. Maximize [MGD15]. Maximizing [BRD09, CDG^{+}09, DKS05]. Maximum [BDJ10, EH19, Gav09a, JMMR23, KLV21, LWZ17, Mit97, WNGK^{+}12]. Maze [KL10a]. Meaningful [DG03]. Means [FS08, HH08, WZ16]. Measure [Ror19, Wil15]. Measured [FOG00]. Measures [DFL^{+}18, GM99]. Measuring [BYM^{+}18]. Mechanical [FPNZ98, JMM98]. Medial [EMM98, GRS08, SFM07]. Median [WZ16]. Medical [WCLS07]. Melzak [Wen02]. Memory [CFM^{+}01, Nek13]. Mesh [AGL09, Ber00, CMO03, FOG00, Sch00, TW00]. Meshes [AM07, Ber00, BBCK05, JH04a, JH04b, MHW00, RSS^{+}05]. Meshing [BE00, CDRR05, MH00, SBBC00]. Method [BMT00, CCD06, Goo98, HYSC18, San09, VB05, CT92]. Methods [ESG98, FPNZ98, LHHHP03, Tou05]. Metric [ACC^{+}12, AHP08, ETT08, KN20, SPPK08, WZ20, Wil15]. Milling [ACM01]. Min [AAK^{+}06, BHP01, EHP18]. Min- [EHP18]. Min-/Max-Volume [EHP18]. Min-Hausdorff-Distance [BHP01]. Min-Sum [AAK^{+}06]. Minimal [BMKS00, DEG^{+}03, GC97, GBRT13]. Minimization [HSKK98]. Minimizing [AACKM11, LWZ17]. Minimum [ACGK17, AGLN03, BFMFP^{+}14, BBL08, BDE02, CDJ^{+}15, Cha02, CL13, CT97, Col04, ECHS11, EC15, Fra08, GKK^{+}10, GSZ11, Jia15, KKY00, MS99, MGR09, TWC06, WLW01]. Minimum-Bends [ECHS11]. Minimum-Dilation [GKK^{+}10]. Minimum-Width [Cha02]. Mining [Tou05]. Minkowski [BBR09, LLCC11, MS07b, MS10]. Mitered [WJA20]. Mixed [RS99]. MMP [TO21]. Mobile [DK06, DK08, GR10, JW24]. Model [GMV99, LYW97]. Modeling [MG98, SPP08, TW06]. Modelling [SOR06]. Models [AMM^{+}98, BBCS99, Goo98]. Modem [DHT15]. Moderate [BL03, CKMK03]. Moderate-Sized [CKMK03]. Modular [RS11]. Modulated [CHW^{+}08, WDBB09]. Molecular [ZWG06]. Monochromatic [DP02]. Monotone [AC01, BBB^{+}19, EW00]. Monotonic [MS07b]. Monotonicity [BV05]. Morphing [Ber05, Bes02]. Morphological [WR07]. Most [AHO^{+}14, CCMS19, SV16]. Motion [CDG^{+}09, Cho99, GR10, HL97, KS10, RS11]. Motorcycle [HH12]. Mountain [CHW^{+}08]. Moving [AGMR98, BDIZ03, DG98, DDE^{+}07, LWZ17, WG21, vKLSW18]. Multi [AACKM11, FN05, FOX08, GR03b, SM06, WIEH05]. Multi-Axis [WIEH05]. Multi-Directional [FOX08]. Multi-Group [SM06]. Multi-Level [FN05]. Multi-Particle [GR03b]. Multicriteria [GGBK21]. Multidimensional [CFL15, EGS08, KS10, Van91]. Multiple [ACM01, HL98, HLM^{+}14]. Multiple-Robot [HL98]. Multiple-Tool [ACM01]. Multisearching [Ata99]. Mutual [ABR14].

n [HDY07]. Natural [Hiy08]. Navigating [CCJV17, CL93]. Navigation [ACFV10]. NC [WIEH05]. NC-Machining [WIEH05]. Near [AMV13]. Nearest [BD05, CVY11, KS11, SV16, Tou05, Wan15, WZ19]. Nearly [BNS10]. Nearly-Opposite [BNS10]. Necklace [SV15]. Necklaces [Ber04]. Neighbor [AM07, CVY11, Hiy08, KS11, SV16, Tou05, Wan15]. Neighborhoods [EFS09]. Neighbors [AMV13, WZ19]. Net [WCLS07]. Network [BC06, MH00]. Networks [AACT17, CLLP09, GSZ11]. News [VR04]. Nice [AH11]. No [kLkHsL^{+}23]. No-Back-Edge-Traversal [kLkHsL^{+}23]. Noisy [ACC^{+}12, MNG04, ULVH10]. Non [GJS09, Kei97, MTT99, Pap99, Pro22, SOR06, Sto21, TSN97, Tat23, WQS05]. Non-Cartesian [SOR06]. Non-Conjugacy [Sto21]. Non-Convex [Tat23]. Non-Crossing [Pap99, TSN97]. Non-Intersection [GJS09]. Non-Piercing [Kei97]. Non-Rational [Pro22]. Non-Uniform [MTT99, WQS05]. Nonintersecting [AC01]. Nonobtuse [Epp97]. Nonparametric [DLMS13]. Nonpositive [Maf14]. Nonrectangular [AB09]. Nonsmooth [Cho99]. Norm [EH19]. Normal [CLR07]. Normal-Compatible [CLR07]. Normal-Map [CLR07]. Normals [MNG04]. Normed [WNGK^{+}12]. Note [FMHT14]. Notice [Lee03]. NP [BG11a, BZ14, GKK^{+}10, Roy16]. NP-Hard [BG11a, BZ14, GKK^{+}10, Roy16]. Null [Dey97]. Null-Homologous [Dey97]. Number [AH19, Col04, EC15, FMHT14, KKY00, KU99, MGD15, MS99, dBHOvK97]. Numerical [For95]. NURBS [BXHN03].

O [Afs13]. Object [CCMS19, DGN09, GMV99]. Objects [AS08b, AS18, APS00, AGR16, BSC00, CW12b, NY98, PL04]. Obnoxious [BMKS00, CW12b]. Obstacle [CT97]. Obstacles [AC01, BL03, CCK^{+}06, KSN99, LYW97]. Obtaining [dFdSdF17]. Obtuse [FMHT14].

Octilinear [MHS07]. Octree [Sch00]. Octree-Based [Sch00]. Offsets [WJA20]. Omino [AH19]. On-Line [GR10, LHHHP03, CL93]. One [LPC00, LHX24]. One-Pass [LHX24]. Onion [BS12]. Online [BBC^{+}02, BDDT17, KS10, LHX24, Jan93]. Onto [RS07]. Operations [HV91, JJ10]. Operator [Xu06]. Operators [SBBC00]. Opposite [BNS10, GBRT13]. Optimal [AFK^{+}10, AKM^{+}17, AAF10, BKC09, BD05, BKST00, Bes02, BG05, CHW02, CM10, CT97, DMS10, DK12, DK99, GC97, HDY07, KG14, KK10, NZ06, RR00, RS11, SV01, Tan02, Wu09, WDBB09, Xu06]. Optimal-Ratio [Wu09]. Optimality [IM12]. Optimally [WZ20]. Optimization [ACKT01, CS06, GR03b, HQYD22, KTT02, LD15]. Optimization-Based [ACKT01]. Optimized [GGBK21]. Oracle [EFKP13]. Oracle-Based [EFKP13]. Order [ABG^{+}09, ACK^{+}16, AKM^{+}17, BMvR16, GR03a]. Order-Preserving [GR03a]. Ordered [GR03a]. Orderings [ACK^{+}16, AKM^{+}17]. Orientation [BZ14]. Orientations [GBRT13]. Oriented [MR03, SI94]. Origin [EEM11]. Orthogonal [AECSU98, BI21, BMT00, BHLO11, BG11b, CY17, KM11, Kei97, MJ12, Nek13, SU13, SM00, WK07, WDBB09]. Orthostacks [DIL10]. Other [CFM^{+}01, Fra08, dFdSdF17]. Outer [DE12]. Outer- [DE12]. Outerplanar [DL07]. Outliers [CW12a, Da 11, HQYD22]. Output [EFKP13, KMW00, NY98]. Output-Sensitive [EFKP13, NY98]. Overlap [CDG^{+}09]. Overlaying [JH04a, JH04b].

p4 [AH19]. p4-Tilings [AH19]. Packed [BK17]. Packing [AS18, BE00, BHLL10, EFK^{+}07, Epp97, SYI00, TWC06]. Pair [KSN99]. Pairs [Pap99]. Pairwise [BK07, WCMS04]. Parabola [CEK^{+}07].

Parallel [BET99, CDKW05, Che98, CWKC98, MGD15, SPP08, Seg99, STÜ07, Zhu97]. Parallelization [CR01]. Parallelizing [TMPD95, TMPD97]. Parameter [BBL08]. Parameterizations [GW04, WJG97]. Parametric [BD05, KTT02, SYI00]. Pareto [CNTV10]. Part [JH04a, JH04b]. Partial [BMSN19, BFL21, Sch16]. Particle [GR03b]. Partition [AGLN03, CM11, MS99]. Partitioning [DP02, GHH^{+}98, NPR17, VR04, Van91]. Partitions [DK12, DD00, DKK09, KK05, MS14]. Parts [DP02]. Pass [LHX24]. patches [BXHN03]. Path [CCK^{+}06, CDK01, CM10, CT97, DN18, GGBK21, HLM99, HZX24, WIEH05]. Paths [ACH^{+}12, AL11, AC01, BL03, CM10, CJVW12, CSY97, DL07, DR20, GBRT13, KSN99, LYW97, Pap99, TSN97, WZ20, WNGK^{+}12]. Pathwidth [Sud04]. Patterns [BBG^{+}11]. Peeling [CLX03]. Permutations [ABD^{+}11]. Persistent [ACW22, CFL15, DG99]. Perspectives [TW00]. Perturbation [HL04]. Perturbations [BDG14]. Phase [BMT00]. Pieces [MS99]. Piecewise [GOG11, HREK07]. Piercing [AK99, DDCN13, Kei97, Seg99]. Placement [BRD09, Gav09a, KT03, MMG01]. Places [BDGW10]. Placing [MGD15]. Planar [AB09, ADF13, ABR14, AAK^{+}06, BS12, BHP16, BCHS07, CW12b, CL17, CK97b, DDE^{+}07, DN18, FW03, Fra08, GR03a, DDL^{+}10, GKS99, HREK07, HL97, IM12, LW04, Maf14, MST13, NY98, CT92, FMR05, Jan93]. Plane [ADD^{+}13, BC06, BDP08, BV13, BK18, CER97, DK12, DE12, DG16, EBGK^{+}07, GJS03, GBRT13, KKY00, KU10, KSN99, LHHHP03, MHN06, SJ99, TSN97, Wan15, WZ16, WZ18, WZ19, Yan23, YLL18]. Planes [Rab05, Vig12]. Planning [Cho99, GR10, GGBK21, HL97, HLM99, KS10, RS11]. Plans [DG13]. Point [AAR97, ACK^{+}16, AKM^{+}17, AGM^{+}12, BD05, BV13, BK07, BK02, BS00, BG14, CHU14, CW12a, CL17, CGG^{+}12, DEH^{+}05, DG98, DK99, DDE^{+}07, DMM02, DP02, DKMM23, EBGK^{+}07, EGS08, EvKSS15, FMHT14, Gav09b, DDL^{+}10, GKS99, Gui22, GJSD97, IM12, JS09, Kan97b, KS13, KBA11, Kir07, LS08, MB02, MNG04, MMG01, MJ12, PS19, Ror19, Roy16, Sit06, ULVH10, Wis00, WTX02, YCCV17, ZP01, CT92, Jan93]. Point-Dominance [GJSD97]. Point-Placement [MMG01]. Point-Set [BV13, DDL^{+}10, Jan93]. Pointed [AHO^{+}14]. Pointer [Afs13]. Pointerless [AM07]. Points [AGMR98, AAK^{+}06, AAF10, BDIZ03, BDBF^{+}14, Bes03, BI22, BM12, BK18, BDGT13, CDKW05, CDWK01, CLL05, Col04, Da 11, DP03, DGRS08, DK06, EC15, FKNN17, GJS03, JW24, Jia15, KK05, KU10, KNN^{+}02, KU99, Löf11, MGD15, SJ99, SW01, WZ18, Zhu04a, dBLM^{+}19, KNA94]. Pointsets [MRM15]. Poisson [DN18]. Poly [ABD^{+}18]. Polycubes [AB09]. Polygon [AACKM11, AHK^{+}14, BRD09, BHP01, BMT99, BVL11, BHLM03, BNS10, CK97a, Che98, CHW02, CvO01, HL98, MGR09, Pap99, SPPK08, VR04, WK07, KNA94]. Polygonal [AFK^{+}10, ABC^{+}15, AC01, BBB^{+}10, CD03, CT97, CGJS11, CMO03, DEH^{+}05, DLMS13, EHP18, HH08, LSS02, LPC00, MS99, PL04, SVY16, STYK01]. Polygones [Gui22]. Polygons [AECSU98, AT18, ABD^{+}11, AMP10, AFN11, BS08, BBB^{+}19, BG05, BHLO11, BV05, CVG^{+}07, CNTV10, CT97, DL06, DH13, DR20, HS02, IST20, Kei97, KS02, KSS02, LR00, MHW00, Nar99, NPR17, Poo09, SS11, Sha01, SM00, Tan99, Tan02, TWC06, THL98, WJA20, Żak10]. Polygons/Trees [Poo09]. Polyhedra [AH11, BHLO11, BG11b, Bin02, BV05, CY17, CDRR05, Guh05, Vig12, Zhu97].

Polyhedral [BSC00, Bar98, GHH^{+}98, GK20, TMPD97, dBHOvK97, TMPD95]. Polyhedron [Res14, TO21, Tat23, WLW01]. Polyline [AAK^{+}06]. Polylines [Ber05, Bes02]. Polymatroid [KTT02]. Polynomial [BGK^{+}09, BL03, KYZ14, SV01]. Polynomial-Time [BGK^{+}09, BL03, KYZ14]. Polynomials [Pro22]. Polyominoes [ABD^{+}18, AB09, Yan23]. Polytopes [CR01, EFKP13, GHH^{+}98]. Popular [BDGW10]. Posed [BBR09]. Position [FKNN17]. Positions [DMM02]. Positively [EHP19]. Possibilities [BHMW11]. Postman [DG98]. Practice [RS99, TW00, FMR05]. Precise [HREK07]. Precision [FR98]. Predicate [MS22]. Predicates [ETT08, MSB19]. Preprocessors [SZP10]. Presence [BvR19]. Preservation [JH04b]. Preserving [AHK^{+}14, GR03a, HHMK14, JJ10, SFM07]. Price [BMvR16]. Pricing [CLLP09]. Primitive [Che10]. Primitives [MST13]. Principal [CWW08]. Printing [GGBK21]. Prisms [KV23]. Probabilistic [BDIZ03, Gui22]. Problem [Als97, AAMT15, BBR09, BGK^{+}09, BV13, BBL08, BGL^{+}97, BS05, BKN^{+}11, BZ14, CARB15, CDJ^{+}15, DFLON12, DDCN13, DBHM^{+}03, DHT15, EFS09, GLL^{+}99, HZX24, LWŻ12, iN23, WKG10, Wen02, WK07, WDBB09, XLYB04, YLL18]. Problems [Afs13, AR19, AHM^{+}06, BMSS11, CS06, Cha12, CDK01, CHW02, CHL^{+}04, CHL^{+}06, CFM^{+}01, DG99, DG98, FLMS18, FSS^{+}97, GR10, GJSD97, HSS05, JS09, KPS13, KK10, Maf14, MNP^{+}00, MJ12, Por09, WZ16, WCLS07, Wu09, ZG06, Dev92]. Product [JPV21, LSB04]. Products [JMM98]. Programming [Bar98, DD00, KNA94]. Projection [ACFV10, HQYD22]. Projections [AH11, BHLO11, EFKP13].

Prone [KL10a]. Properties

[ABG^{+}09, BEW03, BvR19]. Property [CPRS18]. Protein [FOX08]. Provable [CWW08]. Provably [Mit97]. Proximity [HLW13, KL10b, Tou05, dF18, BDH^{+}12]. Pseudo [AHO^{+}14, AAH^{+}15]. Pseudo-Triangulations [AHO^{+}14, AAH^{+}15]. Pseudomanifolds [DMMH11]. Pspace [BG14]. Pspace-Complete [BG14]. Pursuit [ABC^{+}15, BBB^{+}19, GLL^{+}99]. Pursuit-Evasion [ABC^{+}15, GLL^{+}99]. Push [DG13].

Quadrangular [MHW00].

Quadrangulations [PS19]. Quadratic [DNW^{+}09, GW04, WJG97]. Quadrics [WJG97]. Quadrilateral [BE00, MH00, RSS^{+}05]. Quadtrees [BET99, EGS08]. Quality [BET99, Ber00, CDRR05, MH00]. Quantile [MNP^{+}00]. Queries [CEK^{+}07, CVY11, CT97, GJS09, kLkHsL^{+}23]. Query [CDK01, KS13, MMS97]. Query-Sensitive [MMS97].

Radial [ACK^{+}16, AKM^{+}17]. Radiation [CHL^{+}04, CHL^{+}06, CHW^{+}08, WDBB09]. Radius [WZ20]. Random [Gui22, HQYD22]. Randomization [Dev92]. Randomized [CFM^{+}01, MS22]. Range [Afs13, BMSN19, FN05, FPNZ98, KS05, kLkHsL^{+}23, MJ12, Nek13]. Ranges [FN05]. Rank [JJ06]. Ratio [Wu09]. Rational [GW04, HREK07, Pro22, WJG97]. Ratios [BFL21]. Ray [GKS99, Goo98, KYZ14, MMS97]. Rays [DL06]. Reaching [CvO01, Kan97b]. Recognition [Roy16]. Recognizing [BV05, EHP19]. Reconciling [SZP10]. Reconfigurable [RS11]. Reconstructing [ACK^{+}16, AKM^{+}17, BHLO11, EHP19]. Reconstruction [ACC^{+}12, AT18, ACDL02, BBCS99, DW02, DGRS08, FKMW22, Wis00, DEG^{+}03]. Recovery [HQYD22]. Rectangle [FM99, GJSD97, MR03]. Rectangles [Gav09a, Kei97, KBA11, KNN^{+}02, Seg99]. Rectangular [ACS18, DD00, DKK09, MHN06, Por09, Wan09]. Rectilinear [AC01, DMS10, GC97, KSY^{+}01, LYW97, TSN97, WZ18]. Red [AC01, BK18, HSS05]. Red-Blue [HSS05]. Reducing [BBR09]. Reduction [CHW^{+}08, Rab05]. Reference [AAR97]. Refined [vKLSW18]. Refinement [Lab08, MPW05, RW11, STÜ07]. Refinements [MHW00]. Reflector [AAMT15]. Reflex [ACCS04]. Reflex-Free [ACCS04]. Region [DKS05, Gav09a, LSS02, MVV07, STYK01, Wu09]. Regions [ACS18, BK14, CJVW12, CCJV17, NPR17, TSN97, FMR05]. Regular [Guh05]. Related [Afs13, Als97, BMSS11, DG98, FSS^{+}97]. Relations [Wan09]. Removal [GMV99, Lab08]. Reparametrization [SV01]. Reporting [Afs13, CGG^{+}12, Nek13]. Representation [AAH^{+}11, ADM11, JMM98, Kan97a, MG98, DMMH11]. Representations [BBCK05, Sha97a, Sha97b, Sha99]. Representatives [Sto21]. Representing [ALS12]. Resemblance [KC97]. Resilience [KYZ14]. Resolving [Sit06]. Respect [dBLM^{+}19]. Restricted [AGL09]. Restrictions [MH00]. Resultant [EFKP13]. Results [Ben23, KG14]. Reverse [CVY11]. revised [Van91]. Revisited [CDJ^{+}15, DGN09, GJSD97, PX15]. Revolution [WG21]. Right [DE12]. Rigid [CDG^{+}09, Sit06]. Rigidity [JJ06, JJ10, Ngu12, OP10]. Rings [Seg99]. River [Sug92]. Robot [ACFV10, GR10, HL97, HL98, Kan97b, KS10]. Robots [DG13, RS11]. Robust [DLMS13, LDHX20, MNP^{+}00, Müc98, SI94].

Roofs [EHP18]. Room

[KZ10, LPC00, PLC02]. Rooted [KK05]. Rotating [BDBF^{+}14, Vig12]. Rotational [BSC00]. Round [DR02]. Rounding [GM98]. Roundness [DR02, DP03, San09, SJ99]. Routable [NPR17]. Route [WKG10]. Routes [THI99]. Routing [BBC^{+}02, BDDT17, BDM^{+}20, HL97]. Ruled [Aj24, WG21]. Ruler [GSS23]. Rules [HS02].

Sabin [WQS05]. Sail [NS09]. Salesman [EFS09, XLYB04]. Sampled [BYM^{+}18]. Sampling [CFL15, DGRS08, FIS08]. Saw [DH13]. Scalar [BYM^{+}18]. Scale [ULVH10]. Scallop [BBB^{+}19]. Scans [BBCS99]. Scattered [CCJV17]. Scenes [dBHOvK97]. Scheduling [OGB11]. Schemes [MG98]. Sculptured [KMG^{+}01]. Search [FN05, KS05, KS11]. Searchable [KZ10]. Searcher [LPC00]. Searching [BMSN19, FPNZ98, LSS02, LPC00, PLC02, SV16, STYK01, Vig12, Wan15]. Searchlight [OGB11]. Sections [EW00, GW04]. Seen [CCMS19]. Segment [ADS00, ACGK17, BHP01, BMT99, CGG^{+}12, CFM^{+}01, PD13, Wis00]. Segmentation [ACKT01, CWW02, WCLS07]. Segments [AAF10, Bes03, BCD^{+}00, DG99, DK12, KMW00, KS99, MS03, MGR09, PL01, WLW01, XLYB04, XYZK10, Zhu04a]. Seismic [Gui22]. Selected [CP05]. Selecting [Cha01]. Selection [AGR16, LLCC11, ULVH10]. Self [RS11]. Self-Reconfigurable [RS11]. Semi [Ben23, KK05, MS07a]. Semi-Algebraic [MS07a]. Semi-Balanced [KK05]. Semi-Symmetric [Ben23]. Sensing [GIPR12]. Sensitive [EFKP13, KMW00, MMS97, NY98]. Sensors [JW24, KYZ14]. Sentinel [LS08]. Separability [AHM^{+}06, AGM^{+}12, HSS05].

Separating [BCD^{+}00, CDKW05, CER97, DEH^{+}05]. Separation [CEK^{+}07, Guh05]. Separator [FOX08]. Sequences [GM99]. Sequencing [CHL^{+}04, CHL^{+}06]. Service [BMKS00, BGT99]. Set [ACK^{+}16, AKM^{+}17, AEK05, BYM^{+}18, BV13, CDJ^{+}15, CW12a, Col04, DDCN13, DR02, DP03, DMM02, DK06, DKMM23, EvKSS15, Gav09a, DDL^{+}10, GKS99, KBA11, MB02, MGR09, MJ12, Sha99, SJ99, WLW01, Yan23, DEG^{+}03, Jan93]. Sets [AGM^{+}12, BHP01, BDJ10, BCD^{+}00, BK02, CHU14, CGG^{+}12, DEH^{+}05, DK99, DDE^{+}07, DP02, DKMM23, EBGK^{+}07, ESS11, EGS08, FMHT14, KK05, KU10, Kir07, Kra20, PS19, Ror19, Seg99, YCCV17]. SFCDecomp [GGBK21]. Shallow [AS08a]. Shape [CC06, CSU99, MST13]. Shaped [DG13]. Shapes [AAR97, KNN^{+}02]. Sharp [DW02]. Shifting [dFdSdF17]. Shooting [Goo98, MMS97]. Shortcuts [DN18]. Shortest [ACH^{+}12, AL11, ADS00, BMT99, BL03, CCK^{+}06, CJVW12, CT97, CSY97, DR20, KS99, KSN99, Pap99, TSN97, THI99]. Shuffling [DG01]. Sight [BBB^{+}19]. Signed [ABD^{+}11]. Signs [CKMK03]. Similarity [BBR09, Kir07, Sch16, SVY16]. Simple [AT18, ACDL02, BMT99, BG05, BVL11, CK97a, CNTV10, CT97, IST20, KS02, KSS02, Nar99, NPR17, Pap99, THL98, VR04, WTX02, Dev92]. Simplex [Afs13]. Simplices [CHU14, EEM11]. Simplicial [AM07, ALS12, BBCK05, CW12a, EW00, FOG00, LSB04]. Simplification [AHK^{+}14, AGL09, CGJS11, CMO03, DLMS13, HH08, SFM07, WR07]. Simplifying [ALS12]. Simplipoly [CGJS11]. Simultaneous [ADF13, DL07, GHN^{+}03]. Single [CLL05, CL13, Tat23]. Single-Source [CL13, Tat23]. Site [MMR01]. Sites [HDY07, VO98]. Size [BFMFP^{+}14, ELPZ07, RSS^{+}05, RW11, Wil15]. Sized [CKMK03].

Skeleton [HH12]. Skeletons

[BHP16, EHP19, WJA20]. Skew [AAC^{+}99]. Skinny [CCJV17]. Skip [EGS08]. Sliding [BDP08, CS06, KM11, KSY^{+}01]. Sliver [Lab08]. Slopes [DG03]. Small [AKKS14, CDRR05, EBGK^{+}07, KU99, Mit97, NŻ20]. Smallest [AS18, Cha02, DGN09, FG04, NN09]. Smooth [CP05, GOG11]. Smoothing [GLS10, HH08]. Smoothness [CWW02]. Software [ZE02]. Solid [Goo98, SPP08]. Solids [KMG^{+}01, Sha99]. Solution [Ber00, FOG00, Gav09b, VB05]. Solution-Based [Ber00]. Solutions [DD00, HV91, KK10]. Solvation [HYSC18]. Solving [Yan06]. Some [AHM^{+}06, Ben23]. Sorting [Che10]. Source [CL13, Tat23]. Space [BS12, CD03, CSY97, DK12, GGBK21, LHX24, MSB19, Sha97a, Sha97b, Van91, WZ20, WNGK^{+}12]. Space-Efficient [CD03, LHX24]. Space-Filling [GGBK21]. Spaces [Ben23, ES97, HLM99, KN20, Wil15, dBGR19]. Spanner [LW04, XYZK10]. Spanners [BI22, BSX09, BDD^{+}12, DN97, DG16]. Spanning [AACT17, AR19, AGLN03, BvR19, CL13, DGN09, RS99, WLW01]. Sparse [DN97, dBHOvK97]. Spatial [Yan06]. Special [BV05]. Specification [SOR06]. Specified [DMM02, FR98]. Specified-Precision [FR98]. Sphere [AS01, RS07, Xu06]. Spheres [Gav09b]. Spherical [HS02, KS10, Xu06]. Spiralling [KMW00]. splines [BXHN03]. Spread [Wil15]. Square [BFMFP^{+}14, BHMW11, Kan97b]. Square-Tiling [BHMW11]. Squares [MGD15]. Stabbing [CHU14]. Stability [BDG13, BDG14, DK06, For95]. Stable [EMM98, Hiy08]. Stabs [KMW00]. Stage [EFKM08]. Star [LWŻ12]. Static [CHL^{+}04, DBGV06, IM12]. Statistical [MNP^{+}00]. Steiner [AAF10, BZ14, DK06, GC97, KU99, MHS07, Wen02]. Step [CW12a]. stereolithography [FM97]. Stoker [BG11b]. Stone [CEK^{+}07]. Storage [BM02]. Straight [BHP16, EHP19, GR03a, HH12, Tan99]. Straight-Line [GR03a]. Straight-Skeleton [HH12]. Strange [MF06]. Strategy [ABC^{+}15, MMG01]. Streaming [AKKS14, Che10]. Streams [FIS08]. Streets [LOS01]. Stretch [WNGK^{+}12]. Strong [ACFV10, FW03]. Strongly [AACT17, BG14, CDWK01]. Structural [ABG^{+}09, AAH^{+}11]. Structure [ALS12, FG04]. Structures [CW12a, EGS08, KL10b, LSB04]. Study [DGL^{+}00, LHHHP03, TV01]. Sub [BYM^{+}18]. Sub-Level [BYM^{+}18]. Subdivision [BM02, BK18, LD15, ZWG06]. Subdivisions [BBC^{+}02, BDM^{+}20, CL17, KU10]. Subgraph [AACT17]. Subgraphs [JMMR23]. Sublinear [AMV13]. Subsets [DKMM23]. Subspaces [FKMW22]. Subtrajectories [BBG^{+}11]. Successive [CMO03]. Sum [AACKM11, Als97, AAK^{+}06, BBR09, DLOP06, LLCC11, MS10, WZ19, BHP01]. SUM-Hard [BHP01]. Sums [KV23, MS07b]. Superhull [CDWK01]. Superimposing [CC06]. Support [LDHX20]. Surface [Aj24, ACDL02, CM11, DGRS08, GMV99, HV91, JH04a, JH04b, MC91, MNG04, MH00, WCLS07]. Surfaces [CP05, CLRW10, GOG11, Rab05, RS07, SYI00, WQS05, WG21]. Surveillance [BDBF^{+}14]. Survey [JTNM06]. SVM [LHX24]. Sweep [BSC00]. Swept [BSC99, BSC00]. Symmetric [AACT17, Ben23]. Symmetry [GJS03, OP10]. System [GR03b, KMG^{+}01, SM06, VB05, Yan06]. Systems [JTNM06, Sit06, SZP10, TW06].

Tangent [Rab05]. Technique [MS03]. Techniques [CR01, FOG00, MH00].

Template [MH00]. Ten [Yan23]. Terrain [AEK05, DG03, FM01, HLM^{+}14, TMPD97, FMR05, TMPD95]. Terrains [GLS10, MVV07]. Tessellation [BS12, DNW^{+}09]. Testing [ABR14, BK02, CPRS18]. Tethered [HL97]. Tethered-Robot [HL97]. Tetrahedra [CCJV17, LD15]. Tetrahedral [Ber00]. Tetrahedralization [MMG01]. Tetrahedralizations [GOG11]. Tetris [BDH^{+}04]. Their [AT18, BK07, CEK^{+}07, EHP19, Kra20, Cho99]. Theorem [BG11b, Zer12]. Theorems [CPRS18]. Theoretic [ABG^{+}09]. Theory [RS99, TW00]. Therapy [CHL^{+}04, CHL^{+}06, CHW^{+}08, WDBB09]. Thickness [CW12a]. Things [NŻ20]. Three [AS18, BSC00, BMT00, BBCK05, Cha12, DB92, EEM11, HDY07, Kir07, Müc98, SU13]. Three-Dimensional [AS18, BSC00, Kir07, Müc98]. Three-Phase [BMT00]. Throwing [CEK^{+}07]. Tightening [WR07]. Tiles [Nab21]. Tiling [BHMW11, Yan23]. Tilings [AH19, Wan09]. Time [AMV13, BBR09, BGK^{+}09, BL03, KMW00, KS02, KS99, KYZ14, LWŻ12, WZ20, dFdSdF17]. Tolerance [DMOW98, HH08]. Tolerant [BI22, MS14]. Tool [ACM01, GGBK21, LSS98]. Top [Wan15, WZ19]. Top- [Wan15, WZ19]. Topics [CP05]. Topological [APS00, BSC00, CCD06, CLX03, ES97]. Topology [BYM^{+}18, CP05, DNW^{+}09, HHMK14, JH04b, SBBC00, SI94]. Topology-Guided [DNW^{+}09]. Topology-Oriented [SI94]. Topology-Preserving [HHMK14]. Tours [ECHS11, EC15, Löf11]. Tracing [MF06]. Transfinite [Pro22]. Transform [EMM98]. Transformation [GK20, IM12]. Translates [CER97]. Translation [BGT99, KS11]. Translational [BSC99, BHP01]. Translations [KC97]. Transportation [BC06, CLLP09]. trapezoid [CT92]. Traveling [EFS09, XLYB04]. Traversal [BM02, kLkHsL^{+}23]. Tree [AR19, BZ14, Goo98, Sha01, Van91, Wen02]. Trees [AGLN03, AMM^{+}98, AGM^{+}12, BFL21, CM10, CL13, EHP19, Fra08, GC97, GR03a, HLW13, MHS07, Poo09, RS99, Sud04, WNGK^{+}12]. Triangle [AMV13]. Triangles [AK99, BMSS11]. Triangular [Ber00, Rab05]. Triangulated [NPR17]. Triangulating [ES97]. Triangulation [ACH^{+}12, BBL08, BDE02, BS16, BS17, DN18, Epp97, GJ21, HSKK98, Mit97, NZ06, SYI00]. Triangulations [AHO^{+}14, AAH^{+}15, ADM11, AAF10, BET99, BDG13, BSX09, BDDT17, Dev02, DB92, ESS11, GHN^{+}03, IMTI02, KU99, Müc98, Nar99, Xu06, For95]. Truchet [Nab21]. Truck [EFK^{+}07]. TSP [DLOP06]. Tubularity [CWW02]. Turns [Col04, Jia15]. Tverberg [MS14]. Twist [EFK^{+}07]. Two [Als97, BG05, BBCK05, BCD^{+}00, BNS10, CDG^{+}09, CD03, CT97, EEM11, KS05, KK05, KU10, KBA11, Kra20, LYW97, MGD15, MS10, PLC02, Tan02, TWC06, THL98, Wan09, WTX02, ZP01]. Two-Circle [WTX02]. Two-Dimensional [CD03, KS05]. Two-Guard [THL98]. Two-Label [ZP01]. Two-Layer [LYW97]. Type [CPRS18]. Types [ACK^{+}16, AKM^{+}17, Wan09].

Unanchored [Kan97b]. Uncertain [JS09, WZ18]. Uncertainties [GJ21, MJ12]. Uncertainty [CvO01]. Under-Constrained [TW06, ZG06]. Unfolding [CY17, DIL10, Poo09]. Unguarded [Bin02]. Unified [BMT00, KT03]. Uniform [BZ14, MTT99, WQS05]. Uniformly [Gui22]. Unions [CDG^{+}09]. Unique [ACS18]. Unistable [Res14]. Unit [CDJ^{+}15, DFLON12, DDCN13, GSa20, dFdSdF17]. Universal [BS05, FLMS18, KPS13].

Unknown [KL10a, CL93]. Unstable [GRS08]. Unstructured [TW00]. Updates [DG99, Nek13]. Upper [DHT15]. Upward [Fra08]. Using [AGL09, BFS01, CWW08, FS08, GW04, GHH^{+}98, HSKK98, KL10a, MST13, MS22, Pro22, SPP08, GGBK21].

Values [DLOP06]. Variable [Gui22].

Variant [DDCN13]. Various

[AGR16, BKST00, KNN^{+}02]. Vector [HHMK14, LDHX20]. Velocity [DK08]. Verification [WIEH05]. Versions [DBGV06]. Vertex [BI21, BI22, DIL10]. Vertex-Unfolding [DIL10]. Vertices [Gav09b, Rab05]. Via [BDG14, Goo98, DD00, kLkHsL^{+}23, Sch16, SYI00]. View [WKG10]. Viewpoint [DDE^{+}07]. Viewpoints [HLM^{+}14]. Views [dBHOvK97]. Violations [DR20]. Virtual [PW01, SBBC00]. Visibility [AMP10, BRD09, BS00, CK97a, Che98, DDE^{+}07, ELPZ07, FM99, FW03, GLL^{+}99, HLM^{+}14, IST20, Kan97a, KMW00, Roy16, SM00, TMPD97, Wis00, TMPD95]. Visibility-Based [GLL^{+}99]. Visible [BMT99]. Visual [Pet98]. Visualization [LSS98]. VLSI [PL01]. Volume [Ano98, Ano03, Ano04, Ano05, Ano06, Ano07, Ano08, Ano09, Ano10, Ano11, Ano12, Ano13a, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, Ano23, EHP18]. Volumes [BSC99, BSC00, GOG11]. Voronoi [AAC^{+}99, AGMR98, BC06, BKC09, BS12, BBB^{+}10, BK14, BKL17, CC06, DKS05, DG98, DN18, DBGV06, EH19, ETT08, Gav09b, GJ21, GSW08, HREK07, HDY07, HH08, KS05, KKS05, MMR01, NS09, PL01, PL04, PD13, PX15, SPPK08, Sug92, SI94, SV16, VO98]. Voronoi-Based [HH08]. Voronoi-Like [BS12].

Walkability [THL98]. Walkable [Tan99]. Walking [DN18]. Watchman [THI99, WKG10]. Watermarking [HHMK14]. Wavefront [TO21]. Weak [Che98]. Weakly [AT18, BMT99]. Web [AL01, BGT99]. Web-Based [AL01]. Weight [BBL08, BDE02, BK18]. Weight-Equitable [BK18]. Weighted [BI22, BHP16, CJVW12, CCJV17, EHP19, EH19, GJS03, LW04, WZ19]. Well [Sha99, Sit06, ZG06]. Well-Constrained [ZG06]. Well-Formed [Sha99, Sit06]. Which [BK07]. Whose [Mit97]. Wide [BGT99]. Width [Cha02, FOX08, SJ99, Jan93]. Width-Bounded [FOX08]. Window [CCMS19]. Windows [CS06]. Wingspans [KS99]. Within [AFN11]. Without [CM10, MH00, BM02]. Witnesses [CJK^{+}06]. Work [MPW05]. Workspace [HL98]. World [BGT99]. Wrapping [GSS23].

Yao [BDD^{+}12, BvR19]. yields [Dev92].

Zone [LHHHP03].

References

Aichholzer:1999:SVD

[AAC^{+}99] O. Aichholzer, F. Aurenhammer, D. Z. Chen, D. T. Lee, and E. Papadopoulou. Skew Voronoi diagrams. International Journal of Computational Geometry and Applications (IJCGA), 9(3):235–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Abu-Affash:2011:MCP

[AACKM11] A. Karim Abu-Affash, Paz Carmi, Matthew J. Katz, and Gila Morgenstern. Multi cover of a polygon minimizing the sum of areas. International Journal of Computational Geometry and Applications (IJCGA), 21(6):685–698, December 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Abu-Affash:2017:SCS

[AACT17] A. Karim Abu-Affash, Paz Carmi, and Anat Parush Tzur. Strongly connected spanning subgraph for almost symmetric networks. International Journal of Computational Geometry and Applications (IJCGA), 27 (3):207–219, September 2017. CODEN IJCAEV. ISSN 0218-1959.

Aronov:2010:OTP

[AAF10] Boris Aronov, Tetsuo Asano, and Stefan Funke. Optimal triangulations of points and segments with Steiner points. International Journal of Computational Geometry and Applications (IJCGA), 20(1):89–104, February 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Aichholzer:2011:CSA

[AAH^{+}11] Oswin Aichholzer, Franz Aurenhammer, Thomas Hackl, Bert Jüttler, Margot Rabl, and Zbynek Šír. Computational and structural advantages of circular boundary representation. International Journal of Computational Geometry and Applications (IJCGA), 21(1):47– 69, February 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Aichholzer:2015:CPT

[AAH^{+}15] Oswin Aichholzer, Franz Aurenhammer, Thomas Hackl, Clemens Huemer, Alexander Pilz, and Birgit Vogtenhuber. 3-colorability of pseudo-triangulations. International Journal of Computational Geometry and Applications (IJCGA), 25(4):283–??, December 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Aronov:2006:PFP

[AAK^{+}06] Boris Aronov, Tetsuo Asano, Naoki Katoh, Kurt Mehlhorn, and Takeshi Tokuyama. Polyline fitting of planar points under min-sum criteria. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3):97–??, June 2006. CODEN IJCAEV. ISSN 0218-1959.

Andre:2015:FFR

[AAMT15] Julien André, Dominique Attali, Quentin Mérigot, and Boris Thibert. Far-field reflector problem under design constraints. International Journal of Computational Geometry and Applications (IJCGA), 25 (2):??, June 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Aichholzer:1997:MSR

[AAR97] Oswin Aichholzer, Helmut Alt, and Günter Rote. Matching shapes with a reference point. International Journal of Computational Geometry and Applications (IJCGA), 7(4):349– ??, August 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Aleksandrowicz:2009:CDD

[AB09] Gadi Aleksandrowicz and Gill Barequet. Counting d-dimensional polycubes and nonrectangular planar polyominoes. International Journal of Computational Geometry and Applications (IJCGA), 19(3):215–229, June 2009. CODEN IJCAEV. ISSN 0218-1959.

Ames:2015:LSP

[ABC^{+}15] Brendan Ames, Andrew Beveridge, Rosalie Carlson, Claire Djang, Volkan Isler, Stephen Ragain, and Maxray Savage. A leapfrog strategy for pursuitevasion in a polygonal environment. International Journal of Computational Geometry and Applications (IJCGA), 25 (2):??, June 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Aloupis:2011:CSP

[ABD^{+}11] Greg Aloupis, Prosenjit Bose, Erik D. Demaine, Stefan Langerman, Henk Meijer, Mark Overmars, and Godfried T. Toussaint. Computing signed permutations of polygons. International Journal of Computational Geometry and Applications (IJCGA), 21(1):87–100, February 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Aichholzer:2018:FPP

[ABD^{+}18] Oswin Aichholzer, Michael Biro, Erik D. Demaine, Martin L. Demaine, David Eppstein, Sándor P. Fekete, Adam Hesterberg, Irina Kostitsyna, and Christiane Schmidt. Folding polyominoes into (poly)cubes. International Journal of Computational Geometry and Applications (IJCGA), 28(3):??, September 2018. CODEN IJ-CAEV. ISSN 0218-1959.

Abellanas:2009:SGT

[ABG^{+}09] Manuel Abellanas, Prosenjit Bose, Jesús García, Ferran Hurtado, Carlos M. Nicolás, and Pedro Ramos. On structural and graph theoretic properties of higher order Delaunay graphs. International Journal of Computational Geometry and Applications (IJCGA), 19(6):595–615, December 2009. CODEN IJCAEV. ISSN 0218-1959.

Angelini:2014:TMD

[ABR14] Patrizio Angelini, Thomas Bläsius, and Ignaz Rutter. Testing mutual duality of planar graphs. International Journal of Computational Geometry and Applications (IJCGA), 24(4):325–??, December 2014. CODEN IJCAEV. ISSN 0218-1959.

Atallah:2001:CRB

[AC01] Mikhail J. Atallah and Danny Z.

Chen. On connecting red and blue rectilinear polygonal obstacles with nonintersecting monotone rectilinear paths. International Journal of Computational Geometry and Applications (IJCGA), 11(4):373–400, 2001. CODEN IJCAEV. ISSN 0218-1959.

Amenta:2008:GEF

[AC08] Nina Amenta and Otfried Cheong. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 18(6): 507–508, December 2008. CO-DEN IJCAEV. ISSN 0218-1959.

Aanjaneya:2012:MGR

[ACC^{+}12] Mridul Aanjaneya, Frederic Chazal, Daniel Chen, Marc Glisse, Leonidas Guibas, and Dmitriy Morozov. Metric graph reconstruction from noisy data. International Journal of Computational Geometry and Applications (IJCGA), 22(4):305– ??, August 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Ahn:2004:RFH

[ACCS04] Hee-Kap Ahn, Siu-Wing Cheng, Otfried Cheong, and Jack Snoeyink. The reflex-free hull. International Journal of Computational Geometry and Applications (IJCGA), 14(6):453–??, December 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Amenta:2002:SAH

[ACDL02] Nina Amenta, Sunghee Choi, Tamal K. Dey, and Naveen Leekha. A simple algorithm for homeomorphic surface reconstruction. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2):125–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Anghinolfi:2010:CPR

[ACFV10] Andrea Anghinolfi, Luca Costa, Massimo Ferri, and Enrico Viarani. A covering projection for robot navigation under strong anisotropy. International Journal of Computational Geometry and Applications (IJCGA), 20(5):511–525, October 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Alt:2017:MCC

[ACGK17] Helmut Alt, Sergio Cabello, Panos Giannopoulos, and Christian Knauer. Minimum cell connection in line segment arrangements. International Journal of Computational Geometry and Applications (IJCGA), 27(3):159–176, September 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Abellanas:2012:ISP

[ACH^{+}12] Manuel Abellanas, Mercè Claverol, Gregorio Hernández, Ferran Hurtado, Vera Sacristán, Maria Saumell, and Rodrigo I. Silveira. Improving shortest paths in the Delaunay triangulation. International Journal of Computational Geometry and Applications (IJCGA), 22(6):559–??, December 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Aichholzer:2016:RPS

[ACK^{+}16] Oswin Aichholzer, Jean Cardinal, Vincent Kusters, Stefan Langerman, and Pavel Valtr. Reconstructing point set order types from radial orderings. International Journal of Computational Geometry and Applications (IJCGA), 26(3–4): 167–184, September/December 2016. CODEN IJCAEV. ISSN 0218-1959.

Asano:2001:EAO

[ACKT01] Tetsuo Asano, Danny Z. Chen, Naoki Katoh, and T. Tokuyama. Efficient algorithms for optimization-based image segmentation. International Journal of Computational Geometry and Applications (IJCGA), 11(2):145–166, 2001. CODEN IJCAEV. ISSN 0218-1959.

Arya:2001:AAM

[ACM01] Sunil Arya, Siu-Wing Cheng, and David M. Mount. Approximation algorithm for multiple-tool milling. International Journal of Computational Geometry and Applications (IJCGA), 11(3):339–372, 2001. CODEN IJCAEV. ISSN 0218-1959.

Aschner:2018:UCR

[ACS18] Rom Aschner, Paz Carmi, and Yael Stein. Unique coverage with rectangular regions. International Journal of Computational Geometry and Applications (IJCGA), 28(4):341– 363, December 2018. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195918500103.

Adams:2022:PHC

[ACW22] Henry Adams, Sophia Coldren, and Sean Willmot. The persistent homology of cyclic graphs. International Journal of Computational Geometry and Applications (IJCGA), 32(1–2):1– 37, March–June 2022. CO-DEN IJCAEV. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195922500042.

Abel:2013:FEP

[ADD^{+}13] Zachary Abel, Erik D. Demaine, Martin L. Demaine, Sarah Eisenstat, Jayson Lynch, Tao B. Schardl, and Isaac Shapiro-Ellowitz. Folding equilateral plane graphs. International Journal of Computational Geometry and Applications (IJCGA), 23(2):75–??, April 2013. CODEN IJCAEV. ISSN 0218-1959.

Angelini:2013:SEE

[ADF13] Patrizio Angelini, Giuseppe Di Battista, and Fabrizio Frati. Simultaneous embedding of embedded planar graphs. International Journal of Computational Geometry and Applications (IJCGA), 23(2):93–??, April 2013. CODEN IJCAEV. ISSN 0218-1959.

Aleardi:2011:CBR

[ADM11] Luca Castelli Aleardi, Olivier Devillers, and Abdelkrim Mebarki. Catalog-based representation of 2D triangulations. International Journal of Computational Geometry and Applications (IJCGA), 21(4):393–402, August 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Aleksandrov:2000:AFS

[ADS00] L. G. Aleksandrov, H. N. Djidjev, and J.-R. Sack. An O(n log n) algorithm for finding a shortest central link segment. International Journal of Computational Geometry and Applications (IJCGA), 10(2):157–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Abello:1998:IOP

[AECSU98] James Abello, Vladimir Estivill-Castro, Thomas Shermer, and Jorge Urrutia. Illumination of orthogonal polygons with orthogonal floodlights. International Journal of Computational Geometry and Applications (IJCGA), 8(1):25–??, February 1998. CODEN IJ-CAEV. ISSN 0218-1959.

Archambault:2005:CSA

[AEK05] Daniel Archambault, Willam Evans, and David Kirkpatrick. Computing the set of all the distant horizons of a terrain. International Journal of Computational Geometry and Applications (IJCGA), 15(6):547–??, December 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Amenta:1998:GEF

[AF98] Nina Amenta and Steven Fortune. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 8(2):117–??, April 1998. CODEN IJCAEV. ISSN 0218-1959.

Ahn:2010:DOE

[AFK^{+}10] Hee-Kap Ahn, Mohammad Farshi, Christian Knauer, Michiel Smid, and Yajun Wang. Dilation-optimal edge deletion in polygonal cycles. International Journal of Computational Geometry and Applications (IJCGA), 20(1):69–87, February 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Aota:2011:AAL

[AFN11] Hirofumi Aota, Takuro Fukunaga, and Hiroshi Nagamochi. An approximation algorithm for locating maximal disks within convex polygons. International Journal of Computational Geometry and Applications (IJCGA), 21(6):661–684, December 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Afshani:2013:IPM

[Afs13] Peyman Afshani. Improved pointer machine and I/O lower bounds for simplex range reporting and related problems. International Journal of Computational Geometry and Applications (IJCGA), 23(4–5):233– ??, 2013. CODEN IJCAEV. ISSN 0218-1959.

Agarwal:1999:GEF

[Aga99] Pankaj K. Agarwal. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):325–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Andersson:2009:RMS

[AGL09] Mattias Andersson, Joachim Gudmundsson, and Christos Levcopoulos. Restricted mesh simplification using edge contractions. International Journal of Computational Geometry and Applications (IJCGA), 19 (3):247–265, June 2009. CO-DEN IJCAEV. ISSN 0218-1959.

Andersson:2003:BPM

[AGLN03] Mattias Andersson, Joachim Gudmundsson, Christos Levcopoulos, and Giri Narasimhan. Balanced partition of minimum spanning trees. International Journal of Computational Geometry and Applications (IJCGA), 13(4):303–??, August 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Arkin:2012:SPS

[AGM^{+}12] Esther M. Arkin, Delia Garijo, Alberto Márquez, Joseph S. B. Mitchell, and Carlos Seara. Separability of point sets by klevel linear classification trees. International Journal of Computational Geometry and Applications (IJCGA), 22(2):143–??, April 2012. CODEN IJCAEV. ISSN 0218-1959.

Albers:1998:VDM

[AGMR98] Gerhard Albers, Leonidas J. Guibas, Joseph S. B. Mitchell, and Thomas Roos. Voronoi diagrams of moving points. International Journal of Computational Geometry and Applications (IJCGA), 8(3):365–??, June 1998. CODEN IJCAEV. ISSN 0218-1959.

Ashok:2016:SLV

[AGR16] Pradeesha Ashok, Sathish Govindarajan, and Ninad Rajgopal. Selection lemmas for various geometric objects. International Journal of Computational Geometry and Applications (IJCGA), 26(2):67–??, June 2016. CODEN IJCAEV. ISSN 0218-1959.

Alam:2011:CNP

[AH11] Md. Ashraful Alam and Masud Hasan. Computing nice projections of convex polyhedra. International Journal of Computational Geometry and Applications (IJCGA), 21(1):71–85, February 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Amano:2019:NPT

[AH19] Kazuyuki Amano and Yoshinobu Haruyama. On the number of p4-tilings by an n-omino. International Journal of Computational Geometry and Applications (IJCGA), 29(1):3– 19, March 2019. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195919400016.

Aichholzer:2014:GPP

[AHK^{+}14] Oswin Aichholzer, Thomas Hackl, Matias Korman, Alexander Pilz, and Birgit Vogtenhuber. Geodesic-preserving polygon simplification. International Journal of Computational Geometry and Applications (IJCGA), 24(4):307–??, December 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Arkin:2006:SLB

[AHM^{+}06] Esther M. Arkin, Ferran Hurtado, Joseph S. B. Mitchell, Carlos Seara, and Steven S. Skiena. Some lower bounds on geometric separability problems. International Journal of Computational Geometry and Applications (IJCGA), 16(1):1– ??, February 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Aichholzer:2014:FCP

[AHO^{+}14] Oswin Aichholzer, Thomas Hackl, David Orden, Alexander Pilz, Maria Saumell, and Birgit Vogtenhuber. Flips in combinatorial pointed pseudotriangulations with face degree at most four. International Journal of Computational Geometry and Applications (IJCGA), 24(3):197–??, September 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Abellanas:2008:HLM

[AHP08] Manuel Abellanas, Ferran Hurtado, and Belén Palop. The heavy luggage metric. International Journal of Computational Geometry and Applications (IJCGA), 18(4):295–306, August 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Al-jedani:2024:CFR

[Aj24] Awatif Al-jedani. On the curvature functions of ruled surface. International Journal of Computational Geometry and Applications (IJCGA), 34(1–2): 25–41, March–June 2024. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S021819592450002X.

Assa:1999:PDB

[AK99] E. Assa and M. J. Katz. 3piercing of d-dimensional boxes and homothetic triangles. International Journal of Computational Geometry and Applications (IJCGA), 9(3):249–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Ahn:2014:CCS

[AKKS14] Hee-Kap Ahn, Hyo-Sil Kim, Sang-Sub Kim, and Wanbin Son. Computing k centers over streaming data for small k. International Journal of Computational Geometry and Applications (IJCGA), 24(2):107–??, June 2014. CODEN IJCAEV. ISSN 0218-1959.

Aichholzer:2017:OAR

[AKM^{+}17] Oswin Aichholzer, Vincent Kusters, Wolfgang Mulzer, Alexander Pilz, and Manuel Wettstein. An optimal algorithm for reconstructing point set order types from radial orderings. International Journal of Computational Geometry and Applications (IJCGA), 27 (1–2):57–83, March/June 2017.

CODEN IJCAEV. ISSN 0218-1959.

Ahn:2012:CDF

[AKS^{+}12] Hee-Kap Ahn, Christian Knauer, Marc Scherfenberg, Lena Schlipf, and Antoine Vigneron. Computing the discrete Fréchet distance with imprecise input. International Journal of Computational Geometry and Applications (IJCGA), 22(1):27–??, February 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Aoki:2001:TWB

[AL01] Kiyoko F. Aoki and D. T. Lee. Towards Web-based computing. International Journal of Computational Geometry and Applications (IJCGA), 11 (1):71–104, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Ahmed:2011:SDP

[AL11] Mustaq Ahmed and Anna Lubiw. Shortest descending paths: Towards an exact algorithm. International Journal of Computational Geometry and Applications (IJCGA), 21(4):431– 466, August 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Alsuwaiyel:1997:TAS

[Als97] Muhammad H. Alsuwaiyel. Two algorithms for the sum of diameters problem and a related problem. International Journal of Computational Geometry and Applications (IJCGA), 7(5):493–??, October 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Attali:2012:EDS

[ALS12] Dominique Attali, André Lieutier, and David Salinas. Efficient data structure for representing and simplifying simplicial complexes in high dimensions. International Journal of Computational Geometry and Applications (IJCGA), 22 (4):279–??, August 2012. CO-DEN IJCAEV. ISSN 0218-1959.

Atalay:2007:PIH

[AM07] F. Betul Atalay and David M.

Mount. Pointerless implementation of hierarchical simplicial meshes and efficient neighbor finding in arbitrary dimensions. International Journal of Computational Geometry and Applications (IJCGA), 17(6):595– 631, December 2007. CODEN IJCAEV. ISSN 0218-1959.

Arkin:1998:DTG

[AMM^{+}98] Esther M. Arkin, Henk Meijer, Joseph S. B. Mitchell, David Rappaport, and Steven S. Skiena. Decision trees for geometric models. International Journal of Computational Geometry and Applications (IJCGA), 8(3):343–??, June 1998. CODEN IJCAEV. ISSN 0218-1959.

Amit:2010:LGV

[AMP10] Yoav Amit, Joseph S. B.

Mitchell, and Eli Packer. Locating guards for visibility coverage of polygons. International Journal of Computational Geometry and Applications (IJCGA), 20(5):601–630, October 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Arkin:1997:GEF

[AMS97] Esther M. Arkinn, Joseph S. B. Mitchell, and Steven S. Skiena. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):1–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959.

Abdullah:2013:ABN

[AMV13] Amirali Abdullah, John Moeller, and Suresh Venkatasubramanian. Approximate Bregman near neighbors in sublinear time: Beyond the triangle inequality. International Journal of Computational Geometry and Applications (IJCGA), 23(4–5):253–??, 2013. CODEN IJCAEV. ISSN 0218-1959.

Anonymous:1997:AI

[Ano97] Anonymous. Author index. International Journal of Computational Geometry and Applications (IJCGA), 7(6):621–??, December 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:1998:AIV

[Ano98] Anonymous. Author index volume 8 (1998). International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):643–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Anonymous:1999:AI

[Ano99] Anonymous. Author index. International Journal of Computational Geometry and Applications (IJCGA), 9(6):619–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Anonymous:2000:AI

[Ano00] Anonymous. Author index. International Journal of Computational Geometry and Applications (IJCGA), 10(6):653–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Anonymous:2001:AI

[Ano01] Anonymous. Author index. International Journal of Computational Geometry and Applications (IJCGA), 11(6):683–??, 2001. CODEN IJCAEV. ISSN 0218-1959.

Anonymous:2002:AI

[Ano02] Anonymous. Author index. International Journal of Computational Geometry and Applications (IJCGA), 12(6):529–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Anonymous:2003:AIV

[Ano03] Anonymous. Author index volume 13 (2003). International Journal of Computational Geometry and Applications (IJCGA), 13(6):507–??, December 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2004:AIV

[Ano04] Anonymous. Author index volume 14 (2004). International Journal of Computational Geometry and Applications (IJCGA), 14(6):479–??, December 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2005:AIV

[Ano05] Anonymous. Author index volume 15 (2005). International Journal of Computational Geometry and Applications (IJCGA), 15(6):631–??, December 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2006:AIV

[Ano06] Anonymous. Author index volume 16 (2006). International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):617–??, December 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2007:AIV

[Ano07] Anonymous. Author index volume 17 (2007). International Journal of Computational Geometry and Applications (IJCGA), 17(6):633–635, December 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2008:AIV

[Ano08] Anonymous. Author index volume 18 (2008). International Journal of Computational Geometry and Applications (IJCGA), 18(6):627–629, December 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2009:AIV

[Ano09] Anonymous. Author index volume 19 (2009). International Journal of Computational Geometry and Applications (IJCGA), 19(6):617–619, December 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2010:AIV

[Ano10] Anonymous. Author index volume 20 (2010). International Journal of Computational Geometry and Applications (IJCGA), 20(6):751–754, December 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2011:AIV

[Ano11] Anonymous. Author index volume 21 (2011). International Journal of Computational Geometry and Applications (IJCGA), 21(6):699–702, December 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2012:AIV

[Ano12] Anonymous. Author index volume 22 (2012). International Journal of Computational Geometry and Applications (IJCGA), 22(6):593–??, December 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2013:AIV

[Ano13a] Anonymous. Author index volume 23 (2013). International Journal of Computational Geometry and Applications (IJCGA), 23(6):479–??, December 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Asano:2013:GEF

[ANO13b] Takao Asano, Shin-Ichi Nakano, and Yoshio Okamoto. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 23(2):73–??, April 2013. CODEN IJCAEV. ISSN 0218-1959.

Anonymous:2014:AIV

[Ano14] Anonymous. Author index: Volume 24 (2014). International Journal of Computational Geometry and Applications (IJCGA), 24(4):397–??, December 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2015:AIV

[Ano15] Anonymous. Author index volume 25 (2015). International Journal of Computational Geometry and Applications (IJCGA), 25(4):309–??, December 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2016:AIV

[Ano16] Anonymous. Author index volume 26 (2016). International Journal of Computational Geometry and Applications (IJCGA), 26(3–4):231, September/December 2016. CODEN IJCAEV. ISSN 0218-1959.

Anonymous:2017:AIV

[Ano17] Anonymous. Author index volume 27 (2017). International Journal of Computational Geometry and Applications (IJCGA), 27(4):327–??, December 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Anonymous:2018:AIV

[Ano18] Anonymous. Author index volume 28 (2018). International Journal of Computational Geometry and Applications (IJCGA), 28(4):399–400, December 2018. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195918990017.

Anonymous:2019:AIV

[Ano19] Anonymous. Author index volume 29 (2019). International Journal of Computational Geometry and Applications (IJCGA), 29(4):353–354, December 2019. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195919990012.

Anonymous:2020:AIV

[Ano20] Anonymous. Author index volume 30 (2020). International Journal of Computational Geometry and Applications (IJCGA), 30(3–4): ??, September–December 2020. ISSN 0218-1959. URL https:

//www.worldscientific.com/
doi/10.1142/S0218195920990010.

Anonymous:2021:AIV

[Ano21] Anonymous. Author index volume 31 (2021). International Journal of Computational Geometry and Applications (IJCGA), 31(4):??, December 2021. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195921990016.

Anonymous:2022:AIV

[Ano22] Anonymous. Author index volume 32 (2022). International Journal of Computational Geometry and Applications (IJCGA), 32(3–4): ??, September–December 2022. ISSN 0218-1959. URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0218195922990011.

Anonymous:2023:AIV

[Ano23] Anonymous. Author index volume 33 (2023). International Journal of Computational Geometry and Applications (IJCGA), 33(3–4): ??, September–December 2023. ISSN 0218-1959. URL https:

                                        //www.worldscientific.com/
                                        doi/10.1142/S0218195923990017.

Agarwal:1998:CGC

[AO98] Pankaj K. Agarwal and Joseph O’Rourke. Computational geometry column 34. International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):637–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Andersson:2000:ETF

[APS00] L.-E. Andersson, T. J. Peters, and N. F. Stewart. Equivalence of topological form for curvilinear geometric objects. International Journal of Computational Geometry and Applications (IJCGA), 10(6):609–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Andersen:2019:AED

[AR19] Patrick J. Andersen and Charl J. Ras. Algorithms for Euclidean degree bounded spanning tree problems. International Journal of Computational Geometry and Applications (IJCGA), 29(02):121– 160, June 2019. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195919500031.

Andrade:2001:EAC

[AS01] Marcus Vinícius Alvim Andrade and Jorge Stolfi. Exact algorithms for circles on the sphere. International Journal of Computational Geometry and Applications (IJCGA), 11 (3):267–290, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Alon:2008:CFC

[AS08a] Noga Alon and Shakhar Smorodinsky. Conflict-free colorings of shallow discs. International Journal of Computational Geometry and Applications (IJCGA), 18(6):599–604, December 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Alt:2008:CHD

[AS08b] Helmut Alt and Ludmila Scharf. Computing the Hausdorff distance between curved objects. International Journal of Computational Geometry and Applications (IJCGA), 18 (4):307–320, August 2008. CO-DEN IJCAEV. ISSN 0218-

Alt:2018:ASC

[AS18] Helmut Alt and Nadja Scharf. Approximating smallest containers for packing threedimensional convex objects. International Journal of Computational Geometry and Applications (IJCGA), 28(2):??, June 2018. ISSN 0218-1959.

Asano:2009:GEF

[Asa09] Tetsuo Asano. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 19 (2):93, April 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Akitaya:2018:RWS

[AT18] Hugo A. Akitaya and Csaba D. Tóth. Reconstruction of weakly simple polygons from their edges. International Journal of Computational Geometry and Applications (IJCGA), 28(2): ??, June 2018. ISSN 0218-1959.

Atallah:1999:IHB

[Ata99] M. J. Atallah. An improved hypercube bound for multisearching and its applications. International Journal of Computational Geometry and Applications (IJCGA), 9(1):29–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Ahn:2014:GEF

[AV14] Hee-Kap Ahn and Antoine Vigneron. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 24(4):

259–??, December 2014. CO-DEN IJCAEV. ISSN 0218-1959.

Barequet:1998:DPD

[Bar98] Gill Barequet. DCEL: a polyhedral database and programming environment. International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):619–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Barequet:2005:GEF

[Bar05] Gill Barequet. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 15 (4):325–??, August 2005. CO-DEN IJCAEV. ISSN 0218-1959.

Barequet:2013:EF

[Bar13] Gill Barequet. Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 23(4–5): 231–??, 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Bereg:2010:VDP

[BBB^{+}10] Sergey Bereg, Kevin Buchin, Maike Buchin, Marina Gavrilova, and Binhai Zhu. Voronoi diagram of polygonal chains under the discrete Fréchet distance. International Journal of Computational Geometry and Applications (IJCGA), 20(4):471– 484, August 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Berry:2019:LSP

[BBB^{+}19] Lindsay Berry, Andrew Beveridge, Jane Butterfield, Volkan Isler, Zachary Keller, Alana Shine, and Junyi Wang. Lineof-sight pursuit in monotone and scallop polygons. International Journal of Computational Geometry and Applications (IJCGA), 29(4):307– 351, December 2019. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195919500122.

Bose:2002:ORC

[BBC^{+}02] Prosenjit Bose, Andrej Brodnik, Svante Carlsson, Erik D. Demaine, Rudolf Fleischer, Alejandro López-Ortiz, Pat Morin, and J. Ian Munro. Online routing in convex subdivisions. International Journal of Computational Geometry and Applications (IJCGA), 12(4): 283–??, August 2002. CODEN IJCAEV. ISSN 0218-1959.

Blandford:2005:CRS

[BBCK05] Daniel K. Blandford, Guy E. Blelloch, David E. Cardoze, and Clemens Kadow. Compact representations of simplicial meshes in two and three dimensions. International Journal of Computational Geometry and Applications (IJCGA), 15 (1):3–??, February 2005. CO-DEN IJCAEV. ISSN 0218-1959.

Bernardini:1999:ARC

[BBCS99] Fausto Bernardini, Chandrajit L. Bajaj, Jindong Chen, and Daniel R. Schikore. Automatic reconstruction of 3D CAD models from digital scans. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):327– ??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Buchin:2011:DCP

[BBG^{+}11] Kevin Buchin, Maike Buchin, Joachim Gudmundsson, Maarten Löffler, and Jun Luo. Detecting commuting patterns by clustering subtrajectories. International Journal of Computational Geometry and Applications (IJCGA), 21(3):253–282, June 2011. CODEN IJCAEV. ISSN 0218-1959.

Borgelt:2008:FPA

[BBL08] Magdalene Grantson Borgelt, Christian Borgelt, and Christos Levcopoulos. Fixed parameter algorithms for the minimum weight triangulation problem. International Journal of Computational Geometry and Applications (IJCGA), 18(3):185– 220, June 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Bekker:2009:RTC

[BBR09] Henk Bekker, Axel A. Brink, and Jos B. T. M. Roerdink. Reducing the time complexity and identifying ill-posed problem instances of Minkowski sum based similarity calculations. International Journal of Computational Geometry and Applications (IJCGA), 19(5):441– 456, October 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Bae:2006:VDT

[BC06] Sang Won Bae and Kyung-Yong Chwa. Voronoi diagrams for a transportation network on the Euclidean plane. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3):117–??, June 2006. CODEN IJCAEV. ISSN 0218-1959.

Boissonnat:2000:CLC

[BCD^{+}00] J.-D. Boissonnat, J. Czyzowicz, O. Devillers, J. Urrutia, and M. Yvinec. Computing largest circles separating two sets of segments. International Journal of Computational Geometry and Applications (IJCGA), 10 (1):41–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2007:GAA

[BCHS07] Prosenjit Bose, Narcís Coll, Ferran Hurtado, and J. Antoni Sellarès. A general approximation algorithm for planar maps with applications. International Journal of Computational Geometry and Applications (IJCGA), 17(6):529–554, December 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Baran:2005:OAA

[BD05] Ilya Baran and Erik D. Demaine. Optimal adaptive algorithms for finding the nearest and farthest point on a parametric black-box curve. International Journal of Computational Geometry and Applications (IJCGA), 15(4):327–??, August 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Bereg:2014:CSP

[BDBF^{+}14] S. Bereg, J. M. Díaz-Báñez, M. Fort, M. A. Lopez, P. Pérez-Lantero, and J. Urrutia. Continuous surveillance of points by rotating floodlights. International Journal of Computational Geometry and Applications (IJCGA), 24(3):183–??, September 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2012:PAY

[BDD^{+}12] Prosenjit Bose, Mirela Damian, Karim Douïeb, Joseph O’Rourke, Ben Seamone, Michiel Smid, and Stefanie Wuhrer. π/2-angle Yao graphs are spanners. International Journal of Computational Geometry and Applications (IJCGA), 22(1):61–??, February 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2017:COR

[BDDT17] Prosenjit Bose, Jean-Lou De Carufel, Stephane Durocher, and Perouz Taslakian. Competitive online routing on Delaunay triangulations. International Journal of Computational Geometry and Applications (IJCGA), 27(4):241–??, December 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2002:DMW

[BDE02] Prosenjit Bose, Luc Devroye, and William Evans. Diamonds are not a minimum weight triangulation’s best friend. International Journal of Computational Geometry and Applications (IJCGA), 12(6):445–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Boissonnat:2013:SDT

[BDG13] Jean-Daniel Boissonnat, Ramsay Dyer, and Arijit Ghosh. The stability of Delaunay triangulations. International Journal of Computational Geometry and Applications (IJCGA), 23(4–5):303–??, 2013. CODEN IJCAEV. ISSN 0218-1959.

Boissonnat:2014:DSP

[BDG14] Jean-Daniel Boissonnat, Ramsay Dyer, and Arijit Ghosh. Delaunay stability via perturbations. International Journal of Computational Geometry and Applications (IJCGA), 24 (2):125–??, June 2014. CODEN IJCAEV. ISSN 0218-1959.

Butler:2013:CPT

[BDGT13] Steve Butler, Erik Demaine, Ron Graham, and Tomohiro Tachi. Constructing points through folding and intersection. International Journal of Computational Geometry and Applications (IJCGA), 23(1): 49–??, February 2013. CODEN IJCAEV. ISSN 0218-1959.

Benkert:2010:FPP

[BDGW10] Marc Benkert, Bojan Djordjevic, Joachim Gudmundsson, and Thomas Wolle. Finding popular places. International Journal of Computational Geometry and Applications (IJCGA), 20(1):19–42, February 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Breukelaar:2004:THE

[BDH^{+}04] Ron Breukelaar, Erik D. Demaine, Susan Hohenberger, Hendrik Jan Hoogeboom, Walter A. Kosters, and David Liben-Nowell. Tetris is hard, even to approximate. International Journal of Computational Geometry and Applications (IJCGA), 14(1–2):41–??, April 2004. CODEN IJCAEV. ISSN 0218-1959.

Bose:2012:PGT

[BDH^{+}12] Prosenjit Bose, Vida Dujmović, Ferran Hurtado, John Iacono, Stefan Langerman, Henk Meijer, Vera Sacristán, Maria Saumell, and David R. Wood. Proximity Graphs: E, δ, Δ, χ and ω. International Journal of Computational Geometry and Applications (IJCGA), 22(5): 439–??, October 2012. CODEN IJCAEV. ISSN 0218-1959.

Basch:2003:PAD

[BDIZ03] Julien Basch, Harish Devarajan, Piotr Indyk, and Li Zhang. Probabilistic analysis for discrete attributes of moving points. International Journal of Computational Geometry and Applications (IJCGA), 13(1):5– ??, February 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Bereg:2010:MAI

[BDJ10] Sergey Bereg, Adrian Dumitrescu, and Minghui Jiang. Maximum area independent sets in disk intersection graphs. International Journal of Computational Geometry and Applications (IJCGA), 20(2):105– 118, April 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2020:LRC

[BDM^{+}20] Prosenjit Bose, Stephane Durocher, Debajyoti Mondal, Maxime Peabody, Matthew Skala, and Mohammad Abdul Wahid. Local routing in convex subdivisions. International Journal of Computational Geometry and Applications (IJCGA), 30(01):1–17, March 2020. CO-DEN IJCAEV. ISSN 0218-

                                        1959.  URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0218195920500016.

Bereg:2008:SDP

[BDP08] Sergey Bereg, Adrian Dumitrescu, and János Pach. Sliding disks in the plane. International Journal of Computational Geometry and Applications (IJCGA), 18(5):373–387, October 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Bern:2000:QMC

[BE00] M. Bern and D. Eppstein. Quadrilateral meshing by circle packing. International Journal of Computational Geometry and Applications (IJCGA), 10 (4):347–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Benroummane:2023:SRS

[Ben23] Abderrazzak Benroummane.

Some results on semi-symmetric spaces. International Journal of Computational Geometry and Applications (IJCGA), 33(3–4):65–84, September– December 2023. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195923500024.

Berzins:2000:SBM

[Ber00] M. Berzins. Solution-based mesh quality indicators for triangular and tetrahedral meshes. International Journal of Computational Geometry and Applications (IJCGA), 10 (3):333–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Bereg:2004:CHD

[Ber04] Sergey Bereg. Cylindrical hierarchy for deforming necklaces. International Journal of Computational Geometry and Applications (IJCGA), 14(1–2):3–??, April 2004. CODEN IJCAEV. ISSN 0218-1959.

Bereg:2005:AMB

[Ber05] Sergey Bereg. An approximate morphing between polylines. International Journal of Computational Geometry and Applications (IJCGA), 15(2): 193–??, April 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Bespamyatnikh:2002:OMB

[Bes02] Sergei Bespamyatnikh. An optimal morphing between polylines. International Journal of Computational Geometry and Applications (IJCGA), 12(3): 217–??, 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Bespamyatnikh:2003:CCP

[Bes03] Sergei Bespamyatnikh. Computing closest points for segments. International Journal of Computational Geometry and Applications (IJCGA), 13(5): 419–??, October 2003. CODEN IJCAEV. ISSN 0218-1959.

Bern:1999:PCQ

[BET99] M. Bern, D. Eppstein, and S.- H. Teng. Parallel construction of quadtrees and quality triangulations. International Journal of Computational Geometry and Applications (IJCGA), 9(6):517–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2003:PAG

[BEW03] Prosenjit Bose, Hazel Everett, and Stephen Wismath. Properties of arrangement graphs. International Journal of Computational Geometry and Applications (IJCGA), 13(6):447–??, December 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Blazj:2021:ELR

[BFL21] Václav Blažj, Jiří Fiala, and Giuseppe Liotta. On edgelength ratios of partial 2-trees.

International Journal of Computational Geometry and Applications (IJCGA), 31(2–3):141– 162, June–September 2021. ISSN 0218-1959. URL https:

//www.worldscientific.com/
doi/10.1142/S0218195921500072.

Bereg:2014:EDC

[BFMFP^{+}14] S. Bereg, R. Fabila-Monroy, D. Flores-Peñaloza, M. A. Lopez, and P. Pérez-Lantero. Embedding the double circle in a square grid of minimum size. International Journal of Computational Geometry and Applications (IJCGA), 24(3):247–??, September 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Burnikel:2001:EGC

[BFS01] Christoph Burnikel, Stefan Funke, and Michael Seel. Exact geometric computation using cascading. International Journal of Computational Geometry and Applications (IJCGA), 11 (3):245–266, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Bhosle:2005:EAA

[BG05] Amit M. Bhosle and Teofilo F.

Gonzalez. Exact and approximation algorithms for finding an optimal bridge connecting two simple polygons. International Journal of Computational Geometry and Applications (IJCGA), 15(6):609–??, December 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Batog:2011:IBN

[BG11a] Guillaume Batog and Xavier Goaoc. Inflating balls is NPhard. International Journal of Computational Geometry and Applications (IJCGA), 21(4): 403–415, August 2011. CODEN IJCAEV. ISSN 0218-1959.

Biedl:2011:STO

[BG11b] Therese Biedl and Burkay Genç. Stoker’s theorem for orthogonal polyhedra. International Journal of Computational Geometry and Applications (IJCGA), 21(4):383–391, August 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Buchin:2014:DPL

[BG14] Kevin Buchin and Dirk H. P. Gerrits. Dynamic point labeling is strongly Pspace-complete. International Journal of Computational Geometry and Applications (IJCGA), 24(4):373–??, December 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Benkert:2009:PTA

[BGK^{+}09] Marc Benkert, Joachim Gudmundsson, Christian Knauer, René Van Oostrum, and Alexander Wolff. A polynomialtime approximation algorithm for a geometric dispersion problem. International Journal of Computational Geometry and Applications (IJCGA), 19(3): 267–288, June 2009. CODEN IJCAEV. ISSN 0218-1959.

Bose:1997:FP

[BGL^{+}97] Prosenjit Bose, Leonidas Guibas, Anna Lubiw, Mark Overmars, Diane Souvaine, and Jorge Urrutia. The floodlight problem. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):153– 163, February–April 1997. CO-DEN IJCAEV. ISSN 0218-1959.

Bridgeman:1999:GDT

[BGT99] Stina Bridgeman, Ashim Garg, and Roberto Tamassia. A graph drawing and translation service on the World Wide Web. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):419–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Brass:2010:LBA

[BHLL10] Peter Brass, Ferran Hurtado, Benjamin Lafreniere, and Anna Lubiw. A lower bound on the area of a 3-coloured disk packing. International Journal of Computational Geometry and Applications (IJCGA), 20 (3):341–360, June 2010. CO-DEN IJCAEV. ISSN 0218-1959.

Brass:2003:CCA

[BHLM03] Peter Braß, Laura Heinrich-Litan, and Pat Morin. Computing the center of area of a convex polygon. International Journal of Computational Geometry and Applications (IJCGA), 13(5):439–??, October 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Biedl:2011:RCP

[BHLO11] Therese Biedl, Masud Hasan, and Alejandro López-Ortiz. Reconstructing convex polygons and convex polyhedra from edge and face counts in orthogonal projections. International Journal of Computational Geometry and Applications (IJCGA), 21(2):215–239, April 2011. CODEN IJCAEV. ISSN 0218-1959.

Berkoff:2011:PIS

[BHMW11] A. M. Berkoff, J. M. Henle, A. E. Mcdonough, and A. P. Wesolowski. Possibilities and impossibilities in square-tiling. International Journal of Computational Geometry and Applications (IJCGA), 21(5):545– 558, October 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Barequet:2001:PCT

[BHP01] Gill Barequet and Sariel Har-Peled. Polygon containment and translational min-Hausdorff-distance between segment sets are 3 SUM-hard. International Journal of Computational Geometry and Applications (IJCGA), 11(4):465– 474, 2001. CODEN IJCAEV. ISSN 0218-1959.

Biedl:2016:PMW

[BHP16] Therese Biedl, Stefan Huber, and Peter Palfrader. Planar matchings for weighted straight skeletons. International Journal of Computational Geometry and Applications (IJCGA), 26(3–4):211–229, September/ December 2016. CODEN IJ-CAEV. ISSN 0218-1959.

Banerjee:2021:VGD

[BI21] Debangshu Banerjee and R. Inkulu.

Vertex guarding for dynamic orthogonal art galleries. International Journal of Computational Geometry and Applications (IJCGA), 31(2–3):123– 140, June–September 2021. ISSN 0218-1959. URL https:

//www.worldscientific.com/
doi/10.1142/S0218195921500060.

Bhattacharjee:2022:VFT

[BI22] Sukanya Bhattacharjee and R. Inkulu. Vertex faulttolerant geometric spanners for weighted points. International Journal of Computational Geometry and Applications (IJCGA), 32(3–4): 175–199, September–December 2022. ISSN 0218-1959. URL

https://www.worldscientific.
com/doi/10.1142/S021819592250008X.

Binger:2002:PUI

[Bin02] David Binger. Polyhedra with unguarded interiors. International Journal of Computational Geometry and Applications (IJCGA), 12(3):173–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Brass:2002:TCD

[BK02] Peter Brass and Christian Knauer. Testing the congruence of d-dimensional point sets. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2):115–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Boutin:2007:WPC

[BK07] Mireille Boutin and Gregor Kemper. Which point configurations are determined by the distribution of their pairwise distances? International Journal of Computational Geometry and Applications (IJCGA), 17 (1):31–43, February 2007. CO-DEN IJCAEV. ISSN 0218-1959.

Bohler:2014:AVD

[BK14] Cecilia Bohler and Rolf Klein. Abstract Voronoi diagrams with disconnected regions. International Journal of Computational Geometry and Applications (IJCGA), 24(4):347–??, December 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Bringmann:2017:IAF

[BK17] Karl Bringmann and Marvin Künnemann. Improved approximation for Fréchet distance on c-packed curves matching conditional lower bounds. International Journal of Computational Geometry and Applications (IJCGA), 27(1–2):85– 119, March/June 2017. CO-DEN IJCAEV. ISSN 0218-1959.

Buot:2018:WES

[BK18] Jude Buot and Mikio Kano.

Weight-equitable subdivision of red and blue points in the plane. International Journal of Computational Geometry and Applications (IJCGA), 28(1): 39–??, March 2018. CODEN IJ-CAEV. ISSN 0218-1959.

Bae:2009:OCC

[BKC09] Sang Won Bae, Jae-Hoon Kim, and Kyung-Yong Chwa. Optimal construction of the city Voronoi diagram. International Journal of Computational Geometry and Applications (IJCGA), 19(2):95–117, April 2009. CODEN IJCAEV. ISSN 0218-1959.

Bohler:2017:AVD

[BKL17] Cecilia Bohler, Rolf Klein, and Chih-Hung Liu. Abstract Voronoi diagrams from closed bisecting curves. International Journal of Computational Geometry and Applications (IJCGA), 27(3):221–240, September 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Brass:2011:ACP

[BKN^{+}11] Peter Brass, Christian Knauer, Hyeon-Suk Na, Chan-Su Shin, and Antoine Vigneron. The aligned K-center problem. International Journal of Computational Geometry and Applications (IJCGA), 21(2):157–178, April 2011. CODEN IJCAEV. ISSN 0218-1959.

Bespamyatnikh:2000:OFL

[BKST00] S. Bespamyatnikh, K. Kedem, M. Segal, and A. Tamir. Optimal facility location under various distance functions. International Journal of Computational Geometry and Applications (IJCGA), 10(5):523–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Boissonnat:2003:PTA

[BL03] Jean-Daniel Boissonnat and Sylvain Lazard. A polynomialtime algorithm for computing shortest paths of bounded curvature amidst moderate obstacles. International Journal of Computational Geometry and Applications (IJCGA), 13(3): 189–??, June 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2002:IAS

[BM02] Prosenjit Bose and Pat Morin. An improved algorithm for subdivision traversal without extra storage. International Journal of Computational Geometry and Applications (IJCGA), 12 (4):297–??, August 2002. CO-DEN IJCAEV. ISSN 0218-1959.

Binder:2012:BII

[BM12] Thomas Binder and Thomas Martinetz. On the boundedness of an iteration involving points on the hypersphere. International Journal of Computational Geometry and Applications (IJCGA), 22(6):499–??, December 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Ben-Moshe:2000:OFL

[BMKS00] B. Ben-Moshe, M. J. Katz, and M. Segal. Obnoxious facility location: Complete service with minimal harm. International Journal of Computational Geometry and Applications (IJCGA), 10(6):581–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Bint:2019:PER

[BMSN19] Gregory Bint, Anil Maheshwari, Michiel Smid, and Subhas C. Nandy. Partial enclosure range searching. International Journal of Computational Geometry and Applications (IJCGA), 29(1):73– 93, March 2019. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195919500018.

Bose:2011:CEI

[BMSS11] Prosenjit Bose, Mercè Mora, Carlos Seara, and Saurabh Sethia. On computing enclosing isosceles triangles and related problems. International Journal of Computational Geometry and Applications (IJCGA), 21 (1):25–45, February 2011. CO-DEN IJCAEV. ISSN 0218-1959.

Bhattacharya:1999:CSW

[BMT99] B. K. Bhattacharya, A. Mukhopadhyay, and G. T. Toussaint. Computing a shortest weakly externally visible line segment for a simple polygon. International Journal of Computational Geometry and Applications (IJCGA), 9(1):81–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Biedl:2000:TPM

[BMT00] T. C. Biedl, B. P. Madden, and I. G. Tollis. The three-phase method: a unified approach to orthogonal graph drawing. International Journal of Computational Geometry and Applications (IJCGA), 10(6):553–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Bose:2016:PO

[BMvR16] Prosenjit Bose, Pat Morin, and André van Renssen. The price of order. International Journal of Computational Geometry and Applications (IJCGA), 26(3–4):135–149, September/ December 2016. CODEN IJ-CAEV. ISSN 0218-1959.

Brass:2010:GPT

[BNS10] Peter Brass, Hyeon-Suk Na, and Chan-Su Shin. Guarding a polygon from two nearlyopposite directions. International Journal of Computational Geometry and Applications (IJCGA), 20(3):327–339, June 2010. CODEN IJCAEV. ISSN 0218-1959.

Bardhan:2009:GPM

[BRD09] Debabrata Bardhan, Sansanka Roy, and Sandip Das. Guard placement for maximizing Lvisibility exterior to a convex polygon. International Journal of Computational Geometry and Applications (IJCGA), 19 (4):357–370, August 2009. CO-DEN IJCAEV. ISSN 0218-1959.

Bremner:2000:PVG

[BS00] D. Bremner and T. Shermer. Point visibility graphs and O-convex cover. International Journal of Computational Geometry and Applications (IJCGA), 10(1):55–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Brass:2005:LBL

[BS05] Peter Brass and Mehrbod Sharifi. A lower bound for Lebesgue’s universal cover problem. International Journal of Computational Geometry and Applications (IJCGA), 15 (5):537–??, October 2005. CO-DEN IJCAEV. ISSN 0218-1959.

Barequet:2008:MP

[BS08] Gill Barequet and Aya Steiner.

On the matability of polygons. International Journal of Computational Geometry and Applications (IJCGA), 18(5):469– 506, October 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Bae:2012:ODV

[BS12] Sang Won Bae and Chan-Su Shin. The onion diagram: a Voronoi-like tessellation of a planar line space and its applications. International Journal of Computational Geometry and Applications (IJCGA), 22 (1):3–??, February 2012. CO-DEN IJCAEV. ISSN 0218-1959.

Bueno:2016:CTM

[BS16] Lucas Moutinho Bueno and Jorge Stolfi. 3-colored triangulation of 2D maps. International Journal of Computational Geometry and Applications (IJCGA), 26(2):111–133, June 2016. CODEN IJCAEV. ISSN 0218-1959.

Bueno:2017:CTM

[BS17] Lucas Moutinho Bueno and Jorge Stolfi. 4-colored triangulation of 3-maps. International Journal of Computational Geometry and Applications (IJCGA), 27(4):297–??, December 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Baek:1999:CTS

[BSC99] N. Baek, S.-Y. Shin, and K.- Y. Chwa. On computing translational swept volumes. International Journal of Computational Geometry and Applications (IJCGA), 9(3):293–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Baek:2000:TDT

[BSC00] N. Baek, S.-Y. Shin, and K.- Y. Chwa. Three-dimensional topological sweep for computing rotational swept volumes of polyhedral objects. International Journal of Computational Geometry and Applications (IJCGA), 10(2):131–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Bose:2009:DDT

[BSX09] Prosenjit Bose, Michiel Smid, and Daming Xu. Delaunay and diamond triangulations contain spanners of bounded degree. International Journal of Computational Geometry and Applications (IJCGA), 19(2):119–140, April 2009. CODEN IJCAEV. ISSN 0218-1959.

Bose:2005:GMR

[BV05] Prosenjit Bose and Marc Van Kreveld. Generalizing monotonicity: on recognizing special classes of polygons and polyhedra. International Journal of Computational Geometry and Applications (IJCGA), 15(6):591–??, December 2005. CODEN IJCAEV. ISSN 0218-1959.

Biedl:2013:PSE

[BV13] Therese Biedl and Martin Vatshelle. The point-set embeddability problem for plane graphs. International Journal of Computational Geometry and Applications (IJCGA), 23(4–5):357–??, 2013. CODEN IJCAEV. ISSN 0218-1959.

Borgelt:2011:GDC

[BVL11] Magdalene G. Borgelt, Marc Van Kreveld, and Jun Luo. Geodesic disks and clustering in a simple polygon. International Journal of Computational Geometry and Applications (IJCGA), 21(6):595–608, December 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Bose:2019:SPY

[BvR19] Prosenjit Bose and André van Renssen. Spanning properties of Yao and θ-graphs in the presence of constraints. International Journal of Computational Geometry and Applications (IJCGA), 29(02): 95–120, June 2019. CO-DEN IJCAEV. ISSN 0218-

                                        1959.  URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S021819591950002X.

Bajaj:2003:MSC

[BXHN03] Chandrajit L. Bajaj, Guoliang Xu, Robert J. Holt, and Arun N. Netravali. NURBS approximation of A-splines and A-patches. International Journal of Computational Geometry and Applications (IJCGA), 13 (5):359–??, October 2003. CO-DEN IJCAEV. ISSN 0218-1959.

Beketayev:2018:MEA

[BYM^{+}18] Kenes Beketayev, Damir Yeliussizov, Dmitriy Morozov, Gunther H. Weber, and Bernd Hamann. Measuring the error in approximating the sub-level set topology of sampled scalar data. International Journal of Computational Geometry and Applications (IJCGA), 28(1): 57–??, March 2018. CODEN IJCAEV. ISSN 0218-1959.

Brazil:2014:UOS

[BZ14] Marcus Brazil and Martin Zachariasen. The uniform orientation Steiner tree problem is NP-hard. International Journal of Computational Geometry and Applications (IJCGA), 24 (2):87–??, June 2014. CODEN IJCAEV. ISSN 0218-1959.

Carlsson:2015:BMP

[CARB15] John Gunnar Carlsson, Benjamin Armbruster, Saladi Rahul, and Haritha Bellam. A bottleneck matching problem with edge-crossing constraints. International Journal of Computational Geometry and Applications (IJCGA), 25(4):245–??, December 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Chen:2006:SVC

[CC06] Chao Chen and Ho-Lun Cheng.

Superimposing Voronoi complexes for shape deformation. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3):159– ??, June 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Carlsson:2006:ATM

[CCD06] Erik Carlsson, Gunnar Carlsson, and Vin De Silva. An algebraic topological method for feature identification. International Journal of Computational Geometry and Applications (IJCGA), 16(4):291–??, August 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Cheng:2017:NWR

[CCJV17] Siu-Wing Cheng, Man-Kwun Chiu, Jiongxin Jin, and Antoine Vigneron. Navigating weighted regions with scattered skinny tetrahedra. International Journal of Computational Geometry and Applications (IJCGA), 27(1–2):13–32, March/June 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Chang:2006:SPA

[CCK^{+}06] Ee-Chien Chang, Sung Woo Choi, Do Yong Kwon, Hyungju Park, and Chee K. Yap. Shortest path amidst disc obstacles is computable. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):567–??, December 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Carmi:2019:MLO

[CCMS19] Paz Carmi, Farah Chanchary, Anil Maheshwari, and Michiel Smid. The most likely object to be seen through a window. International Journal of Computational Geometry and Applications (IJCGA), 29(4): 269–287, December 2019. CO-DEN IJCAEV. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195919500092.

Chen:2003:SEA

[CD03] Danny Z. Chen and Ovidiu Daescu. Space-efficient algorithms for approximating polygonal curves in twodimensional space. International Journal of Computational Geometry and Applications (IJCGA), 13(2):95–??, April 2003. CODEN IJCAEV. ISSN 0218-1959.

Charlton:2012:GC

[CDD^{+}12] David Charlton, Erik D. Demaine, Martin L. Demaine, Vida Dujmović, Pat Morin, and Ryuhei Uehara. Ghost chimneys. International Journal of Computational Geometry and Applications (IJCGA), 22(3): 207–??, June 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Cabello:2009:MAO

[CDG^{+}09] Sergio Cabello, Mark De Berg, Panos Giannopoulos, Christian Knauer, René Van Oostrum, and Remco C. Veltkamp. Maximizing the area of overlap of two unions of disks under rigid motion. International Journal of Computational Geometry and Applications (IJCGA), 19(6):533–556, December 2009. CODEN IJCAEV. ISSN 0218-1959.

Carmi:2015:MDS

[CDJ^{+}15] Paz Carmi, Gautam K. Das, Ramesh K. Jallu, Subhas C. Nandy, Prajwal R. Prasad, and Yael Stein. Minimum dominating set problem for unit disks revisited. International Journal of Computational Geometry and Applications (IJCGA), 25 (3):??, September 2015. CO-DEN IJCAEV. ISSN 0218-

Chen:2001:GPQ

[CDK01] Danny Z. Chen, Ovidiu Daescu, and Kevin S. Klenk. On geometric path query problems. International Journal of Computational Geometry and Applications (IJCGA), 11(6):617–??, 2001. CODEN IJCAEV. ISSN 0218-1959.

Calinescu:2005:SPA

[CDKW05] Gruia Călinescu, Adrian Dumitrescu, Howard Karloff, and Peng-Jun Wan. Separating points by axis-parallel lines. International Journal of Computational Geometry and Applications (IJCGA), 15(6):575–??, December 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Cheng:2005:QMP

[CDRR05] Siu-Wing Cheng, Tamal K. Dey, Edgar A. Ramos, and Tathagata Ray. Quality meshing of polyhedra with small angles. International Journal of Computational Geometry and Applications (IJCGA), 15(4): 421–??, August 2005. CODEN IJCAEV. ISSN 0218-1959.

Chen:2001:CSC

[CDWK01] Wei Chen, Xiaowen Deng, Koichi Wada, and K. Kawaguchi. Constructing a strongly convex superhull of points. International Journal of Computational Geometry and Applications (IJCGA), 11(5):487–502, 2001. CODEN IJCAEV. ISSN 0218-1959.

Cheong:2007:PSQ

[CEK^{+}07] Otfried Cheong, Hazel Everett, Hyo-Sil Kim, Sylvain Lazard, and René Schott. Parabola separation queries and their application to stone throwing. International Journal of Computational Geometry and Applications (IJCGA), 17(4):349–360, August 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Czyzowicz:1997:STP

[CER97] Jurek Czyzowicz, Hazel Everett, and Jean-Marc Robert. Separating translates in the plane: Combinatorial bounds and an algorithm. International Journal of Computational Geometry and Applications (IJCGA), 7(6):551–??, December 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Cavazza:2015:EMP

[CFL15] Niccolò Cavazza, Massimo Ferri, and Claudia Landi. Estimating multidimensional persistent homology through a finite sampling. International Journal of Computational Geometry and Applications (IJCGA), 25(3):??, September 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Crauser:2001:REM

[CFM^{+}01] A. Crauser, P. Ferragina, K. Mehlhorn, U. Meyer, and E. A. Ramos. Randomized external-memory algorithms for line segment intersection and other geometric problems. International Journal of Computational Geometry and Applications (IJCGA), 11(3):305–337, 2001. CODEN IJCAEV. ISSN 0218-1959.

Cortes:2012:RBS

[CGG^{+}12] Carmen Cortés, Delia Garijo, María Ángeles Garrido, Clara I. Grima, Alberto Márquez, Auxiliadora Moreno-González, Jesús Valenzuela, and María Trinidad Villar. Reporting bichromatic segment intersections from point sets. International Journal of Computational Geometry and Applications (IJCGA), 22(5):421– ??, October 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Chuon:2011:SCB

[CGJS11] Chansophea Chuon, Sumanta Guha, Paul Janecek, and Nguyen Duc Cong Song. Simplipoly: Curvature-based polygonal curve simplification. International Journal of Computational Geometry and Applications (IJCGA), 21(4):417– 429, August 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Chan:2001:ESD

[Cha01] Timothy M. Chan. On enumerating and selecting distances. International Journal of Computational Geometry and Applications (IJCGA), 11(3):291– 304, 2001. CODEN IJCAEV. ISSN 0218-1959.

Chan:2002:ADW

[Cha02] Timothy M. Chan. Approximating the diameter, width, smallest enclosing cylinder, and minimum-width annulus. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2):67–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Chan:2012:TPA

[Cha12] Timothy M. Chan. Three problems about dynamic convex hulls. International Journal of Computational Geometry and Applications (IJCGA), 22 (4):341–??, August 2012. CO-DEN IJCAEV. ISSN 0218-1959.

Chen:1998:DWV

[Che98] Danny Z. Chen. Determining weak visibility of a polygon from an edge in parallel. International Journal of Computational Geometry and Applications (IJCGA), 8(3):277–??, June 1998. CODEN IJCAEV. ISSN 0218-1959.

Chen:2010:GSA

[Che10]  Eric Y. Chen.  Geometric

streaming algorithm with a sorting primitive. International Journal of Computational Geometry and Applications (IJCGA), 20(1):3–18, February 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Chen:2004:GAS

[CHL^{+}04] Danny Z. Chen, Xiaobo S.

Hu, Shuang (Sean) Luan, Chao Wang, and Xiaodong Wu. Geometric algorithms for static leaf sequencing problems in radiation therapy. International Journal of Computational Geometry and Applications (IJCGA), 14(4–5):311–??, October 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Chen:2006:GGA

[CHL^{+}06] Danny Z. Chen, Xiaobo X. Hu, Shuang Luan, Shahid A. Naqvi, Chao Wang, and Cedric X. Yu. Generalized geometric approaches for leaf sequencing problems in radiation therapy. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3):175– ??, June 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Chao:2013:GEF

[CHL13] Kun-Mao Chao, Tsan-Sheng Hsu, and Der-Tsai Lee. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 23(6):425–??, December 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Choset:1999:NAC

[Cho99] Howie Choset. Nonsmooth analysis, convex analysis, and their applications to motion planning. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):447–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Cano:2014:SSP

[CHU14] Javier Cano, Ferran Hurtado, and Jorge Urrutia. Stabbing simplices of point sets with kflats. International Journal of Computational Geometry and Applications (IJCGA), 24(3): 237–??, September 2014. CO-DEN IJCAEV. ISSN 0218-1959.

Chen:2002:OPC

[CHW02] Danny Z. Chen, Xiaobo S. Hu, and Xiaodong Wu. Optimal polygon cover problems and applications. International Journal of Computational Geometry and Applications (IJCGA), 12 (4):309–??, August 2002. CO-DEN IJCAEV. ISSN 0218-1959.

Chen:2008:MRB

[CHW^{+}08] Danny Z. Chen, Xiaobo S.

Hu, Chao Wang, Shuang Luan, and Xiaodong Wu. Mountain reduction, block matching, and applications in intensitymodulated radiation therapy. International Journal of Computational Geometry and Applications (IJCGA), 18(1-2):63– 106, April 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Chwa:2006:GAG

[CJK^{+}06] Kyung-Yong Chwa, Byung-Cheol Jo, Christian Knauer, Esther Moet, René Van Oostrum, and Chan-Su Shin. Guarding art galleries by guarding witnesses. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3):205–??, June 2006. CODEN IJCAEV. ISSN 0218-1959.

Cheng:2012:ASH

[CJVW12] Siu-Wing Cheng, Jiongxin Jin, Antoine Vigneron, and Yajun Wang. Approximate shortest homotopic paths in weighted regions. International Journal of Computational Geometry and Applications (IJCGA), 22 (1):83–??, February 2012. CO-DEN IJCAEV. ISSN 0218-1959.

Cai:1997:CVI

[CK97a] Leizhen Cai and J. Mark Keil. Computing visibility information in an inaccurate simple polygon. International Journal of Computational Geometry and Applications (IJCGA), 7(6):515–??, December 1997. CODEN IJCAEV. ISSN 0218-1959.

Chrobak:1997:CGD

[CK97b] Marek Chrobak and Goos Kant. Convex grid drawings of 3-connected planar graphs. International Journal of Computational Geometry and Applications (IJCGA), 7(3):211–??, June 1997. CODEN IJCAEV. ISSN 0218-1959.

Culver:2003:HAD

[CKMK03] Tim Culver, John Keyser, Dinesh Manocha, and Shankar Krishnan. A hybrid approach for determinant signs of moderate-sized matrices. International Journal of Computational Geometry and Applications (IJCGA), 13(5):399–??, October 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Chan:1993:LAN

[CL93] Kwong-Fai Chan and Tak Wah Lam. An on-line algorithm for navigating in an unknown environment. International Journal of Computational Geometry and Applications (IJCGA), 3(??):227–244, 1993. CODEN IJCAEV. ISSN 0218-1959.

Chen:2009:EF

[CL09] Danny Z. Chen and D. T.

Lee. Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 19(3):213–214, June 2009. CODEN IJCAEV. ISSN 0218-1959.

Cheong:2013:SSD

[CL13] Otfried Cheong and Changryeol Lee. Single-source dilationbounded minimum spanning trees. International Journal of Computational Geometry and Applications (IJCGA), 23(3): 159–??, June 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Cheng:2017:APL

[CL17] Siu-Wing Cheng and Man-Kit Lau. Adaptive point location in planar convex subdivisions. International Journal of Computational Geometry and Applications (IJCGA), 27(1–2):3–12, March/June 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Chen:2005:LPS

[CLL05] Yu-Shin Chen, D. T. Lee, and Chung-Shou Liao. Labeling points on a single line. International Journal of Computational Geometry and Applications (IJCGA), 15(3):261–??, June 2005. CODEN IJCAEV. ISSN 0218-1959.

Cardinal:2009:PGT

[CLLP09] Jean Cardinal, Martine Labbé, Stefan Langerman, and Belén Palop. Pricing geometric transportation networks. International Journal of Computational Geometry and Applications (IJCGA), 19(6):507–520, December 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Chazal:2007:NMB

[CLR07] Frederic Chazal, Andre Lieutier, and Jarek Rossignac. Normal-map between normalcompatible manifolds. International Journal of Computational Geometry and Applications (IJCGA), 17(5):403–421, October 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Chazal:2010:BMH

[CLRW10] Frédéric Chazal, André Lieutier, Jarek Rossignac, and Brian Whited. Ball-map: Homeomorphism between compatible surfaces. International Journal of Computational Geometry and Applications (IJCGA), 20(3):285–306, June 2010. CODEN IJCAEV. ISSN 0218-1959.

Chen:2003:TPA

[CLX03] Danny Z. Chen, Shuang Luan, and Jinhui Xu. Topological peeling and applications. International Journal of Computational Geometry and Applications (IJCGA), 13(2):135–??, April 2003. CODEN IJCAEV. ISSN 0218-1959.

Chen:2010:FMO

[CM10] Danny Z. Chen and Ewa Misiołek. Finding many optimal paths without growing any optimal path trees. International Journal of Computational Geometry and Applications (IJCGA), 20(4):449–469, August 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Chen:2011:FFS

[CM11] Danny Z. Chen and Ewa Misiołek. Free-form surface partition in 3-D. International Journal of Computational Geometry and Applications (IJCGA), 21(6):609–634, December 2011. CODEN IJCAEV. ISSN 0218-1959.

Cohen:2003:SMA

[CMO03] Jonathan Cohen, Dinesh Manocha, and Marc Olano. Successive mappings: An approach to polygonal mesh simplification with guaranteed error bounds. International Journal of Computational Geometry and Applications (IJCGA), 13(1):61– ??, February 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Chepoi:2010:PES

[CNTV10] Victor Chepoi, Karim Nouioua, Edouard Thiel, and Yann Vaxès. Pareto envelopes in simple polygons. International Journal of Computational Geometry and Applications (IJCGA), 20(6):707–721, December 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Cheong:2012:GEF

[CO12] Otfried Cheong and Yoshio Okamoto. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 22(1):1– ??, February 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Collins:2004:CSP

[Col04] Michael J. Collins. Covering a set of points with a minimum number of turns. International Journal of Computational Geometry and Applications (IJCGA), 14(1–2):105–??, April 2004. CODEN IJCAEV. ISSN 0218-1959.

Cazals:2005:DTG

[CP05] Frédéic Cazals and Marc Pouget. Differential topology and geometry of smooth embedded surfaces: Selected topics. International Journal of Computational Geometry and Applications (IJCGA), 15(5): 511–??, October 2005. CODEN IJCAEV. ISSN 0218-1959.

Chakraborty:2018:HTT

[CPRS18] Sourav Chakraborty, Rameshwar Pratap, Sasanka Roy, and Shubhangi Saraf. Helly-type theorems in property testing. International Journal of Computational Geometry and Applications (IJCGA), 28(4):365– 379, December 2018. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195918500115.

Christof:2001:DPT

[CR01] Thomas Christof and Gerhard Reinelt. Decomposition and parallelization techniques for enumerating the facets of combinatorial polytopes. International Journal of Computational Geometry and Applications (IJCGA), 11(4):423–437, 2001. CODEN IJCAEV. ISSN 0218-1959.

Chan:2006:GOP

[CS06] Timothy M. Chan and Bashir S.

Sadjad. Geometric optimization problems over sliding windows. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3): 145–??, June 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Czyzowicz:1999:IS

[CSU99] J. Czyzowicz, I. Stojmenovic, and J. Urrutia. Immobilizing a shape. International Journal of Computational Geometry and Applications (IJCGA), 9(2):181–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Chen:2005:GAD

[CSX05] Danny Z. Chen, Michiel Smid, and Bin Xu. Geometric algorithms for density-based data clustering. International Journal of Computational Geometry and Applications (IJCGA), 15 (3):239–??, June 2005. CODEN IJCAEV. ISSN 0218-1959.

Choi:1997:AES

[CSY97] Joonsoo Choi, Juergen Sellen, and Chee-Keng Yap. Approximate Euclidean shortest paths in 3-space. International Journal of Computational Geometry and Applications (IJCGA), 7(4):271–??, August 1997. CO-DEN IJCAEV. ISSN 0218-1959.

Chiang:1992:DTM

[CT92] Y.-J. Chiang and R. Tamassia. Dynamization of the trapezoid method for planar point location. International Journal of Computational Geometry and Applications (IJCGA), 2(??):311–333, 1992. CODEN IJCAEV. ISSN 0218-1959.

Chiang:1997:OSP

[CT97] Yi-Jen Chiang and Roberto Tamassia. Optimal shortest path and minimum — link path queries between two convex polygons inside a simple polygonal obstacle. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):85–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959.

Chew:2005:GEF

[CÜ05] Paul Chew and Alper Üngör.

Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 15(1):1–??, February 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Cheong:2007:IHP

[CVG^{+}07] Jae-Sook Cheong, A. Frank Van Der Stappen, Ken Goldberg, Mark H. Overmars, and Elon Rimon. Immobilizing hinged polygons. International Journal of Computational Geometry and Applications (IJCGA), 17 (1):45–69, February 2007. CO-DEN IJCAEV. ISSN 0218-1959.

Cheong:2001:RPD

[CvO01] Otfried Cheong and René van Oostrum. Reaching a polygon with directional uncertainty. International Journal of Computational Geometry and Applications (IJCGA), 11(2): 197–214, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Cheong:2011:RNN

[CVY11] Otfried Cheong, Antoine Vigneron, and Juyoung Yon. Reverse nearest neighbor queries in fixed dimension. International Journal of Computational Geometry and Applications (IJCGA), 21(2):179–188, April 2011. CODEN IJCAEV. ISSN 0218-1959.

Chen:2012:FSF

[CW12a] Danny Z. Chen and Haitao Wang. Fitting a step function to a point set with outliers based on simplicial thickness data structures. International Journal of Computational Geometry and Applications (IJCGA), 22(3):215–??, June 2012. CODEN IJCAEV. ISSN 0218-1959.

Chen:2012:LOL

[CW12b] Danny Z. Chen and Haitao Wang. Locating an obnoxious line among planar objects. International Journal of Computational Geometry and Applications (IJCGA), 22(5):391–??, October 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Chen:1998:FCH

[CWKC98] Wei Chen, Koichi Wada, Kimio Kawaguchi, and Danny Z. Chen. Finding the convex hull of discs in parallel. International Journal of Computational Geometry and Applications (IJCGA), 8(3):305–??, June 1998. CODEN IJCAEV. ISSN 0218-1959.

Chen:2002:ISA

[CWW02] Danny Z. Chen, Jie Wang, and Xiaodong Wu. Image segmentation with asteroidality/ tubularity and smoothness constraints. International Journal of Computational Geometry and Applications (IJCGA), 12 (5):413–??, October 2002. CO-DEN IJCAEV. ISSN 0218-1959.

Cheng:2008:PDD

[CWW08] Siu-Wing Cheng, Yajun Wang, and Zhuangzhi Wu. Provable dimension detection using principal component analysis. International Journal of Computational Geometry and Applications (IJCGA), 18(5):415–440, October 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Chang:2017:IAG

[CY17] Yi-Jun Chang and Hsu-Chun Yen. Improved algorithms for grid-unfolding orthogonal polyhedra. International Journal of Computational Geometry and Applications (IJCGA), 27(1–2): 33–56, March/June 2017. CO-DEN IJCAEV. ISSN 0218-1959.

DaFonseca:2011:FFP

[Da 11] Guilherme D. Da Fonseca. Fitting flats to points with outliers. International Journal of Computational Geometry and Applications (IJCGA), 21(5): 559–569, October 2011. CO-DEN IJCAEV. ISSN 0218-1959.

Dey:1992:GTT

[DB92] T. K. Dey and C. Bajaj. On good triangulations in three dimensions. International Journal of Computational Geometry and Applications (IJCGA), 2(1):75–95, 1992. CODEN IJ-CAEV. ISSN 0218-1959.

deBerg:2017:GEF

[dBDE17] Mark de Berg, Adrian Dumitrescu, and Khaled Elbassioni. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 27(1–2):1–2, March/June 2017. CODEN IJ-CAEV. ISSN 0218-1959.

deBerg:2019:FDH

[dBGR19] Mark de Berg, Ade Gunawan, and Marcel Roeloffzen. Faster DBSCAN and HDBSCAN in low-dimensional Euclidean spaces. International Journal of Computational Geometry and Applications (IJCGA), 29(1):21–47, March 2019. ISSN 0218-1959. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0218195919400028.

Diaz-Banez:2006:AVD

[DBGV06] J. M. Díaz-Báñez, F. Gómez, and I. Ventura. The anchored Voronoi diagram: Static, dynamic versions and applications. International Journal of Computational Geometry and Applications (IJCGA), 16(4): 315–??, August 2006. CODEN IJCAEV. ISSN 0218-1959.

Diaz-Banez:2003:LEA

[DBHM^{+}03] J. M. Díaz-Báñez, F. Hurtado, H. Meijer, D. Rappaport, and J. A. Sellarès. The largest empty annulus problem. International Journal of Computational Geometry and Applications (IJCGA), 13(4):317–??, August 2003. CODEN IJ-CAEV. ISSN 0218-1959.

deBerg:1997:SAN

[dBHOvK97] Mark de Berg, Dan Halperin, Mark Overmars, and Marc van Kreveld. Sparse arrangements and the number of views of polyhedral scenes. International Journal of Computational Geometry and Applications (IJCGA), 7(3):175–??, June 1997. CODEN IJCAEV. ISSN 0218-1959.

Diaz-Banez:2014:EF

[DBKU14] J. M. Díaz-Báñez, R. Klein, and J. Urrutia. Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 24(3):173–??, September 2014. CODEN IJ-CAEV. ISSN 0218-1959.

deBerg:2019:FDK

[dBLM^{+}19] Mark de Berg, Tim Leijsen, Aleksandar Markovic, André van Renssen, Marcel Roeloffzen, and Gerhard Woeginger. Fully-dynamic and kinetic conflict-free coloring of intervals with respect to points. International Journal of Computational Geometry and Applications (IJCGA), 29(1):49– 72, March 2019. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S021819591940003X.

deBerg:2002:GEF

[dBS02] Mark de Berg and Stefan Schirra. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2): 1–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

DeMeneses:2000:ESR

[DD00] C. N. De Meneses and C. C. De Souza. Exact solutions of rectangular partitions via integer programming. International Journal of Computational Geometry and Applications (IJCGA), 10(5):477–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

De:2013:AAV

[DDCN13] Minati De, Gautam K. Das, Paz Carmi, and Subhas C. Nandy. Approximation algorithms for a variant of discrete piercing set problem for unit disks. International Journal of Computational Geometry and Applications (IJCGA), 23(6):461–??, December 2013. CODEN IJCAEV. ISSN 0218-1959.

Devillers:2007:MVI

[DDE^{+}07] Olivier Devillers, Vida Dujmović, Hazel Everett, Samuel Hornus, Sue Whitesides, and Steve Wismath. Maintaining visibility information of planar point sets with a moving viewpoint. International Journal of Computational Geometry and Applications (IJCGA), 17(4): 297–304, August 2007. CODEN IJCAEV. ISSN 0218-1959.

Giacomo:2010:CPS

[DDL^{+}10] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, Henk Meijer, and Stephen K. Wismath. Constrained point-set embeddability of planar graphs. International Journal of Computational Geometry and Applications (IJCGA), 20(5):577– 600, October 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Dehkordi:2012:EOP

[DE12] Hooman Reisi Dehkordi and Peter Eades. Every outer-1-plane graph has a right angle crossing drawing. International Journal of Computational Geometry and Applications (IJCGA), 22(6):543–??, December 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Devillers:2003:MSC

[DEG^{+}03] Olivier Devillers, Regina Estkowski, Pierre-Marie Gandoin, Ferran Hurtado, Pedro Ramos, and Vera Sacristán. Minimal set of constraints for 2 d constrained Delaunay reconstruction. International Journal of Computational Geometry and Applications (IJCGA), 13(5): 391–??, October 2003. CODEN IJCAEV. ISSN 0218-1959.

Demaine:2005:SPS

[DEH^{+}05] Erik D. Demaine, Jeff Erickson, Ferran Hurtado, John Iacono, Stefan Langerman, Henk Meijer, Mark Overmars, and Sue Whitesides. Separating point sets in polygonal environments.

International Journal of Computational Geometry and Applications (IJCGA), 15(4):403– ??, August 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Devillers:1992:RYS

[Dev92] O. Devillers. Randomization yields simple O(n log^{∗} n) algorithms for difficult Ω(n) problems. International Journal of Computational Geometry and Applications (IJCGA), 2(1):97– 111, 1992. CODEN IJCAEV. ISSN 0218-1959.

Devillers:2002:DDT

[Dev02] Olivier Devillers. On deletion in Delaunay triangulations. International Journal of Computational Geometry and Applications (IJCGA), 12(3):193–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Dey:1997:EAD

[Dey97]  Tamal K. Dey.  Efficient

algorithms to detect nullhomologous cycles on 2manifolds. International Journal of Computational Geometry and Applications (IJCGA), 7 (3):167–??, June 1997. CODEN IJCAEV. ISSN 0218-1959.

dAmore:2018:LAD

[dF18] Fabrizio d’Amore and Paolo G. Franciosa. A low arithmeticdegree algorithm for computing proximity graphs. International Journal of Computational Geometry and Applications (IJCGA), 28(3):??, September 2018. ISSN 0218-1959.

daFonseca:2017:SCO

[dFdSdF17] Guilherme D. da Fonseca, Vinícius Gusmão Pereira de Sá, and Celina Miraglia Herrera de Figueiredo. Shifting coresets: Obtaining linear-time approximations for unit disk graphs and other geometric intersection graphs. International Journal of Computational Geometry and Applications (IJCGA), 27(4):255–??, December 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Durocher:2018:CDM

[DFL^{+}18] Stephane Durocher, Robert Fraser, Alexandre Leblanc, Jason Morrison, and Matthew Skala. On combinatorial depth measures. International Journal of Computational Geometry and Applications (IJCGA), 28(4):381–398, December 2018. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195918500127.

Das:2012:DUD

[DFLON12] Gautam K. Das, Robert Fraser, Alejandro Lóopez-Ortiz, and Bradford G. Nickerson. On the discrete unit disk cover problem. International Journal of Computational Geometry and Applications (IJCGA), 22(5): 407–??, October 2012. CODEN IJCAEV. ISSN 0218-1959.

Devillers:1998:DBP

[DG98] Olivier Devillers and Mordecai J. Golin. Dog bites postman: Point location in the moving Voronoi diagram and related problems. International Journal of Computational Geometry and Applications (IJCGA), 8(3):321–??, June 1998. CODEN IJCAEV. ISSN 0218-1959.

DAmore:1999:IPS

[DG99] F. D’Amore and R. Giaccio. Intersection problems on segments under boundary updates with application to persistent lists. International Journal of Computational Geometry and Applications (IJCGA), 9(6):553–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Devillers:2001:SB

[DG01] Olivier Devillers and Philippe Guigue. The shuffling buffer. International Journal of Computational Geometry and Applications (IJCGA), 11(5):555– 572, 2001. CODEN IJCAEV. ISSN 0218-1959.

Dakowicz:2003:EMS

[DG03] Maciej Dakowicz and Christopher Gold. Extracting meaningful slopes from terrain contours. International Journal of Computational Geometry and Applications (IJCGA), 13(4): 339–??, August 2003. CODEN IJCAEV. ISSN 0218-1959.

Berg:2013:CPP

[DG13] Mark De Berg and Dirk H. P.

Gerrits. Computing push plans for disk-shaped robots. International Journal of Computational Geometry and Applications (IJCGA), 23(1):29–??, February 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Dumitrescu:2016:LBD

[DG16] Adrian Dumitrescu and Anirban Ghosh. Lower bounds on the dilation of plane spanners. International Journal of Computational Geometry and Applications (IJCGA), 26(2):89–??, June 2016. CODEN IJCAEV. ISSN 0218-1959.

DiBattista:2000:DDA

[DGL^{+}00] G. Di Battista, A. Garg, G. Liotta, A. Parise, R. Tamassia, E. Tassinari, F. Vargiu, and L. Vismara. Drawing directed acyclic graphs: An experimental study. International Journal of Computational Geometry and Applications (IJCGA), 10 (6):623–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Das:2009:SCS

[DGN09] Sandip Das, Partha P. Goswami, and Subhas C. Nandy. Smallest color-spanning object revisited. International Journal of Computational Geometry and Applications (IJCGA), 19(5):457– 478, October 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Dey:2008:CPD

[DGRS08] Tamal K. Dey, Joachim Giesen, Edgar A. Ramos, and Bardia Sadri. Critical points of distance to an ϵ-sampling of a surface and flow-complex-based surface reconstruction. International Journal of Computational Geometry and Applications (IJCGA), 18(1-2):29–61, April 2008. CODEN IJCAEV. ISSN 0218-1959.

Dumitrescu:2013:CPC

[DH13] Adrian Dumitrescu and Masud Hasan. Cutting out polygons with a circular saw. International Journal of Computational Geometry and Applications (IJCGA), 23(2):127–??, April 2013. CODEN IJCAEV. ISSN 0218-1959.

Duque:2015:UBK

[DHT15] Frank Duque and Carlos Hidalgo-Toscano. An upper bound on the k-modem illumination problem. International Journal of Computational Geometry and Applications (IJCGA), 25(4):299–??, December 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Demaine:2010:GVU

[DIL10] Erik D. Demaine, John Iacono, and Stefan Langerman. Grid vertex-unfolding orthostacks. International Journal of Computational Geometry and Applications (IJCGA), 20(3):245– 254, June 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Devillers:1999:OLB

[DK99] O. Devillers and M. J. Katz.

Optimal line bipartitions of point sets. International Journal of Computational Geometry and Applications (IJCGA), 9(1):39–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Durocher:2006:SCS

[DK06] Stephane Durocher and David Kirkpatrick. The Steiner centre of a set of points: Stability, eccentricity, and applications to mobile facility location. International Journal of Computational Geometry and Applications (IJCGA), 16(4):345– ??, August 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Durocher:2008:BVA

[DK08] Stephane Durocher and David Kirkpatrick. Bounded-velocity approximation of mobile Euclidean 2-centres. International Journal of Computational Geometry and Applications (IJCGA), 18(3):161–183, June 2008. CODEN IJCAEV. ISSN 0218-1959.

DeBerg:2012:OBS

[DK12] Mark De Berg and Amirali Khosravi. Optimal binary space partitions for segments in the plane. International Journal of Computational Geometry and Applications (IJCGA), 22(3): 187–??, June 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Dinitz:2009:GRP

[DKK09] Yefim Dinitz, Matthew J. Katz, and Roi Krakovski. Guarding rectangular partitions. International Journal of Computational Geometry and Applications (IJCGA), 19(6):579–594, December 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Durocher:2023:BCS

[DKMM23] Stephane Durocher, J. Mark Keil, Saeed Mehrabi, and Debajyoti Mondal. Bottleneck convex subsets: Finding k large convex sets in a point set. International Journal of Computational Geometry and Applications (IJCGA), 33(1–2):25– 41, March–June 2023. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195922410035.

Dehne:2005:MVR

[DKS05] Frank Dehne, Rolf Klein, and Raimund Seidel. Maximizing a Voronoi region: the convex case. International Journal of Computational Geometry and Applications (IJCGA), 15 (5):463–??, October 2005. CO-DEN IJCAEV. ISSN 0218-1959.

Daescu:2006:CPL

[DL06] Ovidiu Daescu and Jun Luo. Cutting out polygons with lines and rays. International Journal of Computational Geometry and Applications (IJCGA), 16 (2–3):227–??, June 2006. CO-DEN IJCAEV. ISSN 0218-1959.

DiGiacomo:2007:SEO

[DL07] Emilio Di Giacomo and Giuseppe Liotta. Simultaneous embedding of outerplanar graphs, paths, and cycles. International Journal of Computational Geometry and Applications (IJCGA), 17(2):139–160, April 2007. CODEN IJCAEV. ISSN 0218-1959.

Durocher:2013:RNS

[DLMS13] Stephane Durocher, Alexandre Leblanc, Jason Morrison, and Matthew Skala. Robust nonparametric simplification of polygonal chains. International Journal of Computational Geometry and Applications (IJCGA), 23(6):427–??, December 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Dror:2006:TSM

[DLOP06] Moshe Dror, Yusin Lee, James B. Orlin, and Valentin Polishchuk. The TSP and the sum of its marginal values. International Journal of Computational Geometry and Applications (IJCGA), 16(4):333– ??, August 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Doddi:2002:PSL

[DMM02] Srinivas R. Doddi, Madhav V.

Marathe, and Bernard M. E. Moret. Point set labeling with specified positions. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2):29–??, 2002. CODEN IJCAEV. ISSN 0218-1959. URL http://www.

                                        cs.unm.edu/~moret/papers.
                                        html.

Santos:2011:EDL

[DMMH11] Thiago R. Dos Santos, Hans-Peter Meinzer, and Lena Maier-Hein. Extending the doubly linked face list for the representation of 2-pseudomanifolds and 2-manifolds with boundaries. International Journal of Computational Geometry and Applications (IJCGA), 21(4): 467–494, August 2011. CODEN IJCAEV. ISSN 0218-1959.

DeBerg:1998:CAT

[DMOW98] Mark De Berg, Henk Meijer, Mark Overmars, and Gordon Wilfong. Computing the angularity tolerance. International Journal of Computational Geometry and Applications (IJCGA), 8(4):467–??, August 1998. CODEN IJ-CAEV. ISSN 0218-1959.

DeBerg:2010:OBR

[DMS10] Mark De Berg, Elena Mumford, and Bettina Speckmann. Optimal BSPs and rectilinear cartograms. International Journal of Computational Geometry and Applications (IJCGA), 20 (2):203–222, April 2010. CO-DEN IJCAEV. ISSN 0218-1959.

Das:1997:FAC

[DN97] Gautam Das and Giri Narasimhan. A fast algorithm for constructing sparse Euclidean spanners. International Journal of Computational Geometry and Applications (IJCGA), 7(4):297– ??, August 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Devillers:2018:WPP

[DN18] Olivier Devillers and Louis Noizet. Walking in a planar Poisson–Delaunay triangulation: Shortcuts in the Voronoi path. International Journal of Computational Geometry and Applications (IJCGA), 28(3): ??, September 2018. ISSN 0218-1959.

Dillard:2009:TGT

[DNW^{+}09] Scott E. Dillard, Vijay Natarajan, Gunther H. Weber, Valerio Pascucci, and Bernd Hamann. Topology-guided tessellation of quadratic elements. International Journal of Computational Geometry and Applications (IJCGA), 19(2):195–211, April 2009. CODEN IJCAEV. ISSN 0218-1959.

Demaine:2000:CGC

[DO00] E. D. Demaine and J. O’Rourke.

Computational geometry column 37. International Journal of Computational Geometry and Applications (IJCGA), 10 (1):103–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Dumitrescu:2002:PCP

[DP02] Adrian Dumitrescu and János Pach. Partitioning colored point sets into monochromatic parts. International Journal of Computational Geometry and Applications (IJCGA), 12(5): 401–??, October 2002. CODEN IJCAEV. ISSN 0218-1959.

Devillers:2003:CSP

[DP03] Olivier Devillers and Franco P. Preparata. Culling a set of points for roundness or cylindricity evaluations. International Journal of Computational Geometry and Applications (IJCGA), 13(3):231–??, June 2003. CODEN IJCAEV. ISSN 0218-1959.

Devillers:2002:CRE

[DR02] Olivier Devillers and Pedro A. Ramos. Computing roundness is easy if the set is almost round. International Journal of Computational Geometry and Applications (IJCGA), 12(3): 229–??, 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Dutta:2020:ASP

[DR20] Binayak Dutta and Sasanka Roy. Approximate shortest paths in polygons with violations. International Journal of Computational Geometry and Applications (IJCGA), 30 (01):79–95, March 2020. ISSN 0218-1959. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0218195920500041.

Dey:2002:FRC

[DW02] Tamal K. Dey and Rephael Wenger. Fast reconstruction of curves with sharp corners. International Journal of Computational Geometry and Applications (IJCGA), 12(5):353– ??, October 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Ebbers-Baumann:2007:EPS

[EBGK^{+}07] Annette Ebbers-Baumann, Ansgar Grüne, Rolf Klein, Marek Karpinski, Christian Knauer, and Andrzej Lingas. Embedding point sets into plane graphs of small dilation. International Journal of Computational Geometry and Applications (IJCGA), 17(3):201–230, June 2007. CODEN IJCAEV. ISSN 0218-1959.

Estivill-Castro:2015:IFC

[EC15] Vladimir Estivill-Castro. Is it FPT to cover points with tours on minimum number of bends (errata)? International Journal of Computational Geometry and Applications (IJCGA), 25(1):11–??, March 2015. CO-DEN IJCAEV. ISSN 0218-1959. See [ECHS11].

Estivill-Castro:2011:FAM

[ECHS11] Vladimir Estivill-Castro, Apichat Heednacram, and Francis Suraweera. FPT-algorithms for minimum-bends tours. International Journal of Computational Geometry and Applications (IJCGA), 21(2):189–213, April 2011. CODEN IJCAEV. ISSN 0218-1959. See errata [EC15].

Elbassioni:2011:FSC

[EEM11] Khaled Elbassioni, Amr Elmasry, and Kazuhisa Makino.

Finding simplices containing the origin in two and three dimensions. International Journal of Computational Geometry and Applications (IJCGA), 21(5):495–506, October 2011. CODEN IJCAEV. ISSN 0218-1959.

Eisenbrand:2007:PTN

[EFK^{+}07] Friedrich Eisenbrand, Stefan Funke, Andreas Karrenbauer, Joachim Reichel, and Elmar Schömer. Packing a truck — now with a twist! International Journal of Computational Geometry and Applications (IJCGA), 17(5):505–527, October 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Eisenbrand:2008:EAS

[EFKM08] Friedrich Eisenbrand, Stefan Funke, Andreas Karrenbauer, and Domagoj Matijevic. Energy-aware stage illumination. International Journal of Computational Geometry and Applications (IJCGA), 18(1-2): 107–129, April 2008. CODEN IJCAEV. ISSN 0218-1959.

Emiris:2013:OBO

[EFKP13] Ioannis Z. Emiris, Vissarion Fisikopoulos, Christos Konaxis, and Luis Peñaranda. An oraclebased, output-sensitive algorithm for projections of resultant polytopes. International Journal of Computational Geometry and Applications (IJCGA), 23(4–5):397–??, 2013. CODEN IJCAEV. ISSN 0218-1959.

Efrat:2008:GEF

[Efr08] Alon Efrat. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 18 (1-2):1–2, April 2008. CODEN IJCAEV. ISSN 0218-1959.

Elbassioni:2009:AAE

[EFS09] Khaled Elbassioni, Aleksei V.

Fishkin, and René Sitters. Approximation algorithms for the Euclidean traveling salesman problem with discrete and continuous neighborhoods. International Journal of Computational Geometry and Applications (IJCGA), 19(2):173–193, April 2009. CODEN IJCAEV. ISSN 0218-1959.

Eppstein:2008:SQD

[EGS08] David Eppstein, Michael T.

Goodrich, and Jonathan Z. Sun. Skip quadtrees: Dynamic data structures for multidimensional point sets. International Journal of Computational Geometry and Applications (IJCGA), 18(1-2):131– 160, April 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Eder:2019:WVD

[EH19] Günther Eder and Martin Held.

Weighted Voronoi diagrams in the maximum norm. International Journal of Computational Geometry and Applications (IJCGA), 29(03):239– 250, September 2019. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195919500079.

Eder:2018:MMV

[EHP18] Günther Eder, Martin Held, and Peter Palfrader. Min-/ max-volume roofs induced by bisector graphs of polygonal footprints of buildings. International Journal of Computational Geometry and Applications (IJCGA), 28(4):309– 340, December 2018. CO-DEN IJCAEV. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195918500097.

Eder:2019:RGT

[EHP19] Günther Eder, Martin Held, and Peter Palfrader. Recognizing geometric trees as positively weighted straight skeletons and reconstructing their input. International Journal of Computational Geometry and Applications (IJCGA), 29(03): 251–267, September 2019. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195919500080.

Everett:2007:ESV

[ELPZ07] Hazel Everett, Sylvain Lazard, Sylvain Petitjean, and Linqiao Zhang. On the expected size of the 2D visibility complex. International Journal of Computational Geometry and Applications (IJCGA), 17(4):361– 381, August 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Emiris:1998:CIC

[Emi98] Ioannis Z. Emiris. A complete implementation for computing general dimensional convex hulls. International Journal of Computational Geometry and Applications (IJCGA), 8(2):223–??, April 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Evans:1998:SCM

[EMM98] Guy Evans, Alan Middleditch, and Nick Miles. Stable computation of the 2D medial axis transform. International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):577–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Eppstein:1997:FCP

[Epp97] David Eppstein. Faster circle packing with application to nonobtuse triangulation. International Journal of Computational Geometry and Applications (IJCGA), 7(5):485–??, October 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Edelsbrunner:1997:TTS

[ES97] Herbert Edelsbrunner and Nimish R. Shah. Triangulating topological spaces. International Journal of Computational Geometry and Applications (IJCGA), 7(4):365– ??, August 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Ensz:1998:IMG

[ESG98] Mark T. Ensz, Duane W. Storti, and Mark A. Ganter. Implicit methods for geometry creation. International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):509–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

ElOraiby:2011:CTS

[ESS11] Wael El Oraiby, Dominique Schmitt, and Jean-Claude Spehner. Centroid triangulations from k-sets. International Journal of Computational Geometry and Applications (IJCGA), 21(6):635–659, December 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Emiris:2008:PEV

[ETT08] Ioannis Z. Emiris, Elias P. Tsigaridas, and George M. Tzoumas. The predicates for the exact Voronoi diagram of ellipses under the Euclidean metric. International Journal of Computational Geometry and Applications (IJCGA), 18(6): 567–597, December 2008. CO-DEN IJCAEV. ISSN 0218-1959.

Eppstein:2015:IGM

[EvKSS15] David Eppstein, Marc van Kreveld, Bettina Speckmann, and Frank Staals. Improved grid map layout by point set matching. International Journal of Computational Geometry and Applications (IJCGA), 25(2):??, June 2015. CODEN IJCAEV. ISSN 0218-1959.

Edelsbrunner:2000:ASG

[EW00] H. Edelsbrunner and R. Waupotitsch.

Adaptive simplicial grids from cross-sections of monotone complexes. International Journal of Computational Geometry and Applications (IJCGA), 10 (3):267–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Fischer:2004:SEB

[FG04] Kaspar Fischer and Bernd Gärtner. The smallest enclosing ball of balls: Combinatorial structure and algorithms. International Journal of Computational Geometry and Applications (IJCGA), 14(4–5):341– ??, October 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Frahling:2008:SDD

[FIS08] Gereon Frahling, Piotr Indyk, and Christian Sohler. Sampling in dynamic data streams and applications. International Journal of Computational Geometry and Applications (IJCGA), 18(1-2):3–28, April 2008. CODEN IJCAEV. ISSN 0218-1959.

Fekete:2018:CFC

[FK18] Sándor P. Fekete and Phillip Keldenich. Conflict-free coloring of intersection graphs. International Journal of Computational Geometry and Applications (IJCGA), 28(3):??, September 2018. ISSN 0218-1959.

Fasy:2022:RGS

[FKMW22] Brittany Terese Fasy, Rafal Komendarczyk, Sushovan Majhi, and Carola Wenk. On the reconstruction of geodesic subspaces of R^{N}. International Journal of Computational Geometry and Applications (IJCGA), 32(1–2):91– 117, March–June 2022. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195922500066.

Froese:2017:FPG

[FKNN17] Vincent Froese, Iyad Kanj, André Nichterlein, and Rolf Niedermeier. Finding points in general position. International Journal of Computational Geometry and Applications (IJCGA), 27(4):277–??, December 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Fleischer:2006:GEF

[Fle06] Rudolf Fleischer. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3):95–??, June 2006. CODEN IJCAEV. ISSN 0218-1959.

Fekete:2018:UGP

[FLMS18] Sándor P. Fekete, Qian Li, Joseph S. B. Mitchell, and Christian Scheffer. Universal guard problems. International Journal of Computational Geometry and Applications (IJCGA), 28(2):??, June 2018. ISSN 0218-1959.

Fekete:1997:HDS

[FM97] S. P. Fekete and J. S. B.

Mitchell. Histogram decomposition and stereolithography. International Journal of Computational Geometry and Applications (IJCGA), ??(??):??, 1997. CODEN IJCAEV. ISSN 0218-1959.

Fekete:1999:RBV

[FM99] S. P. Fekete and H. Meijer. Rectangle and box visibility graphs in 3D. International Journal of Computational Geometry and Applications (IJCGA), 9(1):1–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Fekete:2001:TDL

[FM01] Sándor P. Fekete and Joseph S. B. Mitchell. Terrain decomposition and layered manufacturing. International Journal of Computational Geometry and Applications (IJCGA), 11 (6):647–??, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Fabila-Monroy:2014:NNO

[FMHT14] Ruy Fabila-Monroy, Clemens Huemer, and Eulàlia Tramuns. Note on the number of obtuse angles in point sets. International Journal of Computational Geometry and Applications (IJCGA), 24(3):177–??, September 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Funke:2005:FPR

[FMR05] Stefan Funke, Theocharis Malamatos, and Rahul Ray. Finding planar regions in a terrain – in practice and with a guarantee. International Journal of Computational Geometry and Applications (IJCGA), 15 (4):379–??, August 2005. CO-DEN IJCAEV. ISSN 0218-1959.

Falconer:2005:MLR

[FN05] Sean M. Falconer and Bradford G. Nickerson. On multilevel k-ranges for range search. International Journal of Computational Geometry and Applications (IJCGA), 15(6):565–??, December 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Freitag:2000:CBA

[FOG00] L. A. Freitag and C. Ollivier-Gooch. A cost/benefit analysis of simplicial mesh improvement techniques as measured by solution efficiency. International Journal of Computational Geometry and Applications (IJCGA), 10(4):361–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Fortune:1995:NSA

[For95] S. J. Fortune. Numerical stability of algorithms for 2D Delaunay triangulations. International Journal of Computational Geometry and Applications (IJCGA), 5(??):193–213, 1995. CODEN IJCAEV. ISSN 0218-1959.

Fortune:1997:EF

[For97] Steven Fortune. Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 7(4):269–??, August 1997. CO-DEN IJCAEV. ISSN 0218-1959.

Fu:2008:MDW

[FOX08] Bin Fu, Sorinel A. Oprisan, and Lizhe Xu. Multi-directional width-bounded geometric separator and protein folding. International Journal of Computational Geometry and Applications (IJCGA), 18(5):389–413, October 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Fjallstrom:1998:ERS

[FPNZ98] Per-Olof Fjällström, Jan Petersson, Larsgunnar Nilsson, and Zhi-Hua Zhong. Evaluation of range searching methods for contact searching in mechanical engineering. International Journal of Computational Geometry and Applications (IJCGA), 8(1):67–??, February 1998. CODEN IJ-CAEV. ISSN 0218-1959.

Farouki:1998:SPC

[FR98] Rida T. Farouki and Rajesh Ramamurthy. Specifiedprecision computation of curve/ curve bisectors. International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):599–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Frati:2008:MAP

[Fra08] Fabrizio Frati. On minimum area planar upward drawings of directed trees and other families of directed acyclic graphs. International Journal of Computational Geometry and Applications (IJCGA), 18(3):251–271, June 2008. CODEN IJCAEV. ISSN 0218-1959.

Frahling:2008:FKM

[FS08] Gereon Frahling and Christian Sohler. A fast k-means implementation using coresets. International Journal of Computational Geometry and Applications (IJCGA), 18(6):605–625, December 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Follert:1997:CLE

[FSS^{+}97] Frank Follert, Elmar Schömer, Jürgen Sellen, Michiel Smid, and Christian Thiel. Computing a largest empty anchored cylinder, and related problems. International Journal of Computational Geometry and Applications (IJCGA), 7(6):563–??, December 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Fink:2003:PSV

[FW03] Eugene Fink and Derick Wood. Planar strong visibility. International Journal of Computational Geometry and Applications (IJCGA), 13(2):173–??, April 2003. CODEN IJCAEV. ISSN 0218-1959.

Gavrilova:2005:GEF

[Gav05] Marina L. Gavrilova. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 15(2):99–??, April 2005. CODEN IJCAEV. ISSN 0218-1959.

Gavrilova:2009:DSM

[Gav09a] Marina L. Gavrilova. Determining a set of maximum inscribed rectangles for label placement in a region. International Journal of Computational Geometry and Applications (IJCGA), 19 (4):341–356, August 2009. CO-DEN IJCAEV. ISSN 0218-1959.

Gavrilova:2009:ESC

[Gav09b] Marina L. Gavrilova. An explicit solution for computing the vertices of the Euclidean d-dimensional Voronoi diagram of spheres in a floating-point arithmetic. International Journal of Computational Geometry and Applications (IJCGA), 19(5):415–424, October 2009. CODEN IJCAEV. ISSN 0218-1959.

Grossman:2013:MCC

[GBRT13] P. A. Grossman, M. Brazil, J. H. Rubinstein, and D. A. Thomas. Minimal curvatureconstrained paths in the plane with a constraint on arcs with opposite orientations. International Journal of Computational Geometry and Applications (IJCGA), 23(3):171–??, June 2013. CODEN IJCAEV. ISSN 0218-1959.

Ganley:1997:ICO

[GC97] Joseph L. Ganley and J. P. Cohoon. Improved computation of optimal rectilinear Steiner minimal trees. International Journal of Computational Geometry and Applications (IJCGA), 7(5):457–??, October 1997. CO-DEN IJCAEV. ISSN 0218-1959.

Gupta:2021:SMO

[GGBK21] Prashant Gupta, Yiran Guo, Narasimha Boddeti, and Bala Krishnamoorthy. SFCDecomp: Multicriteria optimized tool path planning in 3D printing using space-filling curve based domain decomposition. International Journal of Computational Geometry and Applications (IJCGA), 31(4):193– 220, December 2021. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195921500096.

Guibas:1998:PAP

[GHH^{+}98] Leonidas J. Guibas, Dan Halperin, Hirohisa Hirukawa, Jean-Claude Latombe, and Randall H. Wilson. Polyhedral assembly partitioning using maximally covered cells in arrangements of convex polytopes. International Journal of Computational Geometry and Applications (IJCGA), 8(2): 179–??, April 1998. CODEN IJCAEV. ISSN 0218-1959.

Galtier:2003:SEF

[GHN^{+}03] Jerôme Galtier, Ferran Hurtado, Marc Noy, Stéphane Pérennes, and Jorge Urrutia. Simultaneous edge flipping in triangulations. International Journal of Computational Geometry and Applications (IJCGA), 13(2):113–??, April 2003. CODEN IJCAEV. ISSN 0218-1959.

Gupta:2012:CSL

[GIPR12] Rishi Gupta, Piotr Indyk, Eric Price, and Yaron Rachlin. Compressive sensing with local geometric features. International Journal of Computational Geometry and Applications (IJCGA), 22(4):365– ??, August 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Gitik:2021:VDD

[GJ21] Rivka Gitik and Leo Joskowicz.

Voronoi diagram and Delaunay triangulation with independent and dependent geometric uncertainties. International Journal of Computational Geometry and Applications (IJCGA), 31 (2–3):75–121, June–September 2021. CODEN IJCAEV. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195921500059.

Giesen:2003:SFD

[GJS03] Joachim Giesen, Matthias John, and Michel Stöcklin. Symmetry of flow diagrams derived from weighted points in the plane. International Journal of Computational Geometry and Applications (IJCGA), 13 (4):327–??, August 2003. CO-DEN IJCAEV. ISSN 0218-1959.

Gupta:2009:ENI

[GJS09] Prosenjit Gupta, Ravi Janardan, and Michiel Smid. Efficient non-intersection queries on aggregated geometric data. International Journal of Computational Geometry and Applications (IJCGA), 19(6):479– 506, December 2009. CODEN IJCAEV. ISSN 0218-1959.

Gupta:1997:REP

[GJSD97] Prosenjit Gupta, Ravi Janardan, Michiel Smid, and Bhaskar Dasgupta. The rectangle enclosure and point-dominance problems revisited. International Journal of Computational Geometry and Applications (IJCGA), 7(5):437–??, October 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Gupta:2020:ETP

[GK20] Prashant Gupta and Bala Krishnamoorthy. Euler transformation of polyhedral complexes. International Journal of Computational Geometry and Applications (IJCGA), 30(3–4):183–211, September– December 2020. ISSN 0218-

                                        1959.  URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0218195920500090.

Giannopoulos:2010:CGM

[GKK^{+}10] Panos Giannopoulos, Rolf Klein, Christian Knauer, Martin Kutz, and Dániel Marx. Computing geometric minimumdilation graphs is NP-hard. International Journal of Computational Geometry and Applications (IJCGA), 20(2):147–173, April 2010. CODEN IJCAEV. ISSN 0218-1959.

Glozman:1999:CDR

[GKS99] A. Glozman, K. Kedem, and G. Shpitalnik. Computing a double-ray center for a planar point set. International Journal of Computational Geometry and Applications (IJCGA), 9(2):109–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Guibas:1999:VBP

[GLL^{+}99] Leonidas J. Guibas, Jean-Claude Latombe, Steven M. Lavalle, David Lin, and Rajeev Motwani. A visibility-based pursuit-evasion problem. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):471–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Gray:2010:SIT

[GLS10] Chris Gray, Maarten Löffler, and Rodrigo I. Silveira. Smoothing imprecise 1.5D terrains. International Journal of Computational Geometry and Applications (IJCGA), 20(4):381– 414, August 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Guibas:1998:RAD

[GM98] Leonidas J. Guibas and David H. Marimont. Rounding arrangements dynamically. International Journal of Computational Geometry and Applications (IJCGA), 8(2):157–??, April 1998. CODEN IJCAEV. ISSN 0218-1959.

Goldwasser:1999:CMA

[GM99] Michael H. Goldwasser and Rajeev Motwani. Complexity measures for assembly sequences. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):371– ??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Gao:2006:GEF

[GM06] Xiao-Shan Gao and Dominique Michelucci. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6): 377–??, December 2006. CO-DEN IJCAEV. ISSN 0218-1959.

Gudmundsson:2019:FFD

[GMMW19] Joachim Gudmundsson, Majid Mirzanezhad, Ali Mohades, and Carola Wenk. Fast Fréchet distance between curves with long edges. International Journal of Computational Geometry and Applications (IJCGA), 29 (02):161–187, June 2019. ISSN 0218-1959. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0218195919500043.

Grove:1999:OCM

[GMV99] E. F. Grove, T. M. Murali, and J. S. Vitter. The object complexity model for hidden-surface removal. International Journal of Computational Geometry and Applications (IJCGA), 9(2):207–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Gosselin:2011:CCD

[GOG11] Serge Gosselin and Carl Ollivier-Gooch. Constructing constrained Delaunay tetrahedralizations of volumes bounded by piecewise smooth surfaces. International Journal of Computational Geometry and Applications (IJCGA), 21(5):571–594, October 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Goodrich:1998:IRS

[Goo98] Michael T. Goodrich. An improved ray shooting method for constructive solid geometry models via tree contraction. International Journal of Computational Geometry and Applications (IJCGA), 8(1):1–??, February 1998. CODEN IJ-CAEV. ISSN 0218-1959.

Garg:2003:AEO

[GR03a] Ashim Garg and Adrian Rusu.

Area-efficient order-preserving planar straight-line drawings of ordered trees. International Journal of Computational Geometry and Applications (IJCGA), 13(6):487–??, December 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Gavrilova:2003:CDO

[GR03b] Marina L. Gavrilova and Jon Rokne. Collision detection optimization in a multi-particle system. International Journal of Computational Geometry and Applications (IJCGA), 13 (4):279–??, August 2003. CO-DEN IJCAEV. ISSN 0218-1959.

Gabriely:2010:CCM

[GR10] Yoav Gabriely and Elon Rimon. Competitive complexity of mobile robot on-line motion planning problems. International Journal of Computational Geometry and Applications (IJCGA), 20(3):255–283, June 2010. CODEN IJCAEV. ISSN 0218-1959.

Giesen:2008:MAA

[GRS08] Joachim Giesen, Edgar A. Ramos, and Bardia Sadri. Medial axis approximation and unstable flow complex. International Journal of Computational Geometry and Applications (IJCGA), 18(6):533–565, December 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Grechuk:2020:CCC

[GSa20] Bogdan Grechuk and Sittichoke Som-am. A convex cover for closed unit curves has area at least 0.0975. International Journal of Computational Geometry and Applications (IJCGA), 30(02):121– 139, June 2020. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195920500065.

Gagie:2023:RW

[GSS23] Travis Gagie, Mozhgan Saeidi, and Allan Sapucaia. Ruler wrapping. International Journal of Computational Geometry and Applications (IJCGA), 33 (1–2):3–12, March–June 2023. ISSN 0218-1959. URL https:

//www.worldscientific.com/
doi/10.1142/S0218195922410011.

Gorke:2008:CCV

[GSW08] Robert Görke, Chan-Su Shin, and Alexander Wolff. Constructing the city Voronoi diagram faster. International Journal of Computational Geometry and Applications (IJCGA), 18(4):275–294, August 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Guo:2011:GCA

[GSZ11] Zeyu Guo, He Sun, and Hong Zhu. Greedy construction of 2-approximate minimum Manhattan networks. International Journal of Computational Geometry and Applications (IJCGA), 21(3):331–350, June 2011. CODEN IJCAEV. ISSN 0218-1959.

Guha:2005:JSG

[Guh05] Sumanta Guha. Joint separation of geometric clusters and the extreme irregularities of regular polyhedra. International Journal of Computational Geometry and Applications (IJCGA), 15(5):491–??, October 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Guigues:2022:ALE

[Gui22] Vincent Guigues. Algorithms and a library for the exact computation of the cumulative distribution function of the Euclidean distance between a point and a random variable uniformly distributed in disks, balls, or polygones and application to probabilistic seismic hazard analysis. International Journal of Computational Geometry and Applications (IJCGA), 32(3–4): 119–174, September–December 2022. CODEN IJCAEV. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195922500091.

Goldman:2004:UIE

[GW04] Ron Goldman and Wenping Wang. Using invariants to extract geometric characteristics of conic sections from rational quadratic parameterizations. International Journal of Computational Geometry and Applications (IJCGA), 14(3): 161–??, June 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Harrington:2007:OVD

[HDY07] Paul Harrington, Colm Ó. Dúnlaing, and Chee K. Yap. Optimal Voronoi diagram construction with n convex sites in three dimensions. International Journal of Computational Geometry and Applications (IJCGA), 17(6):555–593, December 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Hershberger:2001:GEF

[Her01]  John Hershberger.  Guest

Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 11(3):243–244, 2001. CODEN IJCAEV. ISSN 0218-1959.

Heimlich:2008:BAS

[HH08] Martin Heimlich and Martin Held. Biarc approximation, simplification and smoothing of polygonal curves by means of Voronoi-based tolerance bands. International Journal of Computational Geometry and Applications (IJCGA), 18(3):221– 250, June 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Huber:2012:FSS

[HH12] Stefan Huber and Martin Held.

A fast straight-skeleton algorithm based on generalized motorcycle graphs. International Journal of Computational Geometry and Applications (IJCGA), 22(5):471–??, October 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Huber:2014:TPW

[HHMK14] Stefan Huber, Martin Held, Peter Meerwald, and Roland Kwitt. Topology-preserving watermarking of vector graphics. International Journal of Computational Geometry and Applications (IJCGA), 24(1): 61–??, March 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Hiyoshi:2008:SCN

[Hiy08]  Hisamoto Hiyoshi.  Stable

computation of natural neighbor interpolation. International Journal of Computational Geometry and Applications (IJCGA), 18(4):321–341, August 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Hert:1997:PCR

[HL97] Susan Hert and Vladimir Lumelsky. Planar curve routing for tethered-robot motion planning. International Journal of Computational Geometry and Applications (IJCGA), 7 (3):225–??, June 1997. CODEN IJCAEV. ISSN 0218-1959.

Hert:1998:PAD

[HL98] Susan Hert and Vladimir Lumelsky. Polygon area decomposition for multiple-robot workspace division. International Journal of Computational Geometry and Applications (IJCGA), 8(4):437–??, August 1998. CODEN IJ-CAEV. ISSN 0218-1959.

Halperin:2004:CPA

[HL04] Dan Halperin and Eran Leiserowitz. Controlled perturbation for arrangements of circles. International Journal of Computational Geometry and Applications (IJCGA), 14(4–5):277– ??, October 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Hsu:1999:PPE

[HLM99] David Hsu, Jean-Claude Latombe, and Rajeev Motwani. Path planning in expansive configuration spaces. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):495–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Hurtado:2014:TVM

[HLM^{+}14] Ferran Hurtado, Maarten Löffler, Inês Matos, Vera Sacristán, Maria Saumell, Rodrigo I. Silveira, and Frank Staals. Terrain visibility with multiple viewpoints. International Journal of Computational Geometry and Applications (IJCGA), 24(4):275–??, December 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Hurtado:2013:PDH

[HLW13] Ferran Hurtado, Giuseppe Liotta, and David R. Wood. Proximity drawings of highdegree trees. International Journal of Computational Geometry and Applications (IJCGA), 23(3):213–??, June 2013. CO-DEN IJCAEV. ISSN 0218-1959.

Hong:2011:GEF

[HN11] Seok-Hee Hong and Hiroshi Nagamochi. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 21 (3):251–252, June 2011. CO-DEN IJCAEV. ISSN 0218-1959.

Hong:2018:GEF

[Hon18] Seok-Hee Hong. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 28 (2):??, June 2018. CODEN IJ-CAEV. ISSN 0218-1959.

Huang:2022:RPR

[HQYD22] Jiawei Huang, Ruizhe Qin, Fan Yang, and Hu Ding. Random projection and recovery for high dimensional optimization with arbitrary outliers. International Journal of Computational Geometry and Applications (IJCGA), 32(3–4):201–225, September– December 2022. ISSN 0218-

                                        1959.  URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0218195922500078.

Hanniel:2007:PVC

[HREK07] I. Hanniel, M. Ramanathan, G. Elber, and M.-S. Kim. Precise Voronoi cell extraction of free-form planar piecewise C^{1}-continuous closed rational curves. International Journal of Computational Geometry and Applications (IJCGA), 17(5): 453–486, October 2007. CO-DEN IJCAEV. ISSN 0218-1959.

Ha:2002:EAR

[HS02] Jong-Sung Ha and Sung-Yong Shin. Edge advancing rules for intersecting spherical convex polygons. International Journal of Computational Geometry and Applications (IJCGA), 12 (3):207–??, 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Hong:1998:TBC

[HSKK98] Mu Hong, Thomas W. Sederberg, Krzysztof S. Klimaszewski, and Kazufumi Kaneda. Triangulation of branching contours using area minimization. International Journal of Computational Geometry and Applications (IJCGA), 8(4):389– ??, August 1998. CODEN IJ-CAEV. ISSN 0218-1959.

Hurtado:2005:RBS

[HSS05] Ferran Hurtado, Carlos Seara, and Saurabh Sethia. Red-blue separability problems in 3D. International Journal of Computational Geometry and Applications (IJCGA), 15(2):167–??, April 2005. CODEN IJCAEV. ISSN 0218-1959.

Hoffmann:1991:EES

[HV91] C. M. Hoffmann and P. J. Vermeer. Eliminating extraneous solutions in curve and surface operations. International Journal of Computational Geometry and Applications (IJCGA), 1(??):47–66, 1991. CODEN IJ-CAEV. ISSN 0218-1959.

Hurtado:2012:GEF

[HV12] Ferran Hurtado and Marc Van Kreveld. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 22(4):

277–??, August 2012. CODEN IJCAEV. ISSN 0218-1959.

Hummel:2018:LIM

[HYSC18] Michelle Hatch Hummel, Bihua Yu, Carlos Simmerling, and Evangelos A. Coutsias. Laguerre-intersection method for implicit solvation. International Journal of Computational Geometry and Applications (IJCGA), 28(1):1–??, March 2018. CODEN IJCAEV. ISSN 0218-1959.

Huang:2024:EAC

[HZX24] Ziyun Huang, Yongding Zhu, and Jinhui Xu. An efficient algorithm for the 2central path problem. International Journal of Computational Geometry and Applications (IJCGA), 34(1–2):43– 61, March–June 2024. ISSN 0218-1959. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0218195924500031.

Iacono:2012:SOT

[IM12] John Iacono and Wolfgang Mulzer. A static optimality transformation with applications to planar point location. International Journal of Computational Geometry and Applications (IJCGA), 22(4):327– ??, August 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Imai:2002:ETG

[IMTI02] Hiroshi Imai, Tomonari Masada, Fumihiko Takeuchi, and Keiko Imai. Enumerating triangulations in general dimensions. International Journal of Computational Geometry and Applications (IJCGA), 12(6):455–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Nakano:2023:CPA

[iN23] Shin ichi Nakano. The coverage problem by aligned disks. International Journal of Computational Geometry and Applications (IJCGA), 33(1–2):13– 23, March–June 2023. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195922410023.

Inkulu:2020:DAV

[IST20] R. Inkulu, K. Sowmya, and Nitish P. Thakur. Dynamic algorithms for visibility polygons in simple polygons. International Journal of Computational Geometry and Applications (IJCGA), 30(01):51– 78, March 2020. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S021819592050003X.

Janardan:1993:MWD

[Jan93] R. Janardan. On maintaining the width and diameter of a planar point-set online. International Journal of Computational Geometry and Applications (IJCGA), 3(??):331–344, 1993. CODEN IJCAEV. ISSN 0218-1959. See also 2nd Int. Symp. Algorithms, 1991, pp. 137–149.

Jiao:2004:OSMa

[JH04a] Xiangmin Jiao and Michael T. Heath. Overlaying surface meshes, Part I: Algorithms. International Journal of Computational Geometry and Applications (IJCGA), 14(6):379–??, December 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Jiao:2004:OSMb

[JH04b] Xiangmin Jiao and Michael T. Heath. Overlaying surface meshes, Part II: Topology preservation and feature matching. International Journal of Computational Geometry and Applications (IJCGA), 14(6): 403–??, December 2004. CO-DEN IJCAEV. ISSN 0218-1959.

Jiang:2015:CPM

[Jia15] Minghui Jiang. On covering points with minimum turns. International Journal of Computational Geometry and Applications (IJCGA), 25(1):1–??, March 2015. CODEN IJCAEV. ISSN 0218-1959.

Jackson:2006:RFD

[JJ06] Bill Jackson and Tibor Jordán. On the rank function of the 3-dimensional rigidity matroid. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):415– ??, December 2006. CODEN IJCAEV. ISSN 0218-1959.

Jackson:2010:OPG

[JJ10] Bill Jackson and Tibor Jordán. Operations preserving global rigidity of generic directionlength frameworks. International Journal of Computational Geometry and Applications (IJCGA), 20(6):685–706, December 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Jabbour:1998:DRA

[JMM98] Toni Jabbour, Christian Mascle, and Roland Maranzana. A database for the representation of assembly features in mechanical products. International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):483–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Jana:2023:MBS

[JMMR23] Satyabrata Jana, Anil Maheshwari, Saeed Mehrabi, and Sasanka Roy. Maximum bipartite subgraphs of geometric intersection graphs. International Journal of Computational Geometry and Applications (IJCGA), 33(3–4): 133–157, September–December 2023. ISSN 0218-1959. URL

https://www.worldscientific.
com/doi/10.1142/S021819592350005X.

Johnson:2021:WGD

[JPV21] Matthew Johnson, Daniël Paulusma, and Erik Jan Van Leeuwen. What graphs are 2-dot product graphs? International Journal of Computational Geometry and Applications (IJCGA), 31(01):

1–16, March 2021. CO-DEN IJCAEV. ISSN 0218-

                                        1959.  URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0218195921500011.

Jiang:2009:FPA

[JS09] D. Jiang and N. F. Stewart. Floating-point arithmetic for computational geometry problems with uncertain data. International Journal of Computational Geometry and Applications (IJCGA), 19(4):371–385, August 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Jermann:2006:DGC

[JTNM06] Christophe Jermann, Gilles Trombettoni, Bertrand Neveu, and Pascal Mathis. Decomposition of geometric constraint systems: a survey. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):379–??, December 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Jain:2024:ACB

[JW24] Princy Jain and Haitao Wang. Algorithms for covering barrier points by mobile sensors with line constraint. International Journal of Computational Geometry and Applications (IJCGA), 34(1–2):1– 23, March–June 2024. CO-DEN IJCAEV. ISSN 0218-

                                        1959.  URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0218195924500018.

Kant:1997:MCV

[Kan97a] Goos Kant. A more compact visibility representation. International Journal of Computational Geometry and Applications (IJCGA), 7(3):197–??, June 1997. CODEN IJCAEV. ISSN 0218-1959.

Kantabutra:1997:RPU

[Kan97b] Vitit Kantabutra. Reaching a point with an unanchored robot arm in a square. International Journal of Computational Geometry and Applications (IJCGA), 7(6):539–??, December 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Kim:2011:CPS

[KBA11] Sang-Sub Kim, Sang Won Bae, and Hee-Kap Ahn. Covering a point set by two disjoint rectangles. International Journal of Computational Geometry and Applications (IJCGA), 21 (3):313–330, June 2011. CO-DEN IJCAEV. ISSN 0218-1959.

Kedem:1997:EBR

[KC97] Klara Kedem and Daniel Cohen. Efficient bitmap resemblance under translations. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):57–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959.

Keil:1997:COP

[Kei97] J. Mark Keil. Covering orthogonal polygons with non-piercing rectangles. International Journal of Computational Geometry and Applications (IJCGA), 7(5):473–??, October 1997. CO-DEN IJCAEV. ISSN 0218-1959.

Khopkar:2014:HRC

[KG14] Abhijeet Khopkar and Sathish Govindarajan. Hardness results for computing optimal locally Gabriel graphs. International Journal of Computational Geometry and Applications (IJCGA), 24(2):153–??, June 2014. CODEN IJCAEV. ISSN 0218-1959.

Hon:2023:GEF

[kHsLtT23] Wing kai Hon, Chung shou Liao, and Meng tsung Tsai. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 33(1–2): 1–2, March–June 2023. CO-DEN IJCAEV. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195923020016.

Kim:2009:GEF

[Kim09] Deok-Soo Kim. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 19(5):387–388, October 2009. CODEN IJCAEV. ISSN 0218-1959.

Kirchner:2007:FCS

[Kir07] Stefan Kirchner. An FPTAS for computing the similarity of three-dimensional point sets.

International Journal of Computational Geometry and Applications (IJCGA), 17(2):161– 174, April 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Kaneko:2005:SBP

[KK05] Atsushi Kaneko and Mikio Kano. Semi-balanced partitions of two sets of points and embeddings of rooted forests. International Journal of Computational Geometry and Applications (IJCGA), 15(3):229–??, June 2005. CODEN IJCAEV. ISSN 0218-1959.

Kumar:2010:AOS

[KK10] Pankaj Kumar and Piyush Kumar. Almost optimal solutions to k-clustering problems. International Journal of Computational Geometry and Applications (IJCGA), 20(4):431– 447, August 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Kim:2005:EVD

[KKS05] Donguk Kim, Deok-Soo Kim, and Kokichi Sugihara. Euclidean Voronoi diagram for circles in a circle. International Journal of Computational Geometry and Applications (IJCGA), 15(2):209–??, April 2005. CODEN IJCAEV. ISSN 0218-1959.

Kaneko:2000:AHC

[KKY00] A. Kaneko, M. Kano, and K. Yoshimoto. Alternating Hamilton cycles with minimum number of crossings in the plane. International Journal of Computational Geometry and Applications (IJCGA), 10(1): 73–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Kamphans:2010:LUM

[KL10a] Tom Kamphans and Elmar Langetepe. Leaving an unknown maze using an errorprone compass. International Journal of Computational Geometry and Applications (IJCGA), 20(3):307–325, June 2010. CODEN IJCAEV. ISSN 0218-1959.

Kapoor:2010:PSG

[KL10b] Sanjiv Kapoor and Xiang-Yang Li. Proximity structures for geometric graphs. International Journal of Computational Geometry and Applications (IJCGA), 20(4):415–429, August 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Lee:2023:FDN

[kLkHsL^{+}23] Kuo kai Lee, Wing kai Hon, Chung shou Liao, Kunihiko Sadakane, and Meng tsung Tsai. Fully dynamic noback-edge-traversal forest via 2D-range queries. International Journal of Computational Geometry and Applications (IJCGA), 33(1–2):43– 54, March–June 2023. ISSN 0218-1959. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0218195922410047.

Korotov:2021:IMA

[KLV21] Sergey Korotov, Lars Fredrik Lund, and Jon Eivind Vatne.

Improved maximum angle estimate for longest-edge bisection. International Journal of Computational Geometry and Applications (IJCGA), 31(4): 183–192, December 2021. CO-DEN IJCAEV. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195922500017.

Katz:2011:GOA

[KM11] Matthew J. Katz and Gila Morgenstern. Guarding orthogonal art galleries with sliding cameras. International Journal of Computational Geometry and Applications (IJCGA), 21 (2):241–250, April 2011. CO-DEN IJCAEV. ISSN 0218-1959.

Krishnan:2001:BBE

[KMG^{+}01] S. Krishnan, D. Manocha, M. Gopi, T. Culver, and J. Keyser. BOOLE: a boundary evaluation system for Boolean combinations of sculptured solids. International Journal of Computational Geometry and Applications (IJCGA), 11(1): 105–??, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Keil:2000:VSD

[KMW00] M. Keil, D. M. Mount, and S. K. Wismath. Visibility stabs and depth-first spiralling on line segments in output sensitive time. International Journal of Computational Geometry and Applications (IJCGA), 10 (5):535–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Kerber:2020:MSE

[KN20] Michael Kerber and Arnur Nigmetov. Metric spaces with expensive distances. International Journal of Computational Geometry and Applications (IJCGA), 30(02):141– 165, June 2020. ISSN 0218-

                                        1959.  URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0218195920500077.

Kanamaru:1994:EEG

[KNA94] Naoyoshi Kanamaru, Takao Nishizeki, and Tetsuo Asano. Efficient enumeration of grid points in a convex polygon and its application to integer programming. International Journal of Computational Geometry and Applications (IJCGA), 4(1):69–85, 1994. CODEN IJ-CAEV. ISSN 0218-1959.

Koike:2002:LPR

[KNN^{+}02] Atsushi Koike, Shin-Ichi Nakano, Takao Nishizeki, Takeshi Tokuyama, and Shuhei Watanabe. Labeling points with rectangles of various shapes. International Journal of Computational Geometry and Applications (IJCGA), 12(6):511–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Khandhawit:2013:LBC

[KPS13] Tirasan Khandhawit, Dimitrios Pagonakis, and Sira Sriswasdi. Lower bound for convex hull area and universal cover problems. International Journal of Computational Geometry and Applications (IJCGA), 23(3): 197–??, June 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Kraft:2020:CHD

[Kra20] Daniel Kraft. Computing the Hausdorff distance of two sets from their distance functions. International Journal of Computational Geometry and Applications (IJCGA), 30(01):19– 49, March 2020. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195920500028.

Kirkpatrick:1999:CCS

[KS99] D. Kirkpatrick and J. Snoeyink.

Computing constrained shortest segments: Butterfly wingspans in logarithmic time. International Journal of Computational Geometry and Applications (IJCGA), 9(1):53–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Keil:2002:TBC

[KS02] Mark Keil and Jack Snoeyink.

On the time bound for convex decomposition of simple polygons. International Journal of Computational Geometry and Applications (IJCGA), 12 (3):181–??, 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Kanda:2005:TDR

[KS05] Takeshi Kanda and Kokichi Sugihara. Two-dimensional range search based on the Voronoi diagram. International Journal of Computational Geometry and Applications (IJCGA), 15(2):151–??, April 2005. CODEN IJCAEV. ISSN 0218-1959.

Kobbelt:2007:GEF

[KS07] Leif Kobbelt and Vadim Shapiro. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 17(5): 401–402, October 2007. CO-DEN IJCAEV. ISSN 0218-1959.

Kramer:2010:MOM

[KS10] Josh Brown Kramer and Lucas Sabalka. Multidimensional online motion planning for a spherical robot. International Journal of Computational Geometry and Applications (IJCGA), 20(6):653–684, December 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Knauer:2011:ANN

[KS11] Christian Knauer and Marc Scherfenberg. Approximate nearest neighbor search under translation invariant Hausdorff distance. International Journal of Computational Geometry and Applications (IJCGA), 21 (3):369–381, June 2011. CO-DEN IJCAEV. ISSN 0218-1959.

Kaplan:2013:FLE

[KS13] Haim Kaplan and Micha Sharir. Finding the largest empty disk containing a query point. International Journal of Computational Geometry and Applications (IJCGA), 23(4–5):335–??, 2013. CODEN IJCAEV. ISSN 0218-1959.

Knauer:2016:GEF

[KS16] Christian Knauer and Michiel Smid. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 26(3–4):iii– iv, September/December 2016. CODEN IJCAEV. ISSN 0218-1959.

Kusakari:1999:SPP

[KSN99] Y. Kusakari, H. Suzuki, and T. Nishizeki. A shortest pair of paths on the plane with obstacles and crossing areas. International Journal of Computational Geometry and Applications (IJCGA), 9(2):151–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Kirkpatrick:2002:KCD

[KSS02] David Kirkpatrick, Jack Snoeyink, and Bettina Speckmann. Kinetic collision detection for simple polygons. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2):3–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Kim:2001:LRM

[KSY^{+}01] Sung Kwon Kim, Chan-Su Shin, Tae-Cheon Yang, et al. Labeling a rectilinear map with sliding labels. International Journal of Computational Geometry and Applications (IJCGA), 11(2):167–179, 2001. CODEN IJCAEV. ISSN 0218-1959.

Kakoulis:2003:UAA

[KT03] Konstantinos G. Kakoulis and Ioannis G. Tollis. A unified approach to automatic label placement. International Journal of Computational Geometry and Applications (IJCGA), 13 (1):23–??, February 2003. CO-DEN IJCAEV. ISSN 0218-1959.

Katoh:2002:PPO

[KTT02] Naoki Katoh, Hisao Tamaki, and Takeshi Tokuyama. Parametric polymatroid optimization and its geometric applications. International Journal of Computational Geometry and Applications (IJCGA), 12 (5):429–??, October 2002. CO-DEN IJCAEV. ISSN 0218-1959.

Kranakis:1999:ITS

[KU99] E. Kranakis and J. Urrutia. Isomorphic triangulations with small number of Steiner points. International Journal of Computational Geometry and Applications (IJCGA), 9(2):171–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Kano:2010:BSB

[KU10] M. Kano and Miyuki Uno. Balanced subdivisions with boundary condition of two sets of points in the plane. International Journal of Computational Geometry and Applications (IJCGA), 20(5):527–541, October 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Korotov:2023:DAS

[KV23] Sergey Korotov and Jon Eivind Vatne. On dihedral angle sums of prisms and hexahedra. International Journal of Computational Geometry and Applications (IJCGA), 33(3–4):85–95, September– December 2023. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195923500036.

Kirkpatrick:2014:PTA

[KYZ14] David Kirkpatrick, Boting Yang, and Sandra Zilles. A polynomial-time algorithm for computing the resilience of arrangements of ray sensors. International Journal of Computational Geometry and Applications (IJCGA), 24(3):225–??, September 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Kameda:2010:FAD

[KZ10] Tsunehiko Kameda and John Z.

Zhang. Finding all door locations that make a room searchable. International Journal of Computational Geometry and Applications (IJCGA), 20(2): 175–201, April 2010. CODEN IJCAEV. ISSN 0218-1959.

Labelle:2008:SRL

[Lab08] François Labelle. Sliver removal by lattice refinement. International Journal of Computational Geometry and Applications (IJCGA), 18(6):509–532, December 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Liu:2015:OSI

[LD15] David J. T. Liu and Qiang Du. Optimization of subdivision invariant tetrahedra. International Journal of Computational Geometry and Applications (IJCGA), 25(1):37–??, March 2015. CODEN IJCAEV. ISSN 0218-1959.

Liu:2020:DRS

[LDHX20] Yangwei Liu, Hu Ding, Ziyun Huang, and Jinhui Xu. Distributed and robust support vector machine. International Journal of Computational Geometry and Applications (IJCGA), 30(3–4): 213–233, September–December 2020. ISSN 0218-1959. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0218195920500107.

Lee:2003:CEN

[Lee03] D. T. Lee. Chief Editor’s notice. International Journal of Computational Geometry and Applications (IJCGA), 13(1):1–??, February 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Linhart:2003:ESL

[LHHHP03] Chaim Linhart, Dan Halperin, Iddo Hanniel, and Sariel Har-Peled. An experimental study of on-line methods for zone construction in arrangements of lines in the plane. International Journal of Computational Geometry and Applications (IJCGA), 13(6):463–??, December 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Liu:2024:SEO

[LHX24] Yangwei Liu, Ziyun Huang, and Jinhui Xu. A space-efficient one-pass online SVM algorithm. International Journal of Computational Geometry and Applications (IJCGA), 34(1–2): 63–79, March–June 2024. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195924500043.

Luo:2011:MSS

[LLCC11] Cheng-Wei Luo, Hsiao-Fei Liu, Peng-An Chen, and Kun-Mao Chao. Minkowski sum selection and finding. International Journal of Computational Geometry and Applications (IJCGA), 21(3):283–311, June 2011. CODEN IJCAEV. ISSN 0218-1959.

Lin:1997:ECD

[LM97] Ming Lin and Dinesh Manocha.

Efficient contact determination in dynamic environments. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):123–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959.

Lin:1998:GEF

[LM98] Ming C. Lin and Dinesh Manocha. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 8(4): 385–??, August 1998. CODEN IJCAEV. ISSN 0218-1959.

Loffler:2011:ECT

[Löf11] Maarten Löffler. Existence and computation of tours through imprecise points. International Journal of Computational Geometry and Applications (IJCGA), 21(1):1–24, February 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Lopez-Ortiz:2001:LBS

[LOS01] Alejandro López-Ortiz and Sven Schuierer. Lower bounds for streets and generalized streets. International Journal of Computational Geometry and Applications (IJCGA), 11(4):401–421, 2001. CODEN IJCAEV. ISSN 0218-1959.

Lee:2000:SPR

[LPC00] J.-H. Lee, S.-M. Park, and K.- Y. Chwa. Searching a polygonal room with one door by a 1-searcher. International Journal of Computational Geometry and Applications (IJCGA), 10 (2):201–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Lopez:2000:EAC

[LR00] M. A. Lopez and S. Reisner. Efficient approximation of convex polygons. International Journal of Computational Geometry and Applications (IJCGA), 10 (5):445–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Liu:2008:FPS

[LS08] Yuanxin Liu and Jack Snoeyink. Faraway point: a sentinel point for Delaunay computation. International Journal of Computational Geometry and Applications (IJCGA), 18(4):343– 355, August 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Lienhardt:2004:CPS

[LSB04] Pascal Lienhardt, Xavier Skapin, and Antoine Bergey. Cartesian product of simplicial and cellular structures. International Journal of Computational Geometry and Applications (IJCGA), 14(3):115–??, June 2004. CODEN IJCAEV. ISSN 0218-1959.

Lee:1998:GDV

[LSS98] D. T. Lee, Chin-Fang Shen, and Dennis S. Sheu. Geosheet: a distributed visualization tool for geometric algorithms. International Journal of Computational Geometry and Applications (IJCGA), 8(2):119–??, April 1998. CODEN IJCAEV. ISSN 0218-1959.

Lavalle:2002:ASP

[LSS02] Steven M. Lavalle, Borislav H.

Simov, and Giora Slutzki. An algorithm for searching a polygonal region with a flashlight. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2):87–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Li:2004:ECL

[LW04] Xiang-Yang Li and Yu Wang.

Efficient construction of low weighted bounded degree planar spanner. International Journal of Computational Geometry and Applications (IJCGA), 14(1–2):69–??, April 2004. CO-DEN IJCAEV. ISSN 0218-1959.

Lingas:2012:LTA

[LWŻ12] Andrzej Lingas, Agnieszka Wasylewicz, and Paweł Żyliński. Linear-time 3-approximation algorithm for the r-star covering problem. International Journal of Computational Geometry and Applications (IJCGA), 22(2):103–??, April 2012. CODEN IJCAEV. ISSN 0218-1959.

Lee:2017:MMM

[LWZ17] Victor C. S. Lee, Haitao Wang, and Xiao Zhang. Minimizing the maximum moving cost of interval coverage. International Journal of Computational Geometry and Applications (IJCGA), 27(3):187–205, September 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Lee:1997:FRP

[LYW97] D. T. Lee, C. D. Yang, and C. K. Wong. Finding rectilinear paths among obstacles in a two-layer interconnection model. International Journal of Computational Geometry and Applications (IJCGA), 7(6):581–??, December 1997. CODEN IJCAEV. ISSN 0218-1959.

Maftuleac:2014:ADP

[Maf14] Daniela Maftuleac. Algorithms for distance problems in planar complexes of global nonpositive curvature. International Journal of Computational Geometry and Applications (IJCGA), 24 (1):1–??, March 2014. CODEN IJCAEV. ISSN 0218-1959.

Malandain:2002:CDP

[MB02] Grégoire Malandain and Jean-Daniel Boissonnat. Computing the diameter of a point set. International Journal of Computational Geometry and Applications (IJCGA), 12(6):489–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Manocha:1991:NAS

[MC91] Dinesh Manocha and John F.

Canny. A new approach to surface intersection. International Journal of Computational Geometry and Applications (IJCGA), 1(4):491– 516, 1991. CODEN IJCAEV. ISSN 0218-1959. URL ftp:/

/ftp.cs.unc.edu/pub/users/
manocha/PAPERS/INTERSECT/
IJCGA.pdf;   ftp://ftp.cs.
unc.edu/pub/users/manocha/
PAPERS/INTERSECT/IJCGA.ps.
Z.

Michelucci:2006:IBT

[MF06] Dominique Michelucci and Sebti Foufou. Interval-based tracing of strange attractors. International Journal of Computational Geometry and Applications (IJCGA), 16(1):27– ??, February 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Menon:1998:FFM

[MG98] Jai Menon and Baining Guo. Free-form modeling in bilateral Brep and CSG representation schemes. International Journal of Computational Geometry and Applications (IJCGA), 8(5–6):537–??, October–December 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Mahapatra:2015:PTA

[MGD15] Priya Ranjan Sinha Mahapatra, Partha P. Goswami, and Sandip Das. Placing two axisparallel squares to maximize the number of enclosed points. International Journal of Computational Geometry and Applications (IJCGA), 25(4):263–??, December 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Mukhopadhyay:2009:ISI

[MGR09] Asish Mukhopadhyay, Eugene Greene, and S. V. Rao. On intersecting a set of isothetic line segments with a convex polygon of minimum area. International Journal of Computational Geometry and Applications (IJCGA), 19(6):557–577, December 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Muller-Hannemann:2000:HQQ

[MH00] M. Müller-Hannemann. High quality quadrilateral surface meshing without template restrictions: a new approach based on network flow techniques. International Journal of Computational Geometry and Applications (IJCGA), 10 (3):285–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Miura:2006:IRD

[MHN06] Kazuyuki Miura, Hiroki Haga, and Takao Nishizeki. Inner rectangular drawings of plane graphs. International Journal of Computational Geometry and Applications (IJCGA), 16 (2–3):249–??, June 2006. CO-DEN IJCAEV. ISSN 0218-1959.

Muller-Hannemann:2007:HAO

[MHS07] Matthias Müller-Hannemann and Anna Schulze. Hardness and approximation of octilinear Steiner trees. International Journal of Computational Geometry and Applications (IJCGA), 17(3):231–260, June 2007. CODEN IJCAEV. ISSN 0218-1959.

Muller-Hannemann:2000:QRC

[MHW00] M. Müller-Hannemann and K. Weihe. Quadrangular refinements of convex polygons with an application to finite-element meshes. International Journal of Computational Geometry and Applications (IJCGA), 10(1):1–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Mitchell:1997:FCT

[Mit97] Scott A. Mitchell. Finding a covering triangulation whose maximum angle is provably small. International Journal of Computational Geometry and Applications (IJCGA), 7(1– 2):5–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959.

Mitchell:2000:HFI

[Mit00] S. A. Mitchell. High fidelity interval assignment. International Journal of Computational Geometry and Applications (IJCGA), 10(4):399–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Mitchell:2004:EF

[Mit04] Joseph S. B. Mitchell. Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 14(4–5):231–??, October 2004. CODEN IJCAEV. ISSN 0218-1959.

Myers:2012:PSD

[MJ12] Yonatan Myers and Leo Joskowicz. Point set distance and orthogonal range problems with dependent geometric uncertainties. International Journal of Computational Geometry and Applications (IJCGA), 22(6): 517–??, December 2012. CO-DEN IJCAEV. ISSN 0218-1959.

Murphy:2001:PPS

[MMG01] Michael Murphy, David M. Mount, and Carl W. Gable. A point-placement strategy for conforming Delaunay tetrahedralization. International Journal of Computational Geometry and Applications (IJCGA), 11 (6):669–??, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Memarsadeghi:2007:FII

[MMNM07] Nargess Memarsadeghi, David M.

Mount, Nathan S. Netanyahu, and Jacqueline Le Moigne. A fast implementation of the ISO-DATA clustering algorithm. International Journal of Computational Geometry and Applications (IJCGA), 17(1):71–103, February 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Mehlhorn:2001:FSA

[MMR01] Kurt Mehlhorn, Stefan Meiser, and Ronald Rasch. Furthest site abstract Voronoi diagrams. International Journal of Computational Geometry and Applications (IJCGA), 11(6):583–??, 2001. CODEN IJCAEV. ISSN 0218-1959.

Mitchell:1997:QSR

[MMS97] Joseph S. B. Mitchell, David M.

Mount, and Subhash Suri. Query-sensitive ray shooting. International Journal of Computational Geometry and Applications (IJCGA), 7(4):317– ??, August 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Mitra:2004:ESN

[MNG04] Niloy J. Mitra, An Nguyen, and Leonidas Guibas. Estimating surface normals in noisy point cloud data. International Journal of Computational Geometry and Applications (IJCGA), 14(4–5):261–??, October 2004.

CODEN IJCAEV. ISSN 0218-1959.

Mount:2000:QAR

[MNP^{+}00] D. M. Mount, N. S. Netanyahu, C. D. Piatko, R. Silverman, and A. Y. Wu. Quantile approximation for robust statistical estimation and k-enclosing problems. International Journal of Computational Geometry and Applications (IJCGA), 10 (6):593–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Mitchell:2001:CGC

[MO01] Joseph S. B. Mitchell and Joseph O’Rourke. Computational geometry column 42. International Journal of Computational Geometry and Applications (IJCGA), 11(5):573–582, 2001. CODEN IJCAEV. ISSN 0218-1959.

Miller:2005:WWD

[MPW05] Gary L. Miller, Steven E. Pav, and Noel J. Walkington. When and why Delaunay refinement algorithms work. International Journal of Computational Geometry and Applications (IJCGA), 15(1):25–??, February 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Mukhopadhyay:2003:CLE

[MR03] Asish Mukhopadhyay and S. V. Rao. On computing a largest empty arbitrarily oriented rectangle. International Journal of Computational Geometry and Applications (IJCGA), 13(3):

257–??, June 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Meijer:2005:GEF

[MR05] Henk Meijer and David Rappaport. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 15(6):545–??, December 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Mustafa:2015:CCP

[MRM15] Nabil H. Mustafa, Saurabh Ray, and Nabil H. Mustafa. k-centerpoints conjectures for pointsets in R^{d}. International Journal of Computational Geometry and Applications (IJCGA), 25(3):??, September 2015. CODEN IJ-CAEV. ISSN 0218-1959.

Martini:1999:MNP

[MS99] H. Martini and V. Soltan. Minimum number of pieces in a convex partition of a polygonal domain. International Journal of Computational Geometry and Applications (IJCGA), 9(6):599–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Mehlhorn:2003:IFT

[MS03] Kurt Mehlhorn and Michael Seel. Infimaximal frames: a technique for making lines look like segments. International Journal of Computational Geometry and Applications (IJCGA), 13(3):241–??, June 2003. CODEN IJCAEV. ISSN 0218-1959.

Michelucci:2006:ICC

[MS06] Dominique Michelucci and Pascal Schreck. Incidence constraints: a combinatorial approach. International Journal of Computational Geometry and Applications (IJCGA), 16 (5–6):443–??, December 2006. CODEN IJCAEV. ISSN 0218-1959.

Milenkovic:2007:AAA

[MS07a] Victor Milenkovic and Elisha Sacks. An approximate arrangement algorithm for semialgebraic curves. International Journal of Computational Geometry and Applications (IJCGA), 17(2):175–198, April 2007. CODEN IJCAEV. ISSN 0218-1959.

Milenkovic:2007:MCM

[MS07b] Victor Milenkovic and Elisha Sacks. A monotonic convolution for Minkowski sums. International Journal of Computational Geometry and Applications (IJCGA), 17(4):383–396, August 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Milenkovic:2010:TAM

[MS10] Victor Milenkovic and Elisha Sacks. Two approximate Minkowski sum algorithms. International Journal of Computational Geometry and Applications (IJCGA), 20(4):485– 509, August 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Mulzer:2014:ATT

[MS14] Wolfgang Mulzer and Yannik Stein. Algorithms for tolerant Tverberg partitions. International Journal of Computational Geometry and Applications (IJCGA), 24(4):261–??, December 2014. CODEN IJ-CAEV. ISSN 0218-1959.

Milenkovic:2022:EPE

[MS22] Victor Milenkovic and Elisha Sacks. Efficient predicate evaluation using randomized degeneracy detection. International Journal of Computational Geometry and Applications (IJCGA), 32(1–2):39– 54, March–June 2022. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195922500054.

Milenkovic:2019:FDD

[MSB19] Victor Milenkovic, Elisha Sacks, and Nabeel Butt. Fast detection of degenerate predicates in free space construction. International Journal of Computational Geometry and Applications (IJCGA), 29(03): 219–237, September 2019. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195919500067.

Milenkovic:2013:PSM

[MST13] Victor Milenkovic, Elisha Sacks, and Steven Trac. Planar shape manipulation using approximate geometric primitives. International Journal of Computational Geometry and Applications (IJCGA), 23(1):1– ??, February 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Miller:1999:DGG

[MTT99] G. L. Miller, D. Talmor, and S.-H. Teng. Data generation for geometric algorithms on non-uniform distributions. International Journal of Computational Geometry and Applications (IJCGA), 9(6):577–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Mucke:1998:RIT

[Müc98]  Ernst P. Mücke.  A ro-

bust implementation for threedimensional Delaunay triangulations. International Journal of Computational Geometry and Applications (IJCGA), 8(2):255–??, April 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Moet:2007:RIT

[MVV07] Esther Moet, Marc Van Kreveld, and René Van Oostrum. Region intervisibility in terrains. International Journal of Computational Geometry and Applications (IJCGA), 17(4): 331–347, August 2007. CODEN IJCAEV. ISSN 0218-1959.

Nabli:2021:TTC

[Nab21] Hédi Nabli. Truchet tiles and combinatorial arabesque. International Journal of Computational Geometry and Applications (IJCGA), 31(01):17– 37, March 2021. ISSN 0218-

                                        1959.  URL https://www.

worldscientific.com/doi/
10.1142/S0218195921500035.

Narasimhan:1999:HTS

[Nar99] G. Narasimhan. On Hamiltonian triangulations in simple polygons. International Journal of Computational Geometry and Applications (IJCGA), 9(3):261–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Nekrich:2013:EMO

[Nek13] Yakov Nekrich. External memory orthogonal range reporting with fast updates. International Journal of Computational Geometry and Applications (IJCGA), 23(2):141–??, April 2013. CODEN IJCAEV. ISSN 0218-1959.

Nguyen:2012:EGR

[Ngu12] Viet-Hang Nguyen. 1-extensions and global rigidity of generic direction-length frameworks. International Journal of Computational Geometry and Applications (IJCGA), 22(6):577–??, December 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Nielsen:2009:ASE

[NN09] Frank Nielsen and Richard Nock. Approximating smallest enclosing balls with applications to machine learning. International Journal of Computational Geometry and Applications (IJCGA), 19(5):389–414, October 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Nollenburg:2017:PGD

[NPR17] Martin Nöllenburg, Roman Prutkin, and Ignaz Rutter. Partitioning graph drawings and triangulated simple polygons into greedily routable regions. International Journal of Computational Geometry and Applications (IJCGA), 27(1–2):121– 158, March/June 2017. CO-DEN IJCAEV. ISSN 0218-1959.

Nishida:2009:BSV

[NS09] Tetsushi Nishida and Kokichi Sugihara. Boat-sail Voronoi diagram and its application. International Journal of Computational Geometry and Applications (IJCGA), 19(5):425–440, October 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Nielsen:1998:OSC

[NY98] Franck Nielsen and Mariette Yvinec. An output-sensitive convex hull algorithm for planar objects. International Journal of Computational Geometry and Applications (IJCGA), 8(1):39–??, February 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Nowakowski:2006:BOT

[NZ06] Richard J. Nowakowski and Norbert Zeh. Boundaryoptimal triangulation flooding. International Journal of Computational Geometry and Applications (IJCGA), 16(2–3):271– ??, June 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Nilsson:2020:HKE

[NŻ20] Bengt J. Nilsson and Paweł Żyliński. How to keep an eye on small things. International Journal of Computational Geometry and Applications (IJCGA), 30(02): 97–120, June 2020. CO-DEN IJCAEV. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195920500053.

Obermeyer:2011:CAS

[OGB11] Karl J. Obermeyer, Anurag Ganguli, and Francesco Bullo. A complete algorithm for searchlight scheduling. International Journal of Computational Geometry and Applications (IJCGA), 21(1):101–130, February 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Owen:2010:FSR

[OP10] J. C. Owen and S. C. Power.

Frameworks symmetry and rigidity. International Journal of Computational Geometry and Applications (IJCGA), 20(6):723–750, December 2010. CODEN IJCAEV. ISSN 0218-1959.

ORourke:1997:CGCa

[O’R97a] Joseph O’Rourke. Computational geometry column 30. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):165–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959.

ORourke:1997:CGCb

[O’R97b] Joseph O’Rourke. Computational geometry column 31. International Journal of Computational Geometry and Applications (IJCGA), 7(4):379–??, August 1997. CODEN IJ-CAEV. ISSN 0218-1959.

ORourke:1997:CGCc

[O’R97c] Joseph O’Rourke. Computational geometry column 32. International Journal of Computational Geometry and Applications (IJCGA), 7(5):509–??, October 1997. CODEN IJ-CAEV. ISSN 0218-1959.

ORourke:1998:CGC

[O’R98] Joseph O’Rourke. Computational geometry column 33. International Journal of Computational Geometry and Applications (IJCGA), 8(3):381–??, June 1998. CODEN IJCAEV. ISSN 0218-1959.

ORourke:1999:CGCb

[O’R99a] J. O’Rourke. Computational geometry column 36. International Journal of Computational Geometry and Applications (IJCGA), 9(6):615–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

ORourke:1999:CGCa

[O’R99b] Joseph O’Rourke. Computational geometry column 35. International Journal of Computational Geometry and Applications (IJCGA), 9(4–5):513–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2000:CGCa

[O’R00a] J. O’Rourke. Computational geometry column 38. International Journal of Computational Geometry and Applications (IJCGA), 10(2):221–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2000:CGCb

[O’R00b] J. O’Rourke. Computational geometry column 39. International Journal of Computational Geometry and Applications (IJCGA), 10(4):441–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2000:CGCc

[O’R00c] J. O’Rourke. Computational geometry column 40. International Journal of Computational Geometry and Applications (IJCGA), 10(6):649–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2001:CGC

[O’R01] Joseph O’Rourke. Computational geometry column 41. International Journal of Computational Geometry and Applications (IJCGA), 11(2):239–242, 2001. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2002:CGC

[O’R02] Joseph O’Rourke. Computational geometry column 43. International Journal of Computational Geometry and Applications (IJCGA), 12(3):263–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2003:CGC

[O’R03] Joseph O’Rourke. Computational geometry column 44. International Journal of Computational Geometry and Applications (IJCGA), 13(3):273–??, June 2003. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2004:CGCa

[O’R04a] Joseph O’Rourke. Computational geometry column 45. International Journal of Computational Geometry and Applications (IJCGA), 14(3):227–??, June 2004. CODEN IJCAEV. ISSN 0218-1959.

ORourke:2004:CGCb

[O’R04b] Joseph O’Rourke. Computational geometry column 46. International Journal of Computational Geometry and Applications (IJCGA), 14(6):475–??, December 2004. CODEN IJ-CAEV. ISSN 0218-1959.

ORourke:2006:CGC

[O’R06] Joseph O’Rourke. Computational geometry column 47. International Journal of Computational Geometry and Applications (IJCGA), 16(4):373–??, August 2006. CODEN IJ-CAEV. ISSN 0218-1959.

ORourke:2007:CGC

[O’R07] Joseph O’Rourke. Computational geometry column 48. International Journal of Computational Geometry and Applications (IJCGA), 17(4):397–399, August 2007. CODEN IJ-CAEV. ISSN 0218-1959.

ODunlaing:2000:HCE

[ÓWW00] C. Ó Dúnlaing, C. Watt, and D. Wilkins. Homeomorphism of 2-complexes is equivalent to graph isomorphism. International Journal of Computational Geometry and Applications (IJCGA), 10(5):453–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Papadopoulou:1999:PNC

[Pap99]  E. Papadopoulou.  k-pairs

non-crossing shortest paths in a simple polygon. International Journal of Computational Geometry and Applications (IJCGA), 9(6):533–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Papadopoulou:2013:FLS

[PD13] Evanthia Papadopoulou and Sandeep Kumar Dey. On the farthest line-segment Voronoi diagram. International Journal of Computational Geometry and Applications (IJCGA), 23(6):443–??, December 2013. CODEN IJCAEV. ISSN 0218-1959.

Petitjean:1998:CGA

[Pet98] Sylvain Petitjean. A computational geometric approach to visual hulls. International Journal of Computational Geometry and Applications (IJCGA), 8(4):407–??, August 1998. CO-DEN IJCAEV. ISSN 0218-1959.

Papadopoulou:2001:VDS

[PL01] Evantha Papadopoulou and D. T. Lee. The L_{∞} Voronoi diagram of segments and VLSI applications. International Journal of Computational Geometry and Applications (IJCGA), 11 (5):503–528, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Papadopoulou:2004:HVD

[PL04] Evanthia Papadopoulou and D. T. Lee. The Hausdorff Voronoi diagram of polygonal objects: a divide and conquer approach. International Journal of Computational Geometry and Applications (IJCGA), 14(6):421–??, December 2004. CODEN IJCAEV. ISSN 0218-1959.

Park:2002:SRT

[PLC02] Sang-Min Park, Jae-Ha Lee, and Kyung-Yong Chwa. Searching a room by two guards. International Journal of Computational Geometry and Applications (IJCGA), 12(4):339– ??, August 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Poon:2009:ULP

[Poo09] Sheung-Hung Poon. On unfolding lattice polygons/trees and diameter-4 trees. International Journal of Computational Geometry and Applications (IJCGA), 19(3):289–321, June 2009. CODEN IJCAEV. ISSN 0218-1959.

Porschen:2009:RCP

[Por09] Stefan Porschen. On rectangular covering problems. International Journal of Computational Geometry and Applications (IJCGA), 19(4):325–340, August 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Provatidis:2022:NRR

[Pro22] Christopher G. Provatidis.

Non-rational and rational transfinite interpolation using Bernstein polynomials. International Journal of Computational Geometry and Applications (IJCGA), 32(1–2):55– 89, March–June 2022. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195922500030.

Pilz:2019:CQB

[PS19] Alexander Pilz and Carlos Seara. Convex quadrangulations of bichromatic point sets. International Journal of Computational Geometry and Applications (IJCGA), 29(4):289– 299, December 2019. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195919500109.

Pajarola:2001:VGC

[PW01] Renato Pajarola and Peter Widmayer. Virtual geoexploration: Concepts and design choices. International Journal of Computational Geometry and Applications (IJCGA), 11(1):1–14, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Papadopoulou:2015:HVD

[PX15] Evanthia Papadopoulou and Jinhui Xu. The L_{∞} Hausdorff Voronoi diagram revisited. International Journal of Computational Geometry and Applications (IJCGA), 25(2):??, June 2015. CODEN IJCAEV. ISSN 0218-1959.

Rababah:2005:DRT

[Rab05] Abedallah Rababah. L_{2} degree reduction of triangular Bézier surfaces with common tangent planes at vertices. International Journal of Computational Geometry and Applications (IJCGA), 15(5):477–??, October 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Reshetov:2014:UPF

[Res14]  Alexander Reshetov.  A

unistable polyhedron with 14 faces. International Journal of Computational Geometry and Applications (IJCGA), 24(1): 39–??, March 2014. CODEN IJCAEV. ISSN 0218-1959.

Rokne:2009:GEF

[Rok09] Jon Rokne. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 19 (4):323–324, August 2009. CO-DEN IJCAEV. ISSN 0218-1959.

Rorabaugh:2019:BCM

[Ror19] Danny Rorabaugh. A bound on a convexity measure for point sets. International Journal of Computational Geometry and Applications (IJCGA), 29(4): 301–306, December 2019. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195919500110.

Roy:2016:PVG

[Roy16] Bodhayan Roy. Point visibility graph recognition is NP-hard. International Journal of Computational Geometry and Applications (IJCGA), 26(1):1–??, March 2016. CODEN IJCAEV. ISSN 0218-1959.

Ratschek:2000:EOC

[RR00] H. Ratschek and J. Rokne. Exact and optimal convex hulls in 2D. International Journal of Computational Geometry and Applications (IJCGA), 10 (2):109–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Richards:1999:MST

[RS99] D. Richards and J. S. Salowe. Mixed spanning trees in theory and practice. International Journal of Computational Geometry and Applications (IJCGA), 9(3):277–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Rahi:2007:MCS

[RS07] Sahand Jamal Rahi and Kim Sharp. Mapping complicated surfaces onto a sphere. International Journal of Computational Geometry and Applications (IJCGA), 17(4):305–329, August 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Reif:2011:AOK

[RS11] John Reif and Sam Slee. Asymptotically optimal kinodynamic motion planning for a class of modular selfreconfigurable robots. International Journal of Computational Geometry and Applications (IJCGA), 21(2):131–155, April 2011. CODEN IJCAEV. ISSN 0218-1959.

Ramaswami:2005:CQM

[RSS^{+}05] Suneeta Ramaswami, Marcelo Siqueira, Tessa Sundaram, Jean Gallier, and James Gee. Constrained quadrilateral meshes of bounded size. International Journal of Computational Geometry and Applications (IJCGA), 15(1):55–??, February 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Rand:2011:DRA

[RW11] Alexander Rand and Noel Walkington. Delaunay refinement algorithms for estimating local feature size in 2D and 3D. International Journal of Computational Geometry and Applications (IJCGA), 21(5):507– 543, October 2011. CODEN IJ-CAEV. ISSN 0218-1959.

Sangwin:2009:DBM

[San09] C. J. Sangwin. Deceiving the V - block method for assessment of departure from roundness. International Journal of Computational Geometry and Applications (IJCGA), 19(6):521–532, December 2009. CODEN IJ-CAEV. ISSN 0218-1959.

Sheffer:2000:VTO

[SBBC00] A. Sheffer, M. Bercovier, T. Blacker, and J. Clements. Virtual topology operators for meshing. International Journal of Computational Geometry and Applications (IJCGA), 10 (3):309–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Schneiders:2000:OBH

[Sch00] R. Schneiders. Octree-based hexahedral mesh generation. International Journal of Computational Geometry and Applications (IJCGA), 10(4):383–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Scheffer:2016:MFC

[Sch16] Christian Scheffer. More flexible curve matching via the partial Fréchet similarity. International Journal of Computational Geometry and Applications (IJCGA), 26(1):33–??, March 2016. CODEN IJCAEV. ISSN 0218-1959.

Segal:1999:PSA

[Seg99] M. Segal. On piercing sets of axis-parallel rectangles and rings. International Journal of Computational Geometry and Applications (IJCGA), 9(3):219–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Sud:2007:HPM

[SFM07] Avneesh Sud, Mark Foskey, and Dinesh Manocha. Homotopypreserving medial axis simplification. International Journal of Computational Geometry and Applications (IJCGA), 17(5):423–451, October 2007. CODEN IJCAEV. ISSN 0218-1959.

Shapiro:1997:EMG

[Sha97a] Vadim Shapiro. Errata: “Maintenance of Geometric Representations Through Space Decompositions”. International Journal of Computational Geometry and Applications (IJCGA), 7(4):383–??, August 1997. CO-DEN IJCAEV. ISSN 0218-1959. See [Sha97b].

Shapiro:1997:MGR

[Sha97b] Vadim Shapiro. Maintenance of geometric representations through space decompositions. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):21–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959. See erratum [Sha97a].

Shapiro:1999:WFS

[Sha99] V. Shapiro. Well-formed set representations of solids. International Journal of Computational Geometry and Applications (IJCGA), 9(2):125–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Shapiro:2001:CDT

[Sha01] Vadim Shapiro. A convex deficiency tree algorithm for curved polygons. International Journal of Computational Geometry and Applications (IJCGA), 11 (2):215–238, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Sugihara:1994:RTO

[SI94] Kokichi Sugihara and Masao Iri. A robust topologyoriented incremental algorithm for Voronoi diagrams. International Journal of Computational Geometry and Applications (IJCGA), 4(2):179–228, June 1994. CODEN IJCAEV. ISSN 0218-1959.

Sitharam:2006:WFS

[Sit06] Meera Sitharam. Well-formed systems of point incidences for resolving collections of rigid bodies. International Journal of Computational Geometry and Applications (IJCGA), 16 (5–6):591–??, December 2006. CODEN IJCAEV. ISSN 0218-1959.

Smid:1999:WRS

[SJ99] M. Smid and R. Janardan. On the width and roundness of a set of points in the plane. International Journal of Computational Geometry and Applications (IJCGA), 9(1):97–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Sugihara:2008:GEF

[SK08] Kokichi Sugihara and Deok-Soo Kim. Guest Editors’ foreword. International Journal of Computational Geometry and Applications (IJCGA), 18(4):273– 274, August 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Srinivasaraghavan:2000:OEV

[SM00] G. Srinivasaraghavan and A. Mukhopadhyay. Orthogonal edge visibility graphs of polygons with holes. International Journal of Computational Geometry and Applications (IJCGA), 10(1):79–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Schreck:2006:GCS

[SM06] Pascal Schreck and Pascal Mathis. Geometrical constraint system decomposition: a multi-group approach. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):431–??, December 2006. CODEN IJ-CAEV. ISSN 0218-1959.

SerrE:2006:NCM

[SOR06] Philippe Serré, Auxkin Ortuzar, and Alain Rivière. Non-Cartesian modelling for analysis of the consistency of a geometric specification for conceptual design. International Journal of Computational Geometry and Applications (IJCGA), 16 (5–6):549–??, December 2006. CODEN IJCAEV. ISSN 0218-1959.

Scorzelli:2008:PSM

[SPP08] Giorgio Scorzelli, Alberto Paoluzzi, and Valerio Pascucci. Parallel solid modeling using BSP dataflow. International Journal of Computational Geometry and Applications (IJCGA), 18(5):441–467, October 2008. CODEN IJ-CAEV. ISSN 0218-1959.

Shin:2008:VDP

[SPPK08] Hayong Shin, Seyoun Park, Eonjin Park, and Deok-Soo Kim. Voronoi diagram of a polygon in chessboard metric and maskless lithographic applications. International Journal of Computational Geometry and Applications (IJCGA), 18 (4):357–371, August 2008. CO-DEN IJCAEV. ISSN 0218-1959.

Scharf:2011:IPL

[SS11] Ludmila Scharf and Marc Scherfenberg. Inducing polygons of line arrangements. International Journal of Computational Geometry and Applications (IJCGA), 21(3):351–368, June 2011. CODEN IJCAEV. ISSN 0218-1959.

Stoimenow:2021:ENC

[Sto21] Alexander Stoimenow. Exchangeability and non-conjugacy of braid representatives. International Journal of Computational Geometry and Applications (IJCGA), 31(01):39– 73, March 2021. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195921500047.

Spielman:2007:PDR

[STÜ07] Daniel A. Spielman, Shang-Hua Teng, and Alper Üngör. Parallel Delaunay refinement: Algorithms and analyses. International Journal of Computational Geometry and Applications (IJCGA), 17(1):1–30, February 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Suzuki:2001:SPR

[STYK01] Ichiro Suzuki, Yuichi Tazoe, Masafumi Yamashita, and T. Kameda. Searching a polygonal region from the boundary. International Journal of Computational Geometry and Applications (IJCGA), 11(5):529– 553, 2001. CODEN IJCAEV. ISSN 0218-1959.

Shirakawa:2013:CDT

[SU13] Toshihiro Shirakawa and Ryuhei Uehara. Common developments of three incongruent orthogonal boxes. International Journal of Computational Geometry and Applications (IJCGA), 23(1):65–??, February 2013. CODEN IJ-CAEV. ISSN 0218-1959.

Suderman:2004:PLD

[Sud04] Matthew Suderman. Pathwidth and layered drawings of trees. International Journal of Computational Geometry and Applications (IJCGA), 14(3):203–??, June 2004. CODEN IJCAEV. ISSN 0218-1959.

Sugihara:1992:VDR

[Sug92] K. Sugihara. Voronoi diagrams in a river. International Journal of Computational Geometry and Applications (IJCGA), 2(1):29–48, 1992. CODEN IJ-CAEV. ISSN 0218-1959.

Sugihara:2003:GEF

[Sug03]  Kokichi Sugihara.  Guest

Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 13(4):277–??, August 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Sendra:2001:ORP

[SV01] J. Rafael Sendra and Carlos Villarino. Optimal reparametrization of polynomial algebraic curves. International Journal of Computational Geometry and Applications (IJCGA), 11(4): 439–453, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Silveira:2010:FCD

[SV10] Rodrigo I. Silveira and René Van Oostrum. Flooding countries and destroying dams. International Journal of Computational Geometry and Applications (IJCGA), 20(3):361–380, June 2010. CODEN IJCAEV. ISSN 0218-1959.

Speckmann:2015:ANM

[SV15] Bettina Speckmann and Kevin Verbeek. Algorithms for necklace maps. International Journal of Computational Geometry and Applications (IJCGA), 25(1):15–??, March 2015. CO-DEN IJCAEV. ISSN 0218-1959.

Suri:2016:MLV

[SV16] Subhash Suri and Kevin Verbeek. On the most likely Voronoi diagram and nearest neighbor searching. International Journal of Computational Geometry and Applications (IJCGA), 26(3–4): 151–166, September/December 2016. CODEN IJCAEV. ISSN 0218-1959.

Schlesinger:2016:FSC

[SVY16] M. I. Schlesinger, E. V. Vodolazskiy, and V. M. Yakovenko. Fréchet similarity of closed polygonal curves. International Journal of Computational Geometry and Applications (IJCGA), 26(1):53–??, March 2016. CODEN IJCAEV. ISSN 0218-1959.

Strijk:2001:LPC

[SW01] Tycho Strijk and Alexander Wolff. Labeling points with circles. International Journal of Computational Geometry and Applications (IJCGA), 11(2): 181–195, 2001. CODEN IJ-CAEV. ISSN 0218-1959. URL

http://ejournals.wspc.com.
sg/journals/ijcga/11/preserved-
docs/1102/S0218195901000444.
pdf;   http://www.math-inf.
uni-greifswald.de/map-labeling/
papers/sw-lpc-01.ps.gz.

Shimada:2000:ATP

[SYI00] K. Shimada, A. Yamada, and T. Itoh. Anisotropic triangulation of parametric surfaces via close packing of ellipsoids. International Journal of Computational Geometry and Applications (IJCGA), 10(4):417–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Sitharam:2010:RCC

[SZP10] Meera Sitharam, Yong Zhou, and Jörg Peters. Reconciling conflicting combinatorial preprocessors for geometric constraint systems. International Journal of Computational Geometry and Applications (IJCGA), 20(6):631–651, December 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Tamassia:2003:GEF

[Tam03]  Roberto Tamassia.  Guest

Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 13(1):3–??, February 2003. CODEN IJ-CAEV. ISSN 0218-1959.

Tan:1999:EGS

[Tan99] X. H. Tan. Edge guards in straight walkable polygons. International Journal of Computational Geometry and Applications (IJCGA), 9(1):63–??, 1999. CODEN IJCAEV. ISSN 0218-1959.

Tan:2002:FOB

[Tan02] Xuehou Tan. Finding an optimal bridge between two polygons. International Journal of Computational Geometry and Applications (IJCGA), 12(3): 249–??, 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Tateiri:2023:EEE

[Tat23] Kazuma Tateiri. Efficient exact enumeration of single-source geodesics on a non-convex polyhedron. International Journal of Computational Geometry and Applications (IJCGA), 33(3–4):97–132, September– December 2023. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195923500048.

Teng:2000:GEF

[Ten00] S.-H. Teng. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 10 (3):225–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Tan:1999:CSI

[THI99] X. Tan, T. Hirata, and Y. Inagaki. Corrigendum to “An Incremental Algorithm for Constructing Shortest Watchman Routes”. International Journal of Computational Geometry and Applications (IJCGA), 9(3):319–??, 1999. CODEN IJ-CAEV. ISSN 0218-1959.

Tseng:1998:TGW

[THL98] L. H. Tseng, P. Heffernan, and D. T. Lee. Two-guard walkability of simple polygons. International Journal of Computational Geometry and Applications (IJCGA), 8(1):85–??, February 1998. CODEN IJ-CAEV. ISSN 0218-1959.

Teng:1995:PAV

[TMPD95] Y. A. Teng, D. Mount, E. Puppo, and L. S. Davis. Parallelizing an algorithm for visibility on polyhedral terrain. International Journal of Computational Geometry and Applications (IJCGA), ??(??):??, 1995. CODEN IJCAEV. ISSN 0218-1959.

Teng:1997:PAV

[TMPD97] Y. Ansel Teng, David Mount, Enrico Puppo, and Larry S. Davis. Parallelizing an algorithm for visibility on polyhedral terrain. International Journal of Computational Geometry and Applications (IJCGA), 7(1–2):75–??, February–April 1997. CODEN IJCAEV. ISSN 0218-1959.

Tateiri:2021:EMA

[TO21] Kazuma Tateiri and Toru Ohmoto. An extended MMP algorithm: Wavefront and cutlocus on a convex polyhedron. International Journal of Computational Geometry and Applications (IJCGA), 31(4):221– 247, December 2021. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195922500029.

Tokuyama:2002:GEF

[Tok02]  Takeshi Tokuyama.  Guest

Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 12(4):267–??, August 2002. CODEN IJ-CAEV. ISSN 0218-1959.

Tokuyama:2010:F

[Tok10] Takeshi Tokuyama. Foreword.

International Journal of Computational Geometry and Applications (IJCGA), 20(1):1–2, February 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Tokuyama:2019:GEF

[Tok19]  Takeshi Tokuyama.  Guest

Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 29(1):1, March 2019. CODEN IJCAEV. ISSN 0218-1959. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0218195919020011.

Toussaint:2005:GPG

[Tou05] Godfried Toussaint. Geometric proximity graphs for improving nearest neighbor methods in instance-based learning and data mining. International Journal of Computational Geometry and Applications (IJCGA), 15(2):101–??, April 2005. CODEN IJCAEV. ISSN 0218-1959.

Takahashi:1997:SNC

[TSN97] Jun-Ya Takahashi, Hitoshi Suzuki, and Takao Nishizeki. Shortest non-crossing rectilinear paths in plane regions. International Journal of Computational Geometry and Applications (IJCGA), 7(5):419–??, October 1997. CODEN IJ-CAEV. ISSN 0218-1959.

Tamassia:2001:CSA

[TV01] Roberto Tamassia and Luca Vismara. A case study in algorithm engineering for geometric computing. International Journal of Computational Geometry and Applications (IJCGA), 11 (1):15–70, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Teng:2000:UMG

[TW00] S.-H. Teng and C. W. Wong.

Unstructured mesh generation: Theory, practice, and perspectives. International Journal of Computational Geometry and Applications (IJCGA), 10(3): 227–??, 2000. CODEN IJ-CAEV. ISSN 0218-1959.

Trombettoni:2006:GFA

[TW06] Gilles Trombettoni and Marta Wilczkowiak. GPDOF — a fast algorithm to decompose under-constrained geometric constraint systems: Application to 3D modeling. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):479–??, December 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Tang:2006:MAC

[TWC06] Kai Tang, Charlie C. L. Wang, and Danny Z. Chen. Minimum area convex packing of two convex polygons. International Journal of Computational Geometry and Applications (IJCGA), 16(1):41–??, February 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Unnikrishnan:2010:SSG

[ULVH10] Ranjith Unnikrishnan, Jean-François Lalonde, Nicolas Vandapel, and Martial Hebert. Scale selection for geometric fitting in noisy point clouds. International Journal of Computational Geometry and Applications (IJCGA), 20(5):543–575, October 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Vanecek:1991:BIM

[Van91] G. Vaněček, Jr. Brep-index: a multidimensional space partitioning tree (revised). International Journal of Computational Geometry and Applications (IJCGA), 1(3):243–262, September 1991. CODEN IJ-CAEV. ISSN 0218-1959.

VanDerMeiden:2005:EMD

[VB05] Hilderick A. Van Der Meiden and Willem F. Bronsvoort. An efficient method to determine the intended solution for a system of geometric constraints. International Journal of Computational Geometry and Applications (IJCGA), 15(3):279–??, June 2005. CODEN IJCAEV. ISSN 0218-1959.

Viglietta:2012:SPR

[Vig12] Giovanni Viglietta. Searching polyhedra by rotating halfplanes. International Journal of Computational Geometry and Applications (IJCGA), 22 (3):243–??, June 2012. CODEN IJCAEV. ISSN 0218-1959.

vanKreveld:2018:RDG

[vKLSW18] Marc van Kreveld, Maarten Löffler, Frank Staals, and Lionov Wiratma. A refined definition for groups of moving entities and its computation. International Journal of Computational Geometry and Applications (IJCGA), 28(2):??, June 2018. ISSN 0218-1959.

Vleugels:1998:AVD

[VO98] Jules Vleugels and Mark Overmars. Approximating Voronoi diagrams of convex sites in any dimension. International Journal of Computational Geometry and Applications (IJCGA), 8(2):201–??, April 1998. CO-DEN IJCAEV. ISSN 0218-1959.

VanKreveld:2004:GNP

[VR04] Marc Van Kreveld and Iris Reinbacher. Good news: Partitioning a simple polygon by compass directions. International Journal of Computational Geometry and Applications (IJCGA), 14(4–5):233–??, October 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Wang:2009:RBT

[Wan09] Yusu Wang. Relations between two common types of rectangular tilings. International Journal of Computational Geometry and Applications (IJCGA), 19 (2):161–172, April 2009. CO-DEN IJCAEV. ISSN 0218-1959.

Wang:2015:AMT

[Wan15] Haitao Wang. Aggregate-MAX top-k nearest neighbor searching in the L_{1} plane. International Journal of Computational Geometry and Applications (IJCGA), 25(1):57–??, March 2015. CODEN IJCAEV. ISSN 0218-1959.

Wu:2007:LNS

[WCLS07] Xiaodong Wu, Danny Z. Chen, Kang Li, and Milan Sonka. The layered net surface problems in discrete geometry and medical image segmentation. International Journal of Computational Geometry and Applications (IJCGA), 17(3):261–296, June 2007. CODEN IJCAEV. ISSN 0218-1959.

Wu:2004:EAA

[WCMS04] Xiaodong Wu, Danny Z. Chen, James J. Mason, and Steven R. Schmid. Efficient approximation algorithms for pairwise data clustering and applications. International Journal of Computational Geometry and Applications (IJCGA), 14(1–2): 85–??, April 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Wu:2009:EAOb

[WDBB09] Xiaodong Wu, Xin Dou, John E. Bayouth, and John M. Buatti. Efficient algorithm for optimal matrix orthogonal decomposition problem in intensity-modulated radiation therapy. International Journal of Computational Geometry and Applications (IJCGA), 19 (3):231–246, June 2009. CO-DEN IJCAEV. ISSN 0218-1959.

Weng:2002:GMC

[Wen02] Jia F. Weng. Generalized Melzak’s construction in the Steiner tree problem. International Journal of Computational Geometry and Applications (IJCGA), 12(6):481–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Wang:2021:RSR

[WG21] Haohao Wang and Ron Goldman. Ruled surfaces of revolution with moving axes and angles. International Journal of Computational Geometry and Applications (IJCGA), 31(2– 3):163–181, June–September 2021. ISSN 0218-1959. URL

https://www.worldscientific.
com/doi/10.1142/S0218195921500084.

Wein:2005:CPV

[WIEH05] Ron Wein, Oleg Ilushin, Gershon Elber, and Dan Halperin. Continuous path verification in multi-axis NC-machining. International Journal of Computational Geometry and Applications (IJCGA), 15(4):351–??, August 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Willerton:2015:SMS

[Wil15] Simon Willerton. Spread: A measure of the size of metric spaces. International Journal of Computational Geometry and Applications (IJCGA), 25 (3):??, September 2015. CO-DEN IJCAEV. ISSN 0218-1959.

Wismath:2000:PLS

[Wis00] S. K. Wismath. Point and line segment reconstruction from visibility information. International Journal of Computational Geometry and Applications (IJCGA), 10(2):189–??, 2000. CODEN IJCAEV. ISSN 0218-1959.

Weiss:2020:MOS

[WJA20] Bastian Weiß, Bert Jüttler, and Franz Aurenhammer. Mitered offsets and skeletons for circular arc polygons. International Journal of Computational Geometry and Applications (IJCGA), 30(3–4):235–256, September– December 2020. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195921500023.

Wang:1997:RQP

[WJG97] Wenping Wang, Barry Joe, and Ronald Goldman. Rational quadratic parameterizations of quadrics. International Journal of Computational Geometry and Applications (IJCGA), 7(6):599–??, December 1997. CODEN IJCAEV. ISSN 0218-1959.

Worman:2007:PDO

[WK07] Chris Worman and J. Mark Keil. Polygon decomposition and the orthogonal art gallery problem. International Journal of Computational Geometry and Applications (IJCGA), 17 (2):105–138, April 2007. CO-DEN IJCAEV. ISSN 0218-

Wang:2010:GWR

[WKG10] Pengpeng Wang, Ramesh Krishnamurti, and Kamal Gupta. Generalized watchman route problem with discrete view cost. International Journal of Computational Geometry and Applications (IJCGA), 20(2): 119–146, April 2010. CODEN IJCAEV. ISSN 0218-1959.

Wang:2001:CAM

[WLW01] Jiaye Wang, Ding Yuan Liu, and Wenping Wang. Computing an almost minimum set of spanning line segments of a polyhedron. International Journal of Computational Geometry and Applications (IJCGA), 11(5):475–485, 2001. CODEN IJCAEV. ISSN 0218-1959.

Wulff-Nilsen:2012:CSF

[WNGK^{+}12] Christian Wulff-Nilsen, Ansgar Grüne, Rolf Klein, Elmar Langetepe, D. T. Lee, Tien-Ching Lin, Sheung-Hung Poon, and Teng-Kai Yu. Computing the stretch factor and maximum detour of paths, trees, and cycles in the normed space. International Journal of Computational Geometry and Applications (IJCGA), 22(1):45–??, February 2012. CODEN IJ-CAEV. ISSN 0218-1959.

Wang:2005:ENU

[WQS05] Huawei Wang, Kaihuai Qin, and Hanqiu Sun. Evaluation of non-uniform Doo–Sabin surfaces. International Journal of Computational Geometry and Applications (IJCGA), 15(3): 299–??, June 2005. CODEN IJ-CAEV. ISSN 0218-1959.

Williams:2007:TMS

[WR07] Jason Williams and Jarek Rossignac. Tightening: Morphological simplification. International Journal of Computational Geometry and Applications (IJCGA), 17(5):487–503, October 2007. CODEN IJ-CAEV. ISSN 0218-1959.

Wolff:2002:SFA

[WTX02] Alexander Wolff, Michael Thon, and Yinfeng Xu. A simple factor-\frac{2}{3} approximation algorithm for two-circle point labeling. International Journal of Computational Geometry and Applications (IJCGA), 12(4):269–??, 2002. CO-DEN IJCAEV. ISSN 0218-

1959. URL http://www.math-
inf.uni-greifswald.de/map-
labeling/papers/wtx-sf23a-
01.pdf;   http://www.math-
inf.uni-greifswald.de/map-
labeling/papers/wtx-sf23a-
01.ps.

Wu:2009:EAOa

[Wu09] Xiaodong Wu. Efficient algorithms for the optimal-ratio region detection problems in discrete geometry with applications. International Journal of Computational Geometry and Applications (IJCGA), 19 (2):141–159, April 2009. CO-DEN IJCAEV. ISSN 0218-

Wang:2016:LCM

[WZ16] Haitao Wang and Jingru Zhang. Line-constrained kmedian, k-means, and k-center problems in the plane. International Journal of Computational Geometry and Applications (IJCGA), 26(3–4): 185–210, September/December 2016. CODEN IJCAEV. ISSN 0218-1959.

Wang:2018:CRC

[WZ18] Haitao Wang and Jingru Zhang. Computing the rectilinear center of uncertain points in the plane. International Journal of Computational Geometry and Applications (IJCGA), 28 (3):??, September 2018. ISSN 0218-1959.

Wang:2019:TKW

[WZ19] Haitao Wang and Wuzhou Zhang. On top-k weighted sum aggregate nearest and farthest neighbors in the L_{1} plane. International Journal of Computational Geometry and Applications (IJCGA), 29(03):189– 218, September 2019. CO-DEN IJCAEV. ISSN 0218-

1959.  URL https://www.
worldscientific.com/doi/
10.1142/S0218195919500055.

Wang:2020:LTA

[WZ20] Haitao Wang and Yiming Zhao.

A linear-time algorithm for discrete radius optimally augmenting paths in a metric space. International Journal of Computational Geometry and Applications (IJCGA), 30(3–4): 167–182, September–December 2020. CODEN IJCAEV. ISSN 0218-1959. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0218195920500089.

Xu:2004:TSP

[XLYB04] Jinhui Xu, Zhiyong Lin, Yang Yang, and Ronald Berezney. Traveling salesman problem of segments. International Journal of Computational Geometry and Applications (IJCGA), 14 (1–2):19–??, April 2004. CO-DEN IJCAEV. ISSN 0218-1959.

Xu:2006:DLB

[Xu06] Guoliang Xu. Discrete Laplace– Beltrami operator on sphere and optimal spherical triangulations. International Journal of Computational Geometry and Applications (IJCGA), 16 (1):75–??, February 2006. CO-DEN IJCAEV. ISSN 0218-1959.

Xu:2010:GSS

[XYZK10] Jinhui Xu, Yang Yang, Yongding Zhu, and Naoki Katoh. A geometric spanner of segments. International Journal of Computational Geometry and Applications (IJCGA), 20(1):43–67, February 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Yang:2006:SSC

[Yan06] Lu Yang. Solving spatial constraints with global distance coordinate system. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):533–??, December 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Yang:2023:TPS

[Yan23] Chao Yang. Tiling the plane with a set of ten polyominoes. International Journal of Computational Geometry and Applications (IJCGA), 33(3– 4):55–64, September–December 2023. CODEN IJCAEV. ISSN 0218-1959. URL https://www.

worldscientific.com/doi/
10.1142/S0218195923500012.

Yon:2017:FLC

[YCCV17] Juyoung Yon, Siu-Wing Cheng, Otfried Cheong, and Antoine Vigneron. Finding largest common point sets. International Journal of Computational Geometry and Applications (IJCGA), 27(3):177–185, September 2017. CODEN IJ-CAEV. ISSN 0218-1959.

Yu:2018:CPP

[YLL18] Hung-I Yu, Tien-Ching Lin, and D. T. Lee. The (1|1)- centroid problem in the plane with distance constraints. International Journal of Computational Geometry and Applications (IJCGA), 28(2):??, June 2018. ISSN 0218-1959.

Zak:2010:DPC

[Żak10] Andrzej Żak. Dissections of polygons into convex polygons. International Journal of Computational Geometry and Applications (IJCGA), 20(2):223– 244, April 2010. CODEN IJ-CAEV. ISSN 0218-1959.

Zomorodian:2002:FSB

[ZE02] Afra Zomorodian and Herbert Edelsbrunner. Fast software for box intersections. International Journal of Computational Geometry and Applications (IJCGA), 12(1–2):143–??, 2002. CODEN IJCAEV. ISSN 0218-1959.

Zernisch:2012:GTK

[Zer12]  Jan B. Zernisch.  A gen-

eralization of a theorem of Kleitman and Krieger. International Journal of Computational Geometry and Applications (IJCGA), 22(2):167–??, April 2012. CODEN IJCAEV. ISSN 0218-1959.

Zhang:2006:WCC

[ZG06] Gui-Fang Zhang and Xiao-Shan Gao. Well-constrained completion and decomposition for under-constrained geometric constraint problems. International Journal of Computational Geometry and Applications (IJCGA), 16(5–6):461–??, December 2006. CODEN IJ-CAEV. ISSN 0218-1959.

Zhang:2007:GEF

[Zha07] Li Zhang. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 17(3): 199–200, June 2007. CODEN IJCAEV. ISSN 0218-1959.

Zhu:1997:ACP

[Zhu97]  Binhai Zhu.  Approximat-

ing convex polyhedra with axis-parallel boxes. International Journal of Computational Geometry and Applications (IJCGA), 7(3):253–??, June 1997. CODEN IJCAEV. ISSN 0218-1959.

Zhu:2004:APC

[Zhu04a] Binhai Zhu. Approximating 3D points with cylindrical segments. International Journal of Computational Geometry and Applications (IJCGA), 14(3): 189–??, June 2004. CODEN IJ-CAEV. ISSN 0218-1959.

Zhu:2004:GEF

[Zhu04b] Binhai Zhu. Guest Editor’s foreword. International Journal of Computational Geometry and Applications (IJCGA), 14(1–2):1–??, April 2004. CO-DEN IJCAEV. ISSN 0218-1959.

Zhu:2001:EAA

[ZP01] Binhai Zhu and C. K. Poon.

Efficient approximation algorithms for two-label point labeling. International Journal of Computational Geometry and Applications (IJCGA), 11(4): 455–464, 2001. CODEN IJ-CAEV. ISSN 0218-1959.

Zhang:2006:BSI

[ZWG06] Ming Zhang, Liqun Wang, and Ronald Goldman. Bézier subdivision for inverse molecular kinematics. International Journal of Computational Geometry and Applications (IJCGA), 16 (5–6):513–??, December 2006. CODEN IJCAEV. ISSN 0218-1959.